T.R | Title | User | Personal Name | Date | Lines |
---|
19.4 | HARDWARE.CMP FILE SETTINGS | STRATA::TDYER | | Thu Nov 09 1995 11:30 | 58 |
|
THIS IS THE B2 HARDWARE.CMP CONFIGURATION FILE. SHOULD YOU SUSPECTED
ANY SOFTWARE CONFIGURATION PROBLEMS, VERIFY THAT ALL SETTINGS MATCH THE
ONES BELLOW.....TIM
1 MUTLI-HEAD TRANSPORT ASSY
1 CARRIER 1
1 CARRIER 2
1 CARRIER 3
1 CARRIER 4
1 CARRIER 5
1 INDEX TABLE
1 LOAD FLIPPER
1 UNLOAD FLIPPER
1 ELEVATOR #1
1 ELEVATOR #2
1 PAD CONDITIONER
1 CONTINUOUSE POLISH
1 NITROGEN DURING CARRIER CLEAN
1 CLOSED LOOP
0 PRESSURE SAVED TO FILES
1 SCRUBBER
1 WAFER TRACKING
1 EIGHT INCH WAFER
0 LOF FILES (WARNING, DO NOT SET THIS TO 1)
0 SKIP AUTOMATION
1 VACUUM CHECK
0 UNLOAD STYLE
1 PREALIGNER CENTER WAFER FLAG
1 WAFER LOSS DETECTION
1 CHILLER
1 CHECK AIR PRESSURE
1 CHECK MAIN VAUUM LINE
0 CHECK ROBOT VACUUM
1 CHECK NITROGEN PRESSURE
1 CHECK DI WATER PRESSURE
1 CHECK MIOSTURE SENSORS
1 PAD CONDITIONER RINSE ON
1 LOAD GRIPPER SENSORS
1 UNLOAD GRIPPER SENSORS
1 SECOND POLISH TABLE
1 VAC CHECK DURING INIT
1 PREALIGNER
1 ADE PREALIGNER=0, FIXTURE=1
1 KENSIGNTON MAPPERS
0 DEMO MODE
0 PAD COND UNITS (1=LBS)
1 MOTORIZED PADDLE
0 CHECK SLURRY
0 DEBUG PRINTF STATEMENTS (1=GOOD FOR TROUBLESHOOTING, NORMAL=0)
0 ELEVATOR #1 LIMIT SWITCHES WORK
0 ELEVATOR #2 LIMIT SWITCHES WORK
1 FRONT LOCKS FOR MULTI HEAD
END
|
19.5 | Mechanical Acceptance Activity Update | ASDG::POIRIER | | Thu Nov 09 1995 20:08 | 9 |
|
* Mechanical Acceptance Activity:
The Marathon Run was stopped due to Elevator #2 load problems.
The robot repeatedly failed to pick-up wafers from slot #1 in
Elevator #2. The Speedfam Reps are currently repairing the
elevator.
John.
|
19.6 | ROBOT AND ASSOCIATES LEVELED AND CALIBRATED | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Nov 10 1995 05:42 | 18 |
|
Things going good....
The robot needed to be leveled to the animation
frame as well as getting it parallel to the load
elevators. The robot, aligner, load flipper and
elevators were then re-calibrated
The water track sensors were a little flaky
so the were tweaked to get them to stablize.
50 wafers were run to verify everything. There
were no problems.
After that was done, the marathon started and the
wafer count is just over 100.
D shift
|
19.7 | Mechanical Acceptance Update | ASDG::POIRIER | | Fri Nov 10 1995 19:59 | 11 |
|
Mechanical Acceptance Update:
The system Flew-Out Wafers near the end of the Acceptance run. One
wafer broke and the Pad is, of course, in need of replacement.
The carriers and pad have been removed from the tool. D shift,
please build-up 5 carriers and replace the pad. See Tim Dyer's
passdown for Meacanical Acceptance instructions.
John.
|
19.8 | Marathon stopped | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Nov 11 1995 07:21 | 9 |
|
Marathon was placed on hold three times due to wafer breakage
on the table. Reps are still trying figure what is going on.
The marathon stands at (0). Reps are currently working on the
index table alignment.
D shift
|
19.9 | Marathon run started | FABSIX::R_GEE | | Sun Nov 12 1995 07:06 | 24 |
| Rep's
- Broke the slurry dispense head removing the carriers.
Went to fixture clean and Teflon welded it back together,
looks like new.
- Ran 200 wafers running the no touch program with only 1
error. "Load Flipper fingers wouldn't release the wafer
at the load cup."
- Started running touch down wafers.
Error #1: "head #1 not arriving to top position", had to
adjust the sensor position.
Error #2: A wafer slid out from under carrier #3, didn't
break. Adjusted load cup to carrier for head #3,
it was slightly out of alignment.
- Started the touch down run counter over, we are now at
100 wafers and should hit 125 by 8 am.
B shift
|
19.10 | Mechanical acceptance completed | FABSIX::R_GEE | | Mon Nov 13 1995 07:17 | 11 |
| * The Reps. made a slight adjustment to the load flipper.
150 more wafers were run on B shift tonight with no problems.
* 4 Carriers were built tonight and are in the carrier rebuilt
room. The exposure was set between .17mm and .19mm, this equals
between .007in and .008in. We had problems bringing the tolerances
in on 1 carrier, it's on the table with the depth gauge on it.
* A rush was placed for 20 more retainer rings, we're out of stock.
B shift & Mitch
|
19.11 | passdown | FABSIX::B_FINN | | Mon Nov 13 1995 19:18 | 18 |
|
Engineering began taking particle checks. System broke wafers on pad.
Replaced pad and carriers. System passed particle checks.
One of the carriers was of the correct exposure tolerance, BUT in the
wrong direction of exposure ?????? Positive 7 instead of neg 7-8.
Rep rebuilt one carrier and are currently running initial uniformity
tests and checking correlation of removal data between spindles.
SBA to send 9 pressure plates was cut early this morning. before this
mishap.
No Reps due in tonight. They will however, work late and be in early
tomorrow.
A shift
|
19.12 | install update | LUDWIG::FINN | | Tue Nov 14 1995 19:33 | 84 |
|
The Reps (Darin), calibrated his meter # 405962. He the set up the
head balance pressure. The down force calibration was then performed with the
following results:
Spindle # 1
Original reading was 164.9 lbs on the stress gauge at a set point of 377 lbs.
Original Input coefficient was .94753.
Original Output coefficient was 3.03909
Spindle 1 was adjusted to a meter reading of 377 lbs @ a set point of 377 lbs
New input coefficient is 0.47243
New output coefficient is 5.87533
************************************************************************
Spindle # 2
Original reading was 158.9 lbs on the stress gauge at a set point of 377 lbs.
Original Input coefficient was 0.91642
Original Output coefficient was 3.18241
Spindle 2 was adjusted to a meter reading of 377.1 lbs @ a set point of 377 lbs
New input coefficient is 0.46774
New output coefficient is 6.08488
*************************************************************************
Spindle # 3
Original reading was 186.0 lbs on the stress gauge at a set point of 377 lbs.
Original Input coefficient was 0.85754
Original Output coefficient was 3.29316
Spindle 3 was adjusted to a meter reading of 377.0 lbs @ a set point of 377 lbs
New input coefficient is 0.46832
New output coefficient is 5.97613
**************************************************************************
Spindle # 4
Original reading was 187.7 lbs on the stress gauge at a set point of 377 lbs.
Original Input coefficient was 0.87714
Original Output coefficient was 3.28990
Spindle 4 was adjusted to a meter reading of 377.0 lbs @ a set point of 377 lbs
New input coefficient is 0.47421
New output coefficient is 5.95756
****************************************************************************
Spindle # 5
Original reading was 158.8 lbs on the stress gauge at a set point of 377 lbs.
Original Input coefficient was 0.97152
Original Output coefficient was 3.18893
Spindle 2 was adjusted to a meter reading of 377.1 lbs @ a set point of 377 lbs
New input coefficient is 0.46201
New output coefficient is 6.42971
****************************************************************************
At the completion of the downforce cal, they conditioned the pad and
went on to balance the spindles offset and cancel the effect of individual
carriers.
Darin feels that the our old meter was not calibrated properly in our
CAL LAB. The display should be zeroed then the sensor calibrated to the
display. I have made copies of the entire procedure and will provide this
to our cal lab with feedback... Also kept a copy for us.
I picked up the orings your were looking for and also found 10 22.5"
pads in your name from Rodel; not sure if they were for the secondary table or
fab 4; pads and rings are in the room, paperwork is in your office.
Brendan/Doc
|
19.13 | b2 process acceptance history file | SUBPAC::KRUPA | | Wed Nov 22 1995 11:04 | 304 |
| CMP.B2 Process Acceptance Passdown, Friday, 11/17/95:
CMP.B2 Process Acceptance Passdown, Friday, 11/17/95:
CMP.B2 Process Acceptance Passdown, Monday, 11/20/95
- The primary process acceptance was completed and report was circulated
with results.
CMP.B2 Process Acceptance Passdown, Friday, 11/17/95:
- We believe the wafer breakage problem has been resolved. The problem was
when we landed down onto the pad we were landing and ramping to full pressure
in the ID position where part of the wafer was "off" the pad. This was
verified by processing ~395 wafers on a full pad w/no breakage.
- 5 new carriers were installed along with an ID/OD pad. The landing location
was modified to land in the middle of the ID/OD pad such that the entire wafer
surface is on the pad when landing and ramping to full pressure. 50 wafers
were cycled for 3min to breakin the pad followed by 10 pre-measured wafers.
The results are:
3min oscillation=(110-145) 3min oscillation=(112-147)
slot removal %dev center slot removal %dev center
---- ----- ---- ----- ---- ----- ----- -----
1 8032 6.3 slow 6 7813 4.5 s
2 8010 3.4 slow 7 7818 4.6 s
3 7989 4.9 slow 8 7851 2.7 m
4 7824 8.4 fast 9 7463 9.6 f
5 7892 4.2 slow 10 7616 2.9 med
- At this point carrier #4 was changed out because it was an outlier.
Five breakins were run at 40sec and then 5 pre-measured wafers to check rate
and uniformity 1 more time before starting the 50 wafer 3 minute polish qual.
The results are:
3min oscillation=(112-147)
slot removal %dev center
---- ----- ---- -----
1 7449 5.2 med
2 7519 3.6 med
3 7365 4.2 med
4 7110 7.0 fast
5 7353 3.1 med
- The 50 wafer 3 min polish qual was begun ~6:00. If the results are
comparable to the 2 min run, basic process acceptance will have been completed
and Speedfam will fly home tomorrow. The remainder of November will be spent
optimizing the recipe for the primary platen, developing a recipe for the
secondary platen for particle reduction, and verifying planarization
capabilities. At this point we are on or ahead of schedule to release the
tool to production on 12/1.
******************************************************************************
CMP.B2 Process Acceptance Passdown, Thursday, 11/16/95:
- Things looked good after the pad change and the 50 wafer 2 min polish QUAL
was begun. The run completed with no issues. The results were good. All
specs were met except the overall average head-to-head rate variation(was 6.0%,
spec is <5.0%). Results summary:
Average %std-dev = 4.65%(SPEC <5.0%), max = 6.05%(SPEC 8.0%), min=3.20%
Note that the MAX and MIN ocurred in the 1st 25 wafer run.
The polish rate increased incrementally through the 1st and 2nd run.
Thereafter the rate was pretty stable.
Average removal rate = 2891A/min(SPEC >2kA/min), max = 3045A/min,
min = 2695A/min
The average head-to-head variation across 25 wafer using Larry's
formula was 6.05%(SPEC <5.0%). If the 1st & 2nd run are dropped
the average head-to-head variation ~3%.
However, the head-to-head within-run removal rate was very consistant
and within spec <5.0%. head-to-head- head-to-head-
run rate-%std-dev. run rate-%std-dev.
1 1.25% 6 2.17
2 2.03% 7 2.19
3 1.65% 8 1.09
4 1.58% 9 1.72
5 1.44% 10 2.45
- The 50 wafer 3 min polish test was begun, but, a wafer broke in carrier #2
on the 2nd run. Since this was the same head location that a wafer broke in
yesterday, EE took the tool to check alignment in the load cups. All heads
were found to be slightly high and head 2 was found to off-level slightly.
This may have been the cause for the breakage. A new pad was installed.
Carriers 2&3 were replaced and 2 retainer rings were replaced on 2 other
carriers.
- After breaking in the pad a few test wafers were run and uniformity went
out the window(~4%-11%). Speedfam believes the high uniformity is due to the
3 used carriers that sat for a while and dried. We do not have any more
plates available, however, to install 5 new carriers. So additional breakins
are being performed to see where we stand before committing the 50 wafer 3 min
polish test.
- 50 wafers were run for 30sec and 5 wafers for 3 min to breakin the
pad/carriers more. Then 5 pre-measured wafers were polished for 2 min and
5 pre-measured wafers were polished for 3 min. The results were:
<---------2min----------> <----------3min--------->
1 5440 5.3 s 1 7446 7.1 s
2 4835 11.8 f 2 6687 7.0 f
3 5224 5.0 s 3 7389 4.7 s
4 5156 9.2 f 4 7294 3.8 s
5 5538 4.8 s 5 7399 3.7 s
- We decided to replace carriers 1&2 with new ones and try 5 more wafers.
Upon doing this we BROKE ANOTHER WAFER!!!
- AT THIS POINT WE STOPPED THE PROCESS ACCEPTANCE TESTING. SPEEDFAM WILL WORK
AROUND THE CLOCK CYCLING WAFERS MECHANICALLY IN AN EFFORT TO ISOLATE THE
SOURCE OF THE WAFER BREAKAGE PROBLEM WE'VE BEEN EXPERIENCING SINCE LAST FRIDAY.
PROCESS ACCEPTANCE WILL NOT RESUME UNTIL WE HAVE CONFIDENCE WE HAVE RESOLVED
THE PROBLEM.
- There are 3 new carriers in the carrier build room and 3 additional
carriers being kept wet in the chase for B1 in case of a crash there.
*******************************************************************************
CMP.B2 Process Acceptance Passdown, Wednesday, 11/15/95:
- The carrier offsets were calculated in the manner described yesterday.
The offsets were: head offset
1 0
2 -1
3 -2
4 7
5 -4
- A 25 wafer verification run was performed for 2min. The results were:
The within-wafer non-uniformity was tight across all 25 wafers.
Average %std-dev = 4.07%(SPEC <5.0%), max = 5.75%(SPEC 8.0%), min=2.40%
Important to note that the 1st half of the run polished center-fast,
while the last half of the run polished edge fast.
The polish rate increased incrementally with each run and thus failed
the rate uniformity spec.
Average removal rate = 3088A/min(SPEC >2kA/min), max = 3364A/min,
min = 2904A/min
The average head-to-head variation across 25 wafer using Larry's
formula was 8.5%(SPEC <5.0%).
However, the head-to-head within-run removal rate was very consistant
and within spec. head-to-head-
run rate-%std-dev.
1 1.52%
2 2.10%
3 1.75%
4 1.97%
5 2.26%(SPEC <5.0%)
- We decided to process the 25 wafers at 3min to see if the trend continues.
On the 4th run a wafer broke in carrier #2. We cleaned up the mess enough to
finish the 5th run, however, the wafers ended up being toast. So we changed
the pad and all 5 carriers.
- Offsets for heads 3&4 were changed to -1 & +8 respectively based on the
last run.
- The pad was conditioned twice for 5 minutes at 100% downforce(77lbs). Ten
dummy wafers were processed followed by ten test wafers at 2min. Provided this
loos ok, the 50 wafer 2 min polish test will be repeated.
- Production, please assist Eric B. tonight in measuring wafers.
******************************************************************************
CMP.B2 Process Passdown, Tuesday, 11/14/95:
- Carrier extension was corrected from +6mils to -.75mils on the problem
carrier. Also, the conditioner recipe was found to be incorrect. The
conditioner was set-up to oscillate when it should not. This was corrected
and the conditioner downforce was increased to 100%. 5 wafers were run to
test rate and uniformity. Downforce was 377lbs and polish time was 120s.
The results the rate is still low and the uniformity
poor. They are shown below:
slot process amt.rem %dev c/e fast rate/min
---- -------- ------- ---- -------- --------
1 primary only 2763 7.4 edge 1382
2 " 3002 9.4 center stripe 1501
3 " 3653 8.1 edge 1827
4 " 3783 9.1 edge 1892
5 " 2972 11.2 edge 1486
- The process we're using on B2 was put on B1 and 5 wafers run for comparison.
The results show B1 removal rate to be much higher. The system was handed over
to EE to re-calibrate the downforce. The results of B1 are below:
slot process amt.rem %dev c/e fast rate/min
---- -------- ------- ---- -------- --------
1 primary only 5345 9.5 center 2673
2 " 4997 6.2 center 2499
3 " 4847 4.3 edge 2424
4 " 4338 8.5 center 2169
5 " 5089 6.6 center 2545
note: B1 had failed regular uniformity MQCs prior to this test.
- Darren found the downforce mis-calibrated. When asking for 377lbs we
were only getting ~170lbs. Also, the conditioner max. downforce was increased
from ~55lbs to 77lbs. This appears to be the cause for the low removal
rates observed in the previous test because the test performed after the
re-calibration of the downforce are more comparable to B1. The results are
shown below.
slot process amt.rem %dev c/e fast rate/min
---- -------- ------- ---- -------- --------
1 primary only ? ? wafer reclaimed before I could measure
2 " 5170 10.4 center 2585
3 " 5358 10.8 center(best profile) 2679
4 " 5197 7.7 center 2599
5 " 5266 9.9 center 2633
- The 25 wafer head offset characterization run was started(5 wafers are
processed, then the carriers are rotated 1 head and 5 more wafers run, etc.).
The process used was 377lbs downforce(touchdown is performed at full force
with 15prm table speed) for 2min at 29/29rpm, 1L slurry, 15"ID/OD pad,
oscillation 110-145 at .3"/sec.
The conditioning recipe was 80% downforce(~62lbs) for 60sec.
The results will be placed in Excel and offsets will be calculated and plugged
in to the tool. Confirmation runs will follow.
******************************************************************************
CMP.B2 Process Passdown, Monday 11/13/95:
- Particles were checked today. Mechanical particles & process particles both
w/and w/out second table were checked on ID(15")/OD pad configuration.
Process used was 377lbs, 29/29rpm, 107-145 oscillation, 75sec primary platen,
10sec secondary platen.
Results from Particle tests
slot process pre post delta
---- -------- ----- ---- -----
25 transfer only, no table 12 15 3
24 " 20 36 16 (Spec <20 adders)
23 " 16 25 9
22 dummy
21 dummy
24 primary table 75sec 38 679 641
23 " 16 451 435 (Spec <5000 adders)
22 " 23 2554 2531
21 dummy
20 dummy
14 primary 75s/2nd 10sec 10 413 403
11 " 21 482 461 (Spec <5000 adders)
6 " 14 937 923
2 " 34 751 717
1 dummy
Uniformity results from process particles without second table
slot process amt.rem %dev c/e fast rate/min
---- -------- ------- ---- -------- --------
24 primary 75s 1741 8.4 edge 1392
23 " 1778 20.3 hill 1422 (Spec >2000)
22 " 1847 12.4 center 1477
21 dummy
20 dummy
14 primary+2nd 1797 7.5 edge 14 38
11 " 2685 7.4 edge hill 2148 (Spec >2000)
6 " 1917 12.0 center 1533
2 " 1804 9.4 edge 1443
1 dummy
- Wafer crashed on pad when water was accidentally flowed on table during
subsequent processing. Installed new ID/OD pad, 5 new carriers and ran break-
ins. Five wafers were run for 120sec to check rate and uniformity. Head #2
was found to have 0 removal rate. It was determined that head #2 exposure was
set at +5.5mils instead of -5.5mils...ie the wafer was not exposed at all.
Speedfam reps are in the process of rebuilding that carrier.
- Plan is to recheck rate and uniformity and the Speedfam PEs will go home.
The carrier offset characterization is scheduled to begin tomorrow morning
provided we do not run into any more snags.
LIST OF LOT NAMES FOR CMP.B2 START-UP:
LOT I.D. PRE-MEASURED TOOL BY WHO? FILE #'S USED?
-------- ------------ ---- ------- -------- -----
PA0830 YES A5 Earl 2195-2219 YES
PA0833 YES A5 Earl 2220-2244 YES
PA0834 YES A5 FRANK 2313-2337 YES
PA0835 YES A5 FRANK 2338-2362 YES
PA0837 yes A5 FRANK 2369-2381 YES
2458-2470
PA0717 YES A5 JOHN YES
PA0831 YES A5 FRANK YES
PA0832 yes NO
PA0836 yes YES
PA0712 YES A5 JOHN YES
PA0797 NO
PA0657 NO
PA0705 NO
PA0656 YES
PA0893 YES
|
19.14 | slurry supply problem corrected | SUBPAC::LANDRY | | Mon Nov 27 1995 18:48 | 13 |
| The slurry supply system went down early this morning. The pump in the
distribution unit down stairs kicked off for excessive pressure. This
is normally caused by a "dead-head" situation. So, we opened the by-
pass valve in the end VMB to allow flow to increase in the loop. The
pump down stairs remained running. As the valve was closed down to
supply pressure to the line, we were able to get the pressure up, but
there wasn't sufficient flow. Found the filter clogged up pretty bad.
The bottom fittings were missing from the filter. I'm sure the air
exposure was a vital detriment in the clog situation. A new filter was
installed and flow was verified. The system is now available for
process development work to resume.
A shifters and Tim D
|
19.15 | passdown | STRATA::FINN | | Tue Nov 28 1995 19:21 | 26 |
|
Load cup 5 failed to descend upon loading. It appears that the
pedestal was not within the loadcup double L bracket. This caused the cup to
ascend, but not to descend. The multihead was allowed to move out to polish
even though loadcup 5 was up, due to the fact that the position sensor
monitors the pedestal and not the actual cup. The ascended cup 5 struck
carrier 5 breaking the wafer at the index table. After cleaning
Loadcup 5 and aligning pedestal within loadcup 5, rep (Bob G.) checked
function of cups up and down several times without issue. Cleaned system
and replaced both primary and secondary pads. Put new carrier on head 5.
Bob feels that this loadcup was improperly placed on top of the pedestal at
the completion of the labeling(stamping) process. Frank confirms that the
system has not been run since that event.
In the late afternoon the system snap the pedestal off the unload
station. No one present knows why this happened. Bob G. will have a
new load pedestal here tomorrow....
Brendan/Bob G./Frank/Tim
|
19.16 | Unload Cup Pedistal replaced / Load cup #3 sensor error | ASDG::POIRIER | | Thu Nov 30 1995 19:33 | 16 |
|
Unload Cup Air Cylinder Pedistal:
The Unload Cup Pedistal was replaced by Bob Gugerty and myself
this afternoon. The Unload Flipper was re-taught a new pick-up
position due to the slight variation in the Pedistal height.
Tested o.k.
Load Cup #3 Load cup sensor problem:
The system errored for "all cups not down" at 7:15 p.m.
Digital input #94 was not up. D shift, take a look at the sensor
on load cup #3 air cylinder.
|
19.17 | CUP SENSOR ADJUSTMENT | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 08:38 | 12 |
|
Load Cup #3 sensor wasn't detecting that the pedistal
was done when it was. I adjusted the sensor and all is
well...
while working on #3, I noticed #4 unload cup could free
spin in the hole. The dowel pin was pushed in. I pushed
it back out and checked the other cups. Frank is aware of
it and will be talking with Bob Gugerty about it.
BJ
|
19.18 | Unload air cylinder extension rod replaced. / Bent | ASDG::POIRIER | | Fri Dec 01 1995 19:39 | 6 |
|
* Unload cup air cylinder extention rod:
The rod was slightly bent which caused the sensor to detect
intermittently. The Extension rod was replaced an the unload
flipper pick-up position re-taught.
|
19.21 | passdown | FABSIX::B_FINN | | Wed Dec 06 1995 19:16 | 8 |
|
System is down for a bad infornor. Replacement is
due in tomorrow. Refer to Tim's passdown.
|
19.22 | Collision between spindle #3 and load cup #3 / Major Damage | ASDG::POIRIER | | Thu Dec 07 1995 19:33 | 24 |
|
* Collision between spindle #3 & load cup #3:
------------------------------------------
The tool experienced a major crash at 7:15 p.m. this evening. Spindle
#3 collided with load cup #3. The entire load cup assembly was torn out
of the index table. The load cup is completely destroyed. The index
table pulpier tub also sustained damage. There is a large piece broken
off the side of the tub. The load cup mounting hole has been distorted
from its original shape.
Unknowns at this time:
1) Check carrier #3 for possible damage.
2) Inspect the whole index table assembly for further damage.
3) Check spindle #3 for any possible damage.
4) Check #3 index table air cylinder assembly for damage.
5) Spindle alignments and index table alignments are most likely out of
spec. Check them.
|
19.23 | System crash | FABSIX::B_JUBINVILLE | | Fri Dec 08 1995 05:17 | 20 |
|
After reviewing the entire crash, we found that the multihead
crashed into load cup #3 twice. Once when the arms swung in
after picking up wafers and then going to the center of the index
table. The other is when the multihead moved to the polish table
after it had the wafers. The wafer on head #3 crashed inside of the
carrier.
Upon removal of the index table, we notice that the pedistal
sensor showed exposed shielding and the wires were shorting
against the sensor block. The tool thought the cup was always
down. The cup came up and got stuck coming down. Since the tool
thought it was down, it moved the multihead to polish.
The sensor was removed and will be replaced when it comes in.
A new index table is being sent in from Texas. The entire cups
assembly on L3 needs to be replaced.
BJ
|
19.24 | Index Table Replacement | ASDG::POIRIER | | Fri Dec 08 1995 19:29 | 9 |
|
* Index Table:
-----------
A replacement Index Table and Load Cup are due to arrive tonight,
by currier, at approx 2:00 A.M. The Load Cup Inserts I.D.'s will be
measured and recorded tomorrow. Tim Dyer will take this information
to a Machine Shop and have each Load Cup Shaft turned down to match
its respective insert. See Tim's Passdown for further details.
|
19.25 | Index Table Replacement/Load Cup Sticking Prob Action Plan | ASDG::POIRIER | | Sat Dec 09 1995 18:51 | 14 |
|
Index Table Replacement:
-----------------------
The replacement Index Table Arrived this morning. It was cleaned and
bolted onto the Index Table Mounting Supports. The damaged proximity
sensor and cable for Air Cylinder #3 were also replaced.
All 10 Load Cup Bushings were measured for runout and taper. The
results varied from one bushing to another. The runout varied from
.001" on the best bushings to .005" on the worst. The measument data
has been forwarded to Tim Dyer per his corrective action plan dated
12/9/95. The Load Cup Shaft measurment activity is planned for
tomorrow. Contact Tim Dyer for complete details.
|
19.26 | Index table alignments | FABSIX::R_GEE | | Mon Dec 11 1995 06:53 | 20 |
| * The field rep. came in to work on the system tonight.
We:
- removed the index table to repair the threads
on 1 of the holddown bolt holes, re-installed
and aligned.
- aligned the index table home sensor.
- set the index table home position and checked
repeatability.
- aligned spindle to load cup positions for all 5 heads.
- checked the load flipper to load cup position.
- checked the unload flipper to unload cup position.
B shift
|
19.27 | cups machined | FABSIX::B_FINN | | Mon Dec 11 1995 19:34 | 7 |
|
Disassembled cups and shafts. Tim had all cups trued to .012 of cup
hole. Cups are back in room. I have begun reassembling them. All components
are labeled.
Brendan\Tim\Bob
|
19.28 | Load / Unload cups | FABSIX::R_GEE | | Tue Dec 12 1995 06:51 | 19 |
| - Cleaned and re-assembled the load & unload cups.
- Placed the cups on the index table in there designated
feed-thru holes. Checked to verify that the air cylinder
mushroom pedestals would clear the cup, looked good.
- Filled with water and checked for smooth operation, all cups
traveled very smooth. Found that all cups were traveling up
to fast, splashing water out of the index table. Turned the
air pressure to all cups down from 45 psi to 25 psi, the same
as CMP.B1.
- Waited 4 hours and checked the up/down travel again, all looks
good.
Didn't run the no touch program due to no carriers available.
B shift
|
19.29 | passdown | LUDWIG::FINN | | Tue Dec 12 1995 19:25 | 9 |
|
Cycled 115 wafers through the system, transfer only no touch downs.
Tim changed pad and is now running "MEC2_acp" quick polish program. These
wafers should be able to be reused several times due to the low removal rate.
Please continue to cycle until tomorrow or wafers are used up. Keep system in
wet mode when not in use, this will make Tim and Frank happy campers.
Brendan
|
19.30 | Load Cup #3 misload/Springs replaced | ASDG::POIRIER | | Wed Dec 13 1995 19:20 | 6 |
|
* Load Cup #3:
-----------
Cup #3 had a mis-load incedent this morning. We disassembled the
cup and replaced the springs. We have since run 50+ wafers without
any problems.
|
19.31 | Released to manufacturing | FABSIX::R_GEE | | Mon Dec 18 1995 07:07 | 4 |
| * Eric B. released the system to manufacturing tonight.
B shift
|
19.32 | tcu resevoir low | SUBPAC::LANDRY | | Tue Dec 19 1995 18:17 | 14 |
| An abnxious little alarm was heard this morning near this system.
Alas, it's only the TCU. The glycol level in the resevoir was slightly
low, causing the alarm to go off/come on. Topped of the resevoir. All
is well.
Later on in the day, the system failed to pickup at wafer (head 4) at
the polish table. The wafer was placed back on the carrier and the
cycle was resumed. When placing the wafer on the carrier, it was quite
evident that there was more than sufficient vacuum, as the wafer was
literally lifted from the placee's (Jim Sadin) hand. Other than that,
the system ran flawlessly.
the A shift
|
19.33 | misloading | STRATA::RIDLON | Eliminate the obvious | Wed Dec 20 1995 03:57 | 17 |
| Machine misloaded three times tonight. Twice it double loaded cup four
and would have done it a third time but the operator saw it and emo'd
the machine. It looks like the load flipper didn't release the wafer
into the load cup and then came back to the aligner with the wafer. It
then picked up the wafer that was on the aligner and the wafer it
brought back and deposited them into the load cup, like I said this
happened twice. The reason it didn't alarm for flipper not open and any
other alarms associated with this is because the load flipper sensors
and unload flipper sensors were disabled. We reenabled them and then
exercised the load flipper manually to ensure it wasn't sticking and
that the sensors were working properly (dig-in 2 and dig-in 3 mainly).
Ran 35 wafers through using the same recipe that production was using
(I/O pmda) with the polish times zeroed and carriers left up. The
problem didn't duplicate. The fab might be closing at 4am (its now
3:55) but if it doesn't we'll continue to run dummies through until the
day shift crew gets in.
|
19.34 | more on misloading | STRATA::RIDLON | Eliminate the obvious | Wed Dec 20 1995 07:38 | 14 |
| Ran come more wafers through the machine to try and duplicate the
misloading problem. Fortunately what appears to be the problem occured.
What happened is the load flipper picked up the wafer from the aligner
and dropped it into the load cup. BUT when it came back to the aligner
the fingers were closed enough so that they hit the wafer that was
there. With the load flipper sensors now enabled the machine alarmed
and halted for load flipper time out at aligner. We removed the cover
on the flipper but everything looked good. Nothing was binding and the
movements were smooth. Air pressure was good at 50 psi. We then put a
gauge in the air line that opens the flipper to see if the pressure
would fluctuate while running but it remained solid at 50psi. Since
it now alarms when it screws up we put the machine back up.
B shift
|
19.35 | load gripper/double loading | STRATA::TDYER | | Wed Dec 20 1995 15:56 | 11 |
|
Continued from B shift on load gripper finger problem. The source of
the double loading wafers has been atributed to the fingers sticking
closed. B shift installed a pressure gauge inline with the grippper.
The gripper was cycled, the fingers opened slowly and made a squeeking
niose, also pressure was at 40. B1 fingers move fast and the pressure
is at 60. Increased B2 gripper pressure to 60. Fingers move quicker
but still squeeked. Cleaned and olied giuld rails. Cycle 50 waferes
with out error. Returned to MFG.
|
19.36 | Vacuum problems and crimped wire | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Dec 21 1995 06:59 | 16 |
|
The tool was going through MQC's due to uniformity and
high particles. Gwen cleaned the entire machine and tried
running MQC's again. It failed again. After she tried
running some more wafers, there was a MAIN VACUUM error.
I found the vacuum pump breaker tripped. I found what
appeared to be dried slurry on the black shield that covers
the capacitor. I touched the the black wire A-71 and an
arc shot out and the machine powered down. Breaker #7, which
pretty much controls the boot-up, was tripped. A-71 had a
bad crimp and the wire simply fell out. I cut it the wire
back and re-crimped it. The tool came back up fine and
the NO_TOUCH program was run to verify everything. MQC's
are back on their way.
BJ
|
19.37 | Load Flipper Mis-Load / SV #65 Sticking Open. | ASDG::POIRIER | | Fri Dec 22 1995 17:14 | 17 |
|
* Load Flipper mis-load:
---------------------
The Load Flipper failed to travel all the way down to the pre-aligner
this afternoon. It was short by at least 1/4". Earl rebooted the
system and the problem was gone.
* Sticking Solenoid Valve:
_______________________
SV #65, Carrier #5 Rinse, stuck open once during C shift today. We
gently tapped on the valve body with a screwdriver handle and the
valve closed.
Note: Brian Jubinville observed both of the above conditions last
night on D shift.
|
19.38 | Start-up for 12/29 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 29 1995 15:10 | 12 |
|
The DI came on but is still out of spec so we didn't want
to keep it to the tool so we shut off the DI feed and opened the
by-pass. Facilities still have some by-passes opened, that's
why we don't have very high pressure.
The slurry lines are all back to normal and slurry is flowing.
We put the filter on but but we didn't run the slurry through
the machine because of the DI situation.
Mitch,John,Leon,BJ
|
19.39 | CMP.B1 START-UP COMPLETE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sun Dec 31 1995 14:35 | 18 |
|
CMP.B2 START UP CHECK SHEET
+++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++
1) RE-CONNECT DI WATER B/C/D-12/31/95
2) POWER ON SYSTEM B/C/D-12/31/95
3) CYCLE/RUN ALL DI FUNCTIONS FOR 15 MINUTE B/C/D-12/31/95
4) WIPE DOWN SYSTEM CLEAN ALL DRIED SLURRY B/C/D-12/31/95
5) DRY TABLES, INSTALL PADS
6) INSTALL SCRUBBER BRUSHES B/C/D-12/31/95
7) INSTALL CONDITIONER B/C/D-12/31/95
8) RUN WET IDLE PROGRAM FOR 15 MIN. (ENOUGH TO CLEAN CARRIERS) B/C/D-12/31/95
9) REPLACE SLURRY FILTERS B/C/D-12/31/95
10) CONTACT ASHLAND CHEMICAL, REQUEST DI TO BE PURGE FROM B/C/D-12/31/95
SLURRY SYSTEM AND CHARGED WITH SLURRY
11) CALIBRATE SLURRY FLOW B/C/D-12/31/95
12) PUT IN WET IDLE AND TURN OVER TO MFG FOR MQCS
|
19.40 | High unif = bad carriers | STRATA::TDYER | | Tue Jan 02 1996 15:32 | 6 |
|
CMP.B2 - Continued with high uniformity. The carriers were removed and checked
by Pro Eng (TJB) on the P2. The curvature was ~ -15u, it should be
closer to -2u. The carriers were replaced, the bad plates marked
as bad.
|
19.41 | polish table/area banging noise | SUBPAC::LANDRY | | Tue Jan 02 1996 19:31 | 12 |
| Late afternoon / early evening, the system began to make a very loud,
banging sound in the polish table area. The sound would occur
consistantly (at about 5 seconds after the carriers landed on the
table and the table spinning at it's lowest speed). The noise would
disappear after that and the run would complete fine. The noise is
definately coming from the polish table or associated areas underneath.
While trying to trace this occurrance, the problem miraculously went
away. The system is running fine......for now.
Doc / Guugs
|
19.42 | table banging/polishing vibration | SUBPAC::LANDRY | | Wed Jan 03 1996 19:32 | 30 |
| Today, due to the lack of pressure plates for the carriers, the
carriers were removed from this system and installed on B1. This was
justified by the polish table banging problem that needs to be resolved
on B2.
With no good plates in-house, we installed a set of carriers that had
DF film (black used on Strasbaughs) to assist in the trouble shooting
of the banging noise. We used the I_O_MQC recipe for the test as that
was the recipe the problem first showed on. Ran a cassette of wafers
with no banging sound. Tim D came in and continued on, the problem
re-occurred. He tried decreasing the ramp time (more stress on the
ramp) to 4 and it happened right off the bat. The time was then
increased to allow a slower ramp to occur and the banging ceased.
BUT there was vibration during the polishing cycle. The pad may (or
may no) be the problem of this vibration as it was stressed along with
the belt that spins the table.
So, the pad has been removed and a new one installed. The five DF'd
carriers will need to be re-installed. Then the system can be broken
in and checked out. If banging occurs at the ramp rate of 5 seconds
(in segment 3 or 4.....it escapes me at this typing which is the right
one), increase the ramp rate to 8 and try again. To recap, we need to
see if the banging stays away at a slower ramp rate and also to see if
the vibration has been eliminated.
Doc / Tim D / Guugs
As I type, Bob Gugurty is making a minor adjustment to the belt tention
to assure proper setup.
|
19.43 | Vibration gone, #5 rinse solenoid back with problems | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Jan 04 1996 08:22 | 17 |
|
Bob loosened the belt and the vibration stayed. He
tightened the belt back up and the noise went away. Bob
ran 50 wafers and the problem with the belts went away.
The tool is still waiting for plates/carriers.
The #5 carrier rinse is staying on again. While it
was spraying, DIG-OUT 83 was NOT lit. I was about to check
the voltage and it the spray stopped before I could grab the
meter. I followed through the wiring and everything on that
end looks good. Please keep an eye on it.
BJ
|
19.44 | 1/3 B SHIFT PASSDOWN - VIBRATION, SOLENOID | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Jan 05 1996 01:15 | 35 |
| rom: FABSIX::R_GEE "Bob Gee, Beeper #184 03-Jan-1996 0808" 3-JAN-1996 08:12:00.39
To: @CMP_ALL
CC: R_GEE
CMP.B2
------
* Rinse for carrier #5 wouldn't shut off.
Replaced the solenoid for carrier #5 rinse, it
was sticking.
* A loud banging noise and vibration was coming from under
the polish table 15 sec. into the polish cycle.
Found the drive belt from the motor pulley to the
idle pulley assembly was slipping at the idle pulley.
Adjusted the idle pulley assembly to keep the belt
from slipping. Had to shim the right side of the bottom
adjustment bracket to get both belts to tighten evenly.
The belts needed to be tightened more then the spec
advise to keep them from jumping the pulley teeth.
The table now vibrates badly, the belt tension needs
to be rechecked before it is run. The cause of the
problems seems to be in the belt from the motor or
in the idle pulley assembly.
Field service was consulted by phone last night
and should be here this morning, weather permitting.
B shift
|
19.45 | Wafer breakage and uniformity | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Jan 05 1996 07:14 | 15 |
|
A wafer slid of carrier #2 and broke. The machine errored
right at the beginning of Step 3 and it appears that the wafer
just wasn't seated properly. Carrier #2 was the only one
damaged. The tool was cleaned up and MQC'ed with (4) heads
since there are no more plate in. One of the wafers came out
with a small scratch and the uniformity was out on one head
(4) more wafers were run and there were no scratches. The
uniformity was out but .8 on one head. Rather than rip the
pad off and do things all over, they are going to run an
exercise wafer in the bad slot and run the tool. Day shift
will check it out and run it if they feel it is ok....
D shift
|
19.46 | wafer loss detection problem | SUBPAC::LANDRY | | Sun Jan 07 1996 19:26 | 10 |
| The system just went back down for a wafer detection problem. It's
a phantom detection, seeing the beams are not being broken by a lost
wafer or the carrier itself.
I have resat the board and it's connector, but a red LED intermittently
illuminates. The lit LED isn't on one of the five heads, rather it is
the next LED down (which has a label saying wafer lost). Could be a
flakey sensor/adjustment problem.
Doc
|
19.47 | wafer loss detect | LUDWIG::RIDLON | Eliminate the obvious | Mon Jan 08 1996 06:06 | 21 |
| Ops tried to run B2 again and it failed again for wafer detection. The
same problem that A-shift saw. At first the wafer lost led was dimly
lit. We couldn't find a cause for this other than something on the
board dragging the 24volts thats on the dig-in 20 line when everything
is fine. When a wafer loss is detected dig-in 20 is supposed to go low
which makes the wafer lost led turn on bright plus opto isolater for
dig-in 20 turns on. We basically banged our heads against the wall
trying to find out why dig-in 20 24v signal was being dragged down to
19.9vdc but couldn't come up with anything. We pulled the line out of
BB#9 pin 28 and the voltage goes up to 24v (actually 23). We found a
spare wafer detect board and put it in but this didn't change anything.
We also had to replace relay k1 on the board to fix the main motor
contactor interlock circuit. Now the wafer loss led stays brightly lit
all the time. We must have shorted something out because U2 (a dual
input nand gate chip) isn't working right. Pin 8 stays low no matter
what pins 9 and 10 do and the motor enable led doesn't light up but the
heads still rotate. This has been a viscous circle and its time to
almost go home. We did call speed fam but when they returned our call
the rep was at home, didn't have prints, and wasn't supposed to be on
call so he suggested that we just bypass it. Not enough work to justify
that.
|
19.48 | wafer detection board | SUBPAC::LANDRY | | Mon Jan 08 1996 17:40 | 12 |
| The wafer detection problem still exists. It is strongly believed that
the wafer detection board has a problem, but it could also have wiring
or 24V power supply problems. This is due to the fact that B1's board
also showed a red LED was lit (yup, the sixth LED down). A board,
cable, and technical assistance should be here tomorrow or the next
day. And yes, B2's board would NOT work in B1.
Currently, the wafer loss detection circuitry has been by-passed for
now. The system is in Wet Mode. NO product should be run without
approval. Start with Tim Dyer should a manufacturing need arise!
Tim / Doc
|
19.49 | wafer detection situation update | SUBPAC::LANDRY | | Tue Jan 09 1996 19:37 | 7 |
| The parts arrived this afternoon for the wafer detection circuit.
Please hold off on installing as we will have the reps perform this
task tomorrow. Yes, the board is bad, but something deeper is
taking it out. We have already lost two boards this week!
Doc
|
19.50 | Wafer detection problem | FABSIX::B_WESTPHAL | | Wed Jan 10 1996 19:37 | 6 |
| CMP.B2 - Down for problems w/ the wafer detect, rotation errors on head #1, and
intermittent problems w/ the gripper. Speedfam reps were on site
addressing these issues. A board for the wafer detection system was
replaced and some wires on Terminal boards were found to be loose. the
problem appears to be fixed and the tool is now at verify. The Reps
will return tomorrow morning to continue working on the other issues.
|
19.51 | Cup alignment, wafer crash | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Jan 11 1996 06:55 | 18 |
|
Carrier #3 was not able to load the wafer everytime.
After checking the alignment, we found the head to load
was out 1/2". How it got that far?? After a while the
load cup #3 was misloading. The dowel pin was pushed in
and was barely holding the cup straight. The cup had rotated
a little and the load flipper arm was hitting on the way
down. The pin was pushed back out and the other cups were
checked as well.
Later this morning, a wafer from #2 slid out and took out
carrier #1. The head was checked and no problems were
observed. The mess was cleaned up and the tool is ready for
break-in.
D shift
|
19.52 | Vacuum System Upgrade | ASDG::POIRIER | | Thu Jan 11 1996 19:21 | 8 |
|
* Vacuum system Upgrade:
---------------------
The Speedfam reps removed the check & needle valves from the pnumatic
cabinet and replaced them with a solenoid bleed-off valve & filter.
The vacuum has increased from 24 in/hg to 29.5 in/hg.
They are planning to perform the same upgrade on CMP.B1 Tomorrow.
|
19.53 | no water for the dog. | STRATA::RIDLON | Eliminate the obvious | Mon Jan 15 1996 01:43 | 12 |
| Last week we noticed that the pad condition spray station rinse
(dig-out 53) wasn't coming on at all. This is the water that fills the
dog dish. On B1 its continuously flowing which makes sense but on B2
its always off. Last week I turned on dig-out 53 and the rinse did come
on and overflowed. This week its off again. I turned it back on and
then changed recipes to see if thats whats turning it off and it is. I
checked the hardware enable screen but there's only "pad condition
rinse - on/off" there and this controls di to the spray nozzle on the
conditioner. The control for this must be in dos. If someone turns it
on please put in notes how you did it.
B shift
|
19.54 | scratches | LUDWIG::RIDLON | Eliminate the obvious | Tue Jan 16 1996 05:56 | 15 |
| Scratches on wafers. MQC's ran twice. Ran a wafer into load cups and
out with no scratches. Changed pad and cleaned conditioner. Ran carrier
blow off, rinse and rotated for about an hour. During the blow off,
after I finished reassembling the conditioner, I put the arm back down
using the service screen. This makes the multihead move back to the
pad ( I didn't know this ) and blow remained on. After moving the
multihead back to the index table the pad was inspected and a few black
pieces were found on the pad. They looked liked bits of wafer. These
were cleaned off with the di gun. After blow off was turned off the
multihead was moved back to the pad. The water in the index table and
the area around the second table had alot of tiny black pieces that
looked like wafers. The water was drained and the table cleaned.
Breakins are running now and mqc's should be done soon.
B shift
|
19.55 | scratches gone......carriers suspected | SUBPAC::LANDRY | | Tue Jan 16 1996 19:29 | 14 |
| Came into today to find the system still scratching wafers. Some
expirements were run (see Tim D's shift passdown for indepth details).
We decided to change the pad, empty and clean the index table, remove
and clean the carriers (totally breaking them down), flush the blow off
out, and scrub the system real good.
Found debris (silicon granduals, etc) inside the carrier gimbal area.
Thouroughly cleaned them out and re-assembled. Reran breakin and
MQC's on the new pad & carriers. One small scratch was found on just
one wafer. The system passed MQC's and is being returned to
manufacturing.
The entire A shift
|
19.56 | Carrier #3 rotation failure / #5 uniformity | FABSIX::R_GEE | | Wed Jan 17 1996 07:08 | 24 |
| -------* Carrier #3 failed to rotate, errored on the screen,dcx board and
the infranor.
- checked the connectors on the infranor, looked good.
- checked the ribbon connector on the dcx board, looked good.
- swapped the module in the dcx board with the one from position
#6 on the board, carrier #5 rotation, no change.
- swapped the infranor with one found in the cabinet, no change.
* Released the system back to production with carrier #3 disabled
do to the critical need for the system.
------* Carrier #5 is running with dummy wafers do to high uniformity
10.6%. Another mqc will be run when time allows.
B shift
|
19.57 | scratches on wafer | FABSIX::B_FINN | | Wed Jan 17 1996 19:32 | 13 |
|
Scratches returned, but was traced to contact with cup/carrier, Cup 3
down sensor was replaced. Rotation problem linked to bad cable see Tims
passdown. System running 4 heads until wip is lowered. System passed its last
MQC. A second MQC is currently being run at P Eng. request. If this fails refer
to Tim's passdown.
IF THERE ARE ANY HARDWARE PROBLEMS ON B2, PAGE DARREN, 800-823-6094
A shift
|
19.58 | Carrier #1 error, low slurry flow | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Jan 19 1996 07:04 | 20 |
|
The Carrier #1 not rotating error has been popping
up just before the carriers go to pick up the wafers.
Assists have been logged lately and I spoke to Darren
about this. A possible dwell time is a possible problem.
I logged the warning so the million assist will cease.
This will have to be taken care of ASAP.
A wafer crash occured in the early morning. An exercise
wafer slid out of head #1 (uh oh) and slammed #5 product
wafer. The tool was cleaned up and break-ins were started.
The moose mating call was evident at the beginning of the
polish cycles. The slurry flow was checked at 700 ml/mn on
both pumps. #1 came in at 500 ml/mn and #2 at 600 ml/mn.
We looked into the slurry lines around the pumps and they
were very worn. We replaced the lines and the flows came in
at 700 ml/mn on #1 and 775 ml/mn on #2. Break-ins were started
back up and Bulwinkle wasn't seen nor heard.....
D shift
|
19.59 | passdown | FABSIX::B_FINN | | Fri Jan 19 1996 19:13 | 8 |
|
Slurry died again. Ball valves opened to provide more slurry to the
end of the line (b2). See Tims passdown if it happens again. Slurry delivery
nozzle broke, Brad poly weld and reinstalled. System is currently up and
running 4 heads. Head 3 needs rotation cable repair when manufacturing puts
the system system down after B1 comes up.
C shift
|
19.60 | Monthly PM started, | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Jan 20 1996 05:32 | 10 |
|
The monthly PM was started but not finished. They can't
give up the tool at this time. The completed items have been
notes in the checksheet I placed.
The carrier #1 not rotating error has not been seen all of
D shift. Was this repaired? If so, please remove the WARNING
on it.
BJ
|
19.61 | FROM SPEEDFAM DUDES | SUBPAC::CMP | | Sat Jan 20 1996 17:11 | 5 |
| CHANGED THE RESOLVER CABLE FOR ARM 3 ROTATION....IT IS NOW AVAILBLE FOR
USE....
IN RESPONSE TO THE CARRIER 1 NO ROTATION...THERE IS NOT ANYTHING I CAN
DO ABOUT THIS RIGHT NOW....WE HAVE TO WAIT UNTIL THE NEXT REVISION OF
SOFTWARE IS AVAILABLE..SORRY
|
19.62 | Slurry Line worn/ resolver cable replaced. | FABSIX::B_WESTPHAL | | Sat Jan 20 1996 17:57 | 4 |
| CMP.B2 - Down for a crack in pump #1's slurry line. The reps. already had the
slurry line replaced by the time I got in. Production let us keep the
tool long enough for the reps. to change head 3's faulty rotation
resolver cable and finish the monthly PM. Tool is back to production.
|
19.63 | Carrier #5 not rotating | FABSIX::R_GEE | | Sun Jan 21 1996 07:06 | 16 |
| Came in tonight and found that carrier #5 was not rotating
during wet mode.
Found the power light on the Infranor for carrier #5 rotation
wasn't lit. Checked all the connections, they looked o.k.
Replaced the Infranor with one from the cabinet. When we flipped
the breaker back on the Infranor smoked and the breaker tripper
off. Now the breaker (CB6) won't reset and we believe that power
supply (PS4) may be blown.
The original Infranor is back in the system, not connected and
the power is off. The carriers are in the load cups wet.
B shift
|
19.64 | blown PS 4 and Hi I infranor | FABSIX::B_FINN | | Sun Jan 21 1996 19:41 | 26 |
|
Problem: Breaker 6 trips when system is powered on. Removed connector
xa3 from power supply to isolate 310 volts supply from load. Breaker still
tripped immediately on power up. Removed connector xa1 from 310 power supply
and cb6 did not trip. Replaced power supply and installed a new high current
motor controller (Infranor, crt12) to replace the original motor controller
which had errored last night.
Please note: High Current Infranors are used for carrier rotation, and Multi
Head. Low current Infranors are used for carrier oscillation.
Tested fix by running 25 wafers, first five wafers had errors at flipper
gripper at prealign station, next 5 ran fine, then 3 errors, then rest of the
lot ran fine. Looks like gripper is intermittently sensing that the
gripper is closed. Maybe amplifier adjustment or sensor position is on
the edge.
All wafers ran fine at the multihead and polish table, no Infranor errors.
System still down for intermittent gripper errors at prealign.
310 volt power supply and low current Infranor tagged "bad". Original Infranor
tagged "needs verification"
|
19.65 | Heads out of alignment | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Jan 26 1996 06:28 | 10 |
|
Several times there were near crashes on head #5.
There was some loading problems with the flipper as well.
Checked out the loading and it was fine. I found #3, #4
and #5 heads out of alignment (#5 was out 1/2" again).
Brought all (5) heads in perfect. Will speak with Reps
about why they keep drifting out of alignment.
BJ
|
19.66 | lift cylinder 5 is the main culprit | SUBPAC::LANDRY | | Fri Jan 26 1996 18:24 | 43 |
| Busy day on this baby..........
The system was misloading wafersfrequently, mostly on head 5.
So, the system was run using the infamous "NO_TOUCH" program.
Right off the bat, head 5 misloaded due to the load cup not
quite making it to it's loading destination smoothly. It
raised up about 3/4ths the way, stopped there for a second or
two, then jerked up into load position. So, the lift cylinder
was wiped clean and lube'd with Braycote. The lift regulator
pressure was put back to 45 PSI (found it <40). The cup lift
movement improved to where we thought this might be acceptable.
Next, we tried again, and lo and behold, the Index table messed
up. Found that when the plexi doors were opened then closed,
the index table would jerk when the DCX boards/contactor kicked
in. When this happened, the home position for the table would
be askewed. It's possible this may have happened last night as
well, due to the heads being way out when checked by D shift.
Upon looking it this, we checked the following things:
- DCX board connections (good)
- Swapped DCX modules (no daughter boards available) (no change)
- Swapped Infranors with a known good one (no change)
_ Checked, cleaned and re-gapped the Index Table home sensor
(no change)
We then tried homing the table. It repeated proper home position
every time. We shouldn't have a problem with running in auto-cycle
as the table home sensor is checked prior to every load sequence.
We will investigate later on to why this is occuring.
So, next we re-visited the cup alignment. Sure enough, several
heads were out, ALL in rotation adjustment rather than head in/out
movement. So, we tweaked these heads in.
Upon re-running, load cup 5 began to fail again. Therefore, the
system will have this head dis-abled until a new cylinder arrives
(due in tomorrow at Packaging Plus, right down the hill in the
Shaws shopping center). The fifth carrier is up by the sink,
keeping wet.
many contributors (A & C shift, Tim, Earl, SFam)
|
19.67 | Lifter replaced. | FABSIX::B_WESTPHAL | | Sat Jan 27 1996 16:07 | 7 |
| CMP.B2 - Down for load cylinder #5 sticking. The system was returned to
production yesterday w/ carrier #5 not enabled due to the air cylinder
sticking. The part came in around noon and was installed, height was
checked, and forty wafers were ran to verify that it was working
properly. The pad was also changed and the tool was wiped down by
production due to a particle problem. The tool was returned for MQCs.
|
19.68 | #5 load cyl. replaced | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Jan 27 1996 16:35 | 13 |
|
this is for CMP.B2 not B1.........
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
-< CMP >-
=============================================================================
Note 19.68 CMP.B2 EE/PE PASSDOWNS 68 of 68
SUBPAC::CMP 2 lines 27-JAN-1996 15:01
-< replace cyl. >-
--------------------------------------------------------------------------------
CMP.B1 - Replaced the cyl. for load cup5 ran wafers, at this time the
cup is coming up and loading wafers. SF
|
19.69 | Alignments / load flipper / rotation stopped | FABSIX::R_GEE | | Sun Jan 28 1996 08:20 | 25 |
| A wafer slid out of carrier #5 at the begining of our shift,
it didn't break. We check:
Carrier to load cup height alignments, o.k.
hub to load cup alignments, o.k.
table repeatability, o.k.
load flipper to load cup alignments, adjusted 25 steps back.
Ran 75 wafers when an error began to appear "wafer not present
at load flipper", it was there but the flag didn't see it.
Adjusted the flag to the right slightly.
Ran 25 wafer when another error the table and all the carriers
stopped. No error on the infranors but dcx board #2, led #3 and
5# lit.
Powered down the system and restarted. Didn't have a
chance to test after that.
B shift
|
19.70 | inverter error | STRATA::FINN | | Sun Jan 28 1996 19:30 | 14 |
|
System ran fine today until after glitch. Around 5pm Jim complained
that his touch up recipe #1 caused the polish table to stop turning when the
Multi head moved out to polish, and carrier rotation started. Carriers would
descend then with the carriers on the pad the table would try to start,
rotation, but error out for "inverter error".
Program looks fine, tried "notouch" program and table works fine, we
are about to try to edit this program to a touch down program, excuse the pun
cowboy fans.
No gripper problems today.
|
19.71 | Carrier #2 fly out | FABSIX::R_GEE | | Tue Jan 30 1996 07:14 | 9 |
| A wafer slid out of carrier #2, no wafers broke. Production
wanted to finish processing the lot running so we removed
the carrier and disabled head #2. The carrier is in water
on the table next to the sink, it looks o.k.
B shift
|
19.72 | index table problem. | SUBPAC::CMP | | Tue Jan 30 1996 16:42 | 5 |
| CMP B2 16:40
Due to a lurching of the index table which seemed to cause
a possible load error on carrier 2,investigating to find out why.
Checked voltage on infranor, swapped infranor power supplies, reseated
cables on motor. we have a motor ordered which will arrive in the A.M.
|
19.73 | index table drift | FABSIX::B_FINN | | Tue Jan 30 1996 19:24 | 57 |
|
Carrier 2 lost wafer at polish on b Shift. Checked carrier two for defects
exposure ect. aok.
* reinstalled carrier on b2
* checked flipper load to all cups aok
* manually loaded 15 wafers found no problem
* loaded 15 wafers to carriers, no problem,
* ran 3 run of no touch down program, no problem,
* Polished 15 wafer in program #42, no problem
Returned to manufacturing and system tossed a wafer from head two on
touchdown first run. Carriers one and two are history. 3 good carriers
are in water, Tim asked if you would change the water to prevent
bacteria growth in the during the night.
* Installed cup to carrier alignment fixtures and "homed table" and
performed alignment. All cups were out in the index table ccw
direction with heads 4 and 5 needing compound adjustments ( carriers
in). When I had it aligned, tested by moving multi head out and
exercising the carrier arms. Returned to load and all alignments held.
Reps joined the party at this point.
Index table was homed and position rechecked. The Bloody Index Table
had drifted.
Tim joins the fun and we verify that the calibration data is actually
being stored in the DOS files.
Reps check with factory who confirm a know software bug " index
table can only be safely calibrated in CW direction". If CCW
direction is used it subtracts the number from zero changing your
home position instead of adding or subtracting from current position.
We over compensated in Index table CCW direction, Homed table, then
brought it back into alignment with the CW direction. Our success
assured, we homed the table and found THE @#*%# table drifted again.
Table is still jumping when DCX is intit /interlock is opened.
Last Friday the infranor and DCX boards were swapped without success.
So Tim has requested that the reps fedx a new index motor which will
be here tomorrow.
ETCH tech have lent us a oscilloscope. We measured the infranor IDC
pin 1 and 2 per Darren. We found 2 volts of noise/sine wave, this
does not appear on other infranors on B2, nor is it on the infranor
for B1 index table. We unhooked the 25 pin connector from the dcx and
the signal cleared up. Dan and rich are now tracing back to DCX board
swapping cables trying to get the bump to move to motor #10.
I cleaned up the broken wafer just before I changed aligned the table
but i have not got to the pad change yet, torn off it by CMP.A1.
A shift
|
19.74 | the case of the nervous index table. | SUBPAC::CMP | | Wed Jan 31 1996 13:16 | 5 |
| CMP.B2 SpeedFam
installed new motor and we found that it indeed cure the shakes
will continue to test and recalibrate.
|
19.75 | Index Table still exists | SUBPAC::LANDRY | | Wed Jan 31 1996 19:32 | 16 |
| Come to find out that the new index table motor installed this morning
has NOT solved the jerking problem. It most definately still exists.
Many things were tried today to isolate the problem. New motor, new
cable, different DCX boards, different DCX slots, DCX power supply
swapped, etc. The problem "appears" to be within the VME interface.
We did some ground checks in several places. Those are at the
Infranors, the incoming power ground wire, the transformer in the
pipe gallery. Also checked the nuetrals and hot connections in the
transformer. All connections were tight, except for a couple of the
Infranors.
See Tim's passdown for any additional instructions on the tool.
A Shift
|
19.76 | noise, noise, everywhere??? | SUBPAC::CMP | | Wed Jan 31 1996 19:44 | 5 |
| CMP.B2 SpeedFam
well we thought we had it,but the problem continued to pound us.
we checked noise levels, grounding and even tried another DCX board.
At this time we are going to order a power supply, and the 2 cables
for the motor.
|
19.77 | Are we there yet | SUBPAC::CMP | | Thu Feb 01 1996 19:22 | 8 |
| CMP.B2 SpeedFam
well were almost there with the tool. Things done today were,
check circuts for the noise, and looked at a voltage difference
at MC6 conntactor. Found that after tracing to infranors that the
one that was swapped last week for the problem was also bad in the same
way. the tool was calibrated with some fancy cuss words and a little
excitement. At this time the tool will be able to run with heads 1-3
until the infranors come in tomorrow.
|
19.78 | INDEX TABLE JUMPING. | FABSIX::B_WESTPHAL | | Thu Feb 01 1996 19:39 | 6 |
| CMP.B2 - Down for the index table jumping after an interlock was actuated. The
problem seems to have been traced to two bad infanors that will not
maintain auxillary power. Four new infanors are on order. The
alignments for the load cups have been completed but the tool still
needs to be cleaned up, new pad and, the three still good carriers
need to be installed.
|
19.79 | B2 almost finished. | SUBPAC::CMP | | Fri Feb 02 1996 14:27 | 7 |
| CMP.B2 SpeedFam
Almost there with the tool! last night the scrubber motor was
taken because B1's died, so we installed the upgrade while the tool
was down and we had a Rep. from infranor arrive to give some info about
the operation of the little beasts. Need to adjust the calibration on
the scrubber and run wafers.
|
19.80 | INFRANORS/SCRUBBER UPGRADE | FABSIX::B_WESTPHAL | | Fri Feb 02 1996 19:40 | 18 |
| CMP.B2 - Down for the index table jumping when an interlock was actuated
(waiting for parts). The infranors came in around 11:00am along w/
a rep. from infranor. The infranors for carrier rot #4 & #5 were
swapped out with the new parts. The problem with the table jumping
has went away. The infranors Irms values going to the motors for all
the infranors on both tools were adjusted under the recommendation of
the Rep. were as follows. All of the carrier rotation Motor controll-
ers were tweaked to 9 volts (checked at test point 6 on the controll-
er). All of the rest of the controllers are set to 10v. The only
exception is carrier rot #1 on CMP.B1 this one has been set to 13v
because of some carriers not rotating errors. If these errors occur
again the infranor should be adjusted to its max and if the problem
persist Tim Dyer and Daniel (Speedfam) should be contacted and the
Infranor swapped w/ the one in Speedfam's office. The scubber was
also upgraded and calibrated. 50 wafers still need to be run to verify
proper handling and scrub operation. Then Tim wants the pad changed
before the tool is returned to production.
|
19.81 | Scrubber RPM | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Feb 03 1996 04:33 | 9 |
|
The tool had run 25 wafers without problems with ANYTHING.
While running the last 25 wafers, I noticed the RPM's on the
scrubber was way too fast. I checked the RMP setting on both
tools and it was the same, 300. I changed B2 to 100. The
pad was removed and break-in's were run. MQC's passed fine.
all set...... BJ
|
19.83 | Scrubber problem with B2 | FABSIX::R_POIRIER | | Sun Feb 04 1996 07:10 | 16 |
|
We had a problem last nite with the scrubber on B2. The
brushes would not spin. We checked all the connections
and the input voltages. We found that the motor was not
getting the 50 Volts. We tried checking continunity from
the motor to motor controller #18. No continunity. We tried
bypassing the cable by running a set of jumpers from the
controller to the motor and we were able to get the motor
to work. The cable must have a break in the wire and needs
to be replaced.
Since we do have the Ontrak scrubbers we didn't feel it
necessary to "Down" CMP.B2
B Shift
|
19.84 | wire problem X 2 = | SUBPAC::LANDRY | | Sun Feb 04 1996 19:21 | 32 |
| A couple of problems with the system today...............
- The scrubber brush rotation problem is indeed due to a broken
wire within the liquid-tight flex tube, just as B shift indicated.
This problem has been a major pain in the buttock on B1 in the
past. This is the first occurance of this problem on B2. As
long as the design/wire used remains as is, we can expect this to
happen several times a year.
The reps had just recently installed an upgrade to this scrubber,
which I believe happen on Friday. New wiring was suppose to have
been part of this upgrade. Notes entered on the scrubber upgrade
do not reflect if this was actually performed or not. So, the
status remains the same, with the system running minus the use
of the scrubber.
- Problem # 2 for the day occured when the system errored for "Main
Vacuum Low" alarm. Found the pump to have shut off. Checked the
breaker, low and behold, it had NOT tripped off. So, began to
inspect the pump/wiring for obvious problems. Bingo! Found a
wire, and it's connector, burnt to a crisp. Cut back all the
burnt portions. Re-attached the wire using a butt connector.
The pump ran well after that.
- Upon repowering the system, twice in a row we got an Inverter error
during initialization. Both times, it was CB #10 (15 VDC power
supply breaker) that had kicked off. Reset the breaker one more
time, and the system run fine. Not sure as to why this kicked off
twice.
A shift
|
19.85 | scrub station | FABSIX::R_GEE | | Mon Feb 05 1996 07:07 | 9 |
| * The scrubber wouldn't rotate.
Replaced the cable and installed a molex connector.
Re-routed the water line so that there is more room
for the cable.
B shift
|
19.86 | The scrubber | SUBPAC::CMP | | Mon Feb 05 1996 10:58 | 8 |
| CMP.B2 SpeedFam
Just want to let everyone know about the scrubber fix done on
Friday, we replaced the motor to the unit because it was used to get B1
up and running, also the cable was replaced on Sunday due to the old
one breaking, THIS IS ONLY A PARTIAL UPGRADE.....We will be completing
the upgrade this week as work permits. For those interested, the rest
of the upgrade {should} take about 3 hours.
|
19.87 | scrubber upgrade | FABSIX::B_WESTPHAL | | Wed Feb 07 1996 19:45 | 2 |
| CMP.B2 - Down for srubber upgrade/pm. The scubber upgrade is complete but pm
needs to be finished.
|
19.88 | Scrubber upgrade, vacuum capacitor, particles | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Feb 08 1996 06:22 | 19 |
|
The tool needed 25 more wafers run for the scrubber
upgrade. The weekly PM was done already so that didn't
need to get done. Everything ran fine except the tool
failed MQC's for particles. We cleaned the entire polish
table area and MQC's came in.
While cleaning tool, we noticed a loud vibration coming
from the vacuum area. It ended up being the pump capacitor.
Actually, the bracket that holds it down snaped off. The
cap didn't need replacing so I'll check with the Speedy guys
and see if they can get one shipped out to us. The cover to
the cap is on the tool box. Please don't toss it....
D shift
|
19.89 | scrubber upgrade | SUBPAC::CMP | | Thu Feb 08 1996 09:05 | 6 |
| CMP.B2 SpeedFam
Just want to let everyone know that the scrubber upgrade is
completed. For those service related type guys, the motor has a new gear
ratio (11.5-1) and the IN and OUT sensors are now located on the side
of the tub, and are connected to BB11 pins 18 & 19 (DG in 88 & 89) and
power is located on the lower part of the BB strip.
|
19.90 | PARTICLES/SCRATCHES | FABSIX::B_WESTPHAL | | Thu Feb 08 1996 19:15 | 7 |
| CMP.B1 - Down for high particles/scratches. The pad was changed and the polish
area was cleaned with a brush. The index table was drained and wiped
out. The PVA brushes on the scubber were swapped out with new ones
from stock. The conditioning arm was disassembled and cleaned and
operations rechecked the MQCs. The particles passed and there were no
scratches but the uniformity failed on head #3 at 10.8. Production is
going to run some more wafers and it should come in.
|
19.91 | Transfer Paddle Interference problem | ASDG::POIRIER | | Sat Feb 10 1996 19:52 | 8 |
|
* Transfer Paddle interference:
----------------------------
The Transfer Paddle is interfering with the 2 teflon bars that were
added to the scrub station as part of the recent scrubber upgrade.
The paddle contacts the bars on its up stroke. This causes the wafer
to jump-up approx 1 ". It appears that a minor adjustment will
clear-up this problem.
|
19.92 | scrubber upgrade modified | STRATA::RIDLON | Eliminate the obvious | Sun Feb 11 1996 01:10 | 16 |
| The thumb screws for the wafer support stantions in the scrub station
were loose. These were tightened and clearance between the teflon bars
and the paddle was checked and found to be ok. However when the tool
was checked a couple of hours later the left teflon bar and two wafer
support rings were in the base of the scrub station. They were put back
on without interrupting the unload cycle. When the paddle lifted up the
next wafer it indeed appeared the paddle was hitting the left teflon
bar. When ops was finished with tool the was checked. With out a wafer
in the scrub station the paddle was raised and had plenty of clearance
on both sides. A wafer was installed and ran thru a scrub manually.
When the paddle came up the left teflon bar lifted. There is no
clearance between the wafer and the teflon bars so the action of
scrubbing them creates a vacuum between them. The teflon bars were
flipped over as a temporary fix for now.
B shift
|
19.93 | wafer squirted out....no damage | SUBPAC::LANDRY | | Sun Feb 11 1996 19:33 | 12 |
|
This morning.............
A wafer slipped out of head # 5 during segment # 3. Wafer
detection did it's job and shut down quickly. No wafer breakage,
nor pad or carrier damage was exhibited. The Opt's folks rebooted
the system and unloaded the remaining wafers manually. They just
resumed running the system, which ran fine for the remainder of the
shift.
A shift
|
19.94 | Slurry flow stopped | FABSIX::R_GEE | | Mon Feb 12 1996 07:16 | 19 |
| * While the system was running 3 banging noises came from
polish table, it was the drive belt slipping. By the
time we got to the system the banging had stopped and the
system seemed to be running fine. On the next run
the multi-head went to the polish table, slurry began
to flow. When the carriers began to drop (segment #3)
the slurry flow stopped, the system was EMO'ed.
The program that was running was I_O_ILDA #33. We checked the
program, the slurry flow was correct. We checked the slurry
using the manual polish program, it looked good, no leaks.
Ran 25 dummies using the I_O_ILDA program and everything
looked o.k. We believe the slurry flow stopped due to
a software problem and was cured by the re-boot.
B shift
|
19.95 | scrubber info | SUBPAC::CMP | | Mon Feb 12 1996 08:44 | 8 |
| CMP B2 SpeedFam
I saw the note on the teflon bar lifting on the scrubber. The
drawing shows the bar to be installed the way it was operating but
if the tool runs better this way then we will let it run and I will
ask about whether the drawing could be incorrect. The bar is there
to keep the wafer from falling into the lower part of the tub if it
were to slip off the flipper.
|
19.96 | ARm cals | LUDWIG::RIDLON | Eliminate the obvious | Tue Feb 13 1996 02:31 | 30 |
| Entered for A shift.
CMP.B2
==========
The system is currently DOWN as scheduled equipment, in the
competent hands of Process/Equipment Engineering, for the
purpose of qualifying the second polish table, the blocky
diamond, etc. The system should remain down until released
by either Frank or Tim.
Some things that took place today.........
* Arm Oscillation Calibration was performed. This procedure
is to set up the arms such that they oscillate out to the
same physical position on the polish table. The arms were
just slightly out, but the proper adjustment was made to
"fine tune" the arms.
* The Monthly PM was performed this afternoon. No problems
were found.
* The pad and all carriers were changed out at this time.
This included a major clean at all the stations to remove
a plethora of slurry buildup.
* I don't have the specifics at this time, but Frank was
making some changes to recipes.
A shift and Guest
|
19.97 | scrubber upgrades | SUBPAC::CMP | | Wed Feb 14 1996 08:16 | 3 |
| CMP.B2 SpeedFam
We finished the scrubber upgrades, so now B1 and B2 are completed.
|
19.98 | 2nd Table Process released to Manufacuring | ASDG::POIRIER | | Thu Feb 15 1996 19:35 | 6 |
|
* 2nd table process released:
---------------------------
Frank Krupa has completed characterization of the 2nd Table process.
The tool has been returned to Manufacturing. See Frank's message
for details.
|
19.99 | Wafer slide out | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Feb 16 1996 07:29 | 7 |
|
A wafer slid out of head #5 for no known reason. The handling
looked ok on the next five wafers which were run as MQC wafers.
If another slide out happens, further alignments will have to
take place.
D shift
|
19.100 | Lost wafers/ scratches | FABSIX::B_WESTPHAL | | Fri Feb 16 1996 19:39 | 12 |
|
CMP.B2 - The tool lost a wafer from head #5 this morning. It broke, so the pad
was changed, alignments checked. All seemed okay except for the inner
portion of the load cup was sticking alittle. The same was
disasembled and sanded. The tool slipped another wafer later in the
afternoon this time it lost at touch down so this time it seemed to
be loading problem. Load cup #1 & #5 were swapped to see if the
loading problem follows the cup. This is still to be determined. The
tool also had a problem w/ scatches on wafers in head #5 earlier today
and this problem followed the load cup to position #1. The load cup
was inspected and there are a couple of embeded particle in the cup.
Ran out of time.
|
19.101 | scratches and no slurry flow | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Feb 17 1996 07:12 | 14 |
|
Several deep pieces of wafers were into load cup #1 and #5.
Everything was cleaned up and a MQC particle check was done.
The tool passed uniformity and particles with no scratches.
The tool would of been up by 9pm if there were MQC wafers
available.
During polish, the slurry flow went to nothing.
We pulled the slurry filter and found a lot of crystalized
slurry and but not as bad as last month. Installed new
filter and checked slurry flow. all set.....
D shift
|
19.102 | Wafer Breakage / Speedfam Rep on the way. | ASDG::POIRIER | | Sat Feb 17 1996 19:31 | 21 |
|
* Wafer Breakage:
--------------
The system flew a wafer out of carrier #5. We found broken wafers
on both the primary & secondary tables. The system was stopped in
segment #7 (lift off). Gwen noticed 2 error messages after the crash,
"Invertor Communication" & "Main Vacuum Low".
We feel that the first wafer broke during 2nd table clean. The system
must have unloaded 4 wafers and loaded five more. It then moved to
the polish table where a wafer flew out of carrier #5 & broke.
In the end, 2 wafers were smashed at 8 were severly scratched.
Nine product wafers & 1 monitor were scrapped.
B shift, please finish the clean-up activity and wait for the
Speedfam rep (Tony) to arrive with further instuctions. We expect
him to arrive at 8:00 p.m. tonight.
C Shift.
|
19.103 | Secondary table by-passed / Left wafer on pad | FABSIX::R_GEE | | Sun Feb 18 1996 06:48 | 18 |
| At the end of "C" shift a wafer broke on the secondary
table. The system was cleaned, carriers re-installed and
a new polish pad installed. Tony was here and checked
the system out, nothing looked wrong. Process Engineering
changed the recipes so the secondary table will be by-passed
until monday morning.
This morning after running 85 breakin wafers the system left
an MQC wafer on the pad after polish from head #3. We believe
that lifting the carriers from the center of the pad is creating
a vacuum between the wafer and the pad causing a wafer to stick
to the pad. Need to talk to Engineering about moving the lift-off
position out or in. The system is up but needs MQC's.
B shift
|
19.104 | multi problems | FABSIX::B_FINN | | Sun Feb 18 1996 19:24 | 34 |
|
Head 2 threw a wafer during touch down on the pad, prior to
release of vacuum. Wafer broke and took out one carrier.
We took the system for checkout. Installed alignment posts
and verified carrier to cup alignment, arm 4 needed to come out by
320 motor counts. Head 5 was moved out 64 counts.
We then verified cup height alignment on all five heads at
there respective carriers. They were fine
.
Next we verified flipper load calibration. Slight adjustments
were made at the flipper to load cup cal.
15 volt power supply tripped twice while rebooting system.
We removed all tape from all interlocks. The process door
switch was bent such that when the door was closed it would not make
the sensor. Sensor block was removed and repaired.
Slurry dispersion tube was clogged. Removed said tube and
cleaned clogged.
Jesus ran first five breakin wafers fine, on second load system
miss load wafer to head 4, did not pass vacuum check at pad.
Head 4 seems to be coming later than the rest, and then
jerking up to the carrier.
Scratches are present on cup four hole in index table. We are
attempting to remove then before adjusting pressure at cup up
# 4 solenoid.
A shift
|
19.105 | Load flipper alignment / unload cup #1 problem | FABSIX::R_GEE | | Mon Feb 19 1996 07:03 | 27 |
|
The system was down do to a wafer being mis-aligned onto head
#4 at the end of "A" shift.
We checked:
- cup #4 up/down motion, it looked good.
- spindle to cup #4 alignment, it looked good.
- had to adjust the load flipper to load cup 15 counts
because the wafers were floating out of the cups.
Replaced the pad do to a bubble under the pad.
Ran 110 verification wafers then MQC's. On the first MQC
the system passed uniformity but failed for a scratch (?) 5.3 mm
on head #1, the second MQC passed.
**** We noticed when the unload flipper goes to pickup the wafer out
of unload cup #1 it is to low, it still pickup the wafer, barely.
We found that unload cup #1 has a load cup piston plate.
It's about an 1/8" bigger causing the cup to raise to high.
This needs to be replaced as soon as possible, we have
none in stock.
B shift
|
19.106 | vacuum switches | FABSIX::B_FINN | | Mon Feb 19 1996 19:32 | 49 |
|
Head 5 threw a wafer. It did not break. Wafer loss detector
picked up the loss. But vacuum sensor did not detect.
Tim requested that the vacuum sensor trip point be turned up
to 21" hg. We tried to do this but there is a difference
between the turn on threshold value vs turn off threshold
value. We needed to spit the difference to prevent false
alarms.
The sensors are set to the following values
Head detects loss detects vac
1 18("hg) 22("hg)
2 18("hg) 22("hg)
3 18.5("hg) 22("hg)
4 18("hg) 22("hg)
5 18("hg) 22("hg)
Cup 5 outer ring is not retracting after being depressed when
in it index table hole. Movement is free is its removed and hand
held. Feeling is that the outer shaft of the cup is binding. Load cup
4 also exhibits this to a lesser degree. Load cups 4 and 5 were
swapped. Load cup 4 works fine in index table hole five position. Load
cup 5 still binds in index table hole 4. Were added to the base of
all load cups, during which time it was observed that the screws/springs
on cups 3 + 5 were backed out two far preventing compression of the
springs. This was fixed. cup 5 outer ring still stuck. Dual springs
were placed on all 3 screws for cup 5. It still stuck. It was
disassembled, but movement was fine. The cup was mic'ed and is in
spec and reasonably true. Tony will check it in other 3 remaining
holes and possibly sand it down some.
System still has one 21 hole carrier. it needs to be swapped with the
29 hole carrier which is labeled and on the ss table near the
changeroom door.
A shift
|
19.107 | Head %3 off line / #1 & #5 unload | FABSIX::R_GEE | | Tue Feb 20 1996 07:15 | 23 |
| - Head #5 was taken off line by field service do to load
problems. Parts are on order to repair the cups. A warning
has been put into RTC.
- Swapped the piston plates for unload cup #1 and #5. #5 is
off line and #1 had the wrong piston plate, it was a load
cup piston plate.
** Unload #5 now has the wrong piston plate, it will need
to be replaced when the parts arrive.
- On the 5th run after completing the polish cycle a wafer became
mis-aligned on carrier #4 giving use a vacuum error on all 4 heads.
We believe it was a pad to wafer suction problem. Re-aligned
the wafer and restarted the system.
- We ran 120 wafers after that with no errors. The system past MQC's
at 1:30am, and has been running product since then.
B shift
|
19.108 | passdown | FABSIX::B_FINN | | Tue Feb 20 1996 19:06 | 25 |
|
Load cup five was disassembled and sent out for truing.
Received shaft five back and reinstalled it with new inner shaft.
Cup five has been installed in the index table between lots, but is
still disabled. System is running 4 heads. Cup 5 needs to be
qualified at next MQC.
All springs have been replaced on bottom of cups.
Unload cup one has had its peddle base replaced (mate for cylinder
pedestal).
All load and unload pedestals are now the correct height.
All LOAD cup inner shafts have been replaced with new ones.
The moisture sensor at the Fluids cabinet has been relocated to allow
for the new slurry bulkhead fittings.
SEE Tim's PD. If System loses a wafer he wants the system to
remain down, five new load cups are due in house tomorrow
A shift
|
19.109 | LOAD FLIPPER PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Feb 21 1996 07:30 | 26 |
|
The tool was running a hot lot and an error for the
UNLOAD flipper not picking up the wafer from the load
cup came up. I verified the wafer was securely in the flipper
fingers. I hit a button to continue as it asked and the
paddle came up from the scrubber with a wafer and stopped.
The UNLOAD flipper came in to do the slam dunk and did not
stop at the "wafer drain" step. It proceeded to head for
the paddle which was still up. The tool was EMO'ed. Circuit
breaker #10 tripped three times in a row.
Later, the tool was running and errored for "LOAD flipper
did not pick-up wafer at align." The load flipper would
not home. After an hour of troubleshooting, I gave Tony
a beep to help me remove the flipper box. We traced the
wiring back assuming the problem was the motor itself. While
Tony was removing the motor, it wasn't connected to the
drive shaft. The collar itself was the problem. I don't
want to get into the details on what exactly was found but
only to say that the screw in the collar was broke and could
not be removed. Tony is going to rush the parts in this
morning. He's getting a whole new motor and gear box sent
in.
BJ
|
19.110 | flipper gripper for the gipper | SUBPAC::CMP | | Wed Feb 21 1996 13:05 | 4 |
| CMP-B2 SPEEDFAM
We took the gear that was broken to the machine shop and had the
old screw shaft removed and a new screw installed, now in the process
of putting flipper unit back together for your processing enjoyment.
|
19.111 | Index Table Load CUps / Flipper Box | ASDG::POIRIER | | Wed Feb 21 1996 19:27 | 39 |
|
Index Table & Flipper Box:
--------------------------
a) 10 new Load Cups arrived and were installed into the Index
Table.
b) 10 new Dowel Pins were installed into the Load Cup Bushings.
(New locking type with raised sides)
c) All 10 Load Cup Bushings were honed to proper size before
installing the new Load Cups. (We now have a new GO/NO GO
gage for sizing the Bushings from Speedfam.
d) The FLipper Box Motor Coller was repaired at Bryco Machine
this morning. The broken screw was drilled-out the threads
were re-tapped. (While you wait service..wow!)
f) The Speedfam reps had prolems with the proper positioning of
the spline on the Flipper Box Motor Coller. As of 7:15 the
problem has been resolved. They have installed the Flipper
Box back into the tool and are now aligning the Flippers to
the load/unload cups.
g) Brad has installed a new N2 Gun on the tool.
From Tim Dyer's Passdown. D Shift, here's the progress thus
------------------------ far:
1) Install new cup locking pins DONE
2) Set ALL cup to carrier head heights NOT DONE
3) Align ALL heads to CUPS with fixtures NOT DONE
4) Align gripper to load/unload cups IN PROGRESS
5) Cycle 50 wafers through NO_TOUCH NOT DONE
6) Cycle 100 wafers through QUIK_TCH NOT DONE
|
19.112 | FLIPPERS ALIGNED, BACK TO SCRATCHES | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Feb 22 1996 07:03 | 13 |
|
We went through the fine adjustments on the load and
unload flipper. We had a rather tough time with the unload
flipper but it finally came. After 150 wafer verify, a new
pad and without any problems, it was turned over to OPS for
production. MQC's failed bad for scratches and NO_TOUCH was
run and ruled everything out except the pad. More wafers are
being run now for testing.
BJ
|
19.113 | Rotating error, scratches | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Feb 24 1996 05:15 | 12 |
|
The tool had problems with scratches again. This time they
were straight and 1" long. It was ok during NO_TOUCH but showed
up during touchdown. After more MQC's were run, the scratches
went away???????
The tool errored for Carrier #1 not rotating several times
tonight.
BJ
|
19.114 | Conditioner Sensor Adj. / Scrubber Down Sensor adj. | ASDG::POIRIER | | Sat Feb 24 1996 19:08 | 14 |
|
CONDITIONER SENSOR PROBLEM:
--------------------------
The system encountered many "Invertor Timeout" errors This Morning.
The Speedfam rep (Tony) was paged around 10:00 a.m. He found the
Conditioner "Wheel to Table" sensor out of adjustment. The Invertor
Timeout errors stopped after the sensor was re-adjusted.
SCRUBBER TIMEOUT ERRORS:
-----------------------
The scrubber was failing to make the "Scrubber Down" sensor Today. This
caused numerous "Scrubber Timeout " errors. Tony adjusted the Scrubber
Down Sensor Sensitivity pot and the problem ceased.
|
19.115 | infranor/pad conditioner replaced | FABSIX::B_FINN | | Sun Feb 25 1996 19:40 | 47 |
|
Wafer breakage on secondary pad. Wafer Broke on head 5 which
took out head 4 in turn. Presume wafer became slightly ajar during
liftoff and broke on touch down at secondary pad. Head 5 shows strong
circular and straight patterned in carrier pad. Head 4 shows only
circular damage. Most of the fragments were found on and around
secondary and index table. System continued until it tried to unload
a broken wafer from cup four. Some collateral damage occurred to
wafers which were next to be polished (prior to unload error).
Fragments of silicon made their way to the polish table and caused
scratches on the five wafers at polish table.
Cleaned up silicon debris and scratched wafers from polish table.
Removed/replaced both secondary and polish pads. Replaced carrier 5 and 4.
removed and replaced ring from carrier 3.
Removed clog from both drains at secondary table. Had inverter error
twice on power up, found 15v ps circuit breaker tripped. Tried to
power up again but errored for " waiting for pad conditioner to home"
error. Trouble shot found inhibit sw for cond. home made, (open), and no
power or errors at Pad Cond infranor but power was on to its power
supply and other infranor on its loop.. Tried to move pad conditioner off
switch but could not. Reset breaker for infranor. Jumpered inhibit
switch and used DCX to manually move it off switch and tested home position
sensor while in DCX mode aok. Tested inhibit switches with multi meter aok.
Had numerous errors on power up. Including inverter errors and head
five not homing.
Found pad condition infranor would not boot up on initial power up but
would work for one cycle of conditioner if circuit breaker was
manually tripped and reset.
Replaced Infranor. System Ran fine.
Broke pad in.
On first run with new carrier we received error of "One or more wafer not on
carrier" at polish pad vacuum check.
Checked all carriers for misload found none but two very close to the
edge of the ring. Checked vacuum gauge found it at 22 which is
threshold value for vacuum sense since adjustments last week. Acknowledged
error and it ran fine. Next three runs ran fine. System is running
breakin wafers.
A shift
|
19.116 | water leak / scrub station up/down travel | FABSIX::R_GEE | | Mon Feb 26 1996 06:41 | 20 |
| The system was down due to a crash on "A" shift. Since
there was no process person to run the Speedfam's we
decided to do the weekly p.m.
- Repaired a water leak coming from under load cup #4's
air cylinder.
- Repaired a water leak under the elevators, found
2 fittings with the ferrules on backwards.
Started running verification wafers and had an error " scrubber
not arriving down". Found the teflon sleeve that connects the
brush spin shaft to the upper brush station stripped. Found
a replacement in the speedfam cabinet. We need to stock this
idem. Also, when you put the top brush station on be sure to
align the sleeve to the square shaft.
B shift
|
19.117 | Scrub station / load flipper / broken wafer | FABSIX::R_GEE | | Tue Feb 27 1996 07:02 | 30 |
| * The scrub station was timing out for not arriving home.
Lubricated the rails and turned up the air pressure
slightly.
* The scrub station wasn't traveling up properly.
Shimmed the upper scrub station Teflon collar and the
lower scrub station drive shaft with Teflon washers.
* At 6:ooam this morning on the 2nd run of touchups a wafer slid
out from under carrier # 4. The wafer slid out at touchdown to
the primary pad, wafer detection saved the carriers. We checked
all the alignments for carrier #4, they all looked good. We checked
the flipper to load cup alignment and it was off. The wafer was
being positioned to far forward resting ontop of the loadcup rim.
We adjusted the vertical position of the hole flipper assembly
by using the 2 top sets of alignment screws. The alignment is
better but we called like field service in to fine tune the
adjustments. Field service was called and should be in between
7:00 and 7:30.
The pad and carrier #4 have been changed.
B shift
|
19.118 | passdown | STRATA::FINN | | Wed Feb 28 1996 19:44 | 7 |
|
Vacuum error on first run of MQC wafers. Found no misplaced wafers,
vacuum was at 26" hg. UnLoaded from manual screen and restarted wafers
loaded and ran fine, Possibly semi dry films. System is down for pad
change. Poor uniformity
A shift
|
19.119 | MQC and Uniformity issues | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Feb 29 1996 08:59 | 10 |
|
Came in having to replace the pad and carriers #1, #2 and #3
for MQC's/wafer count related problems. Between B2 and B1, we
pulled (4) built SpeedFam carriers off the shelf that gave us
18% uniformity. They are on the cart in the maint. chase with
two crashed carriers from B1. After two pad changes, it failed
again for MQC's at 8am. Earl is working at it now.
BJ
|
19.120 | wafers staying on table at lift-off | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Mar 09 1996 04:53 | 10 |
|
Tool was leaving wafers on the polish table after lift
off and was causing vacuum low alarms. This was an issue
before and was corrected in the segments. I called everyone
in process that work with the tool and got no call backs.
We gave a pad change a shot and it seemed to have worked.
the tool passed MQC's and is running product without any
problems.
BJ
|
19.121 | Load flipper | FABSIX::R_GEE | | Sun Mar 10 1996 06:46 | 5 |
| The load flipper wouldn't move. Had to reset the computer.
B Shift
|
19.122 | TABLE STALL OUT, SLURRY FILTER CLOGGED | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Mar 14 1996 04:52 | 10 |
|
The slurry filter became clogged during the polish
cycle of a production run. The table dried up and caused
the table to over-current and stop the machine. The filter
was taken from A1 since there was none in stock and the flow
was checked. I had OPS run MQC's to verify there was no
damage to the pad or tool and everything checked out ok.
BJ
|
19.123 | Table indexed improperly | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Mar 14 1996 22:21 | 8 |
|
After the system was done carrier/wafer rinse on the first
(5) wafers in the run, the index table failed to go to the
proper location. It appears a process door was opened while
it was indexing and caused it to stop indexing.
D shift
|
19.124 | scratches | FABSIX::B_WESTPHAL | | Fri Mar 15 1996 19:43 | 7 |
|
CMP.B2 - The tool had a problem with scratches on wafers from head #1. The
load/unload cups were wiped out and the index table drained, the
carrier inspected and the final polish pad was changed out to adress
this problem. Still had what appeared to be scratches but process
found it to be dried slurry. Tool was allowed to run.
|
19.125 | Wafers being left on table at lift off | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Mar 16 1996 06:59 | 11 |
| i
System left a wafer on the table after lift-off (4) times
tonight. It only happened with PMD lots, not with metal lots.
With all of the problems we are having right now, we decided
to put the PMD's on the back burner and just run the metal
lots on it. A pad change fixed this problem last week.
BJ
|
19.126 | unload flipper | FABSIX::R_GEE | | Sun Mar 17 1996 06:15 | 7 |
| The unload flipper wasn't picking up the wafers out of
unload #1 & #3. Had to calibrate the unload flipper 10
steps forward.
B shift
|
19.127 | Scratches | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Mar 20 1996 03:30 | 7 |
|
System failed twice for scratches. Found the top
outside cleaning brush had fallen off its mount and
was inside of the tub. How long it was there is the
question.
BJ
|
19.128 | Unload Track #2 double slotting wafers. | ASDG::POIRIER | | Thu Mar 21 1996 20:32 | 9 |
|
* Unload Track #2:
---------------
The system has been double slotting wafers on track #2 all day.
Ops has been running the tool one cassette at a time to avoid
this problem. We ran like this due to CMP.B1 being down for the
cable upgrade. D Shift will work this problem as soon as B1
sucessfully completes MQC's. D shift suspects the water level in
tank may be incorrect.
|
19.129 | slurry ring drain tubes....... | SUBPAC::LANDRY | | Sat Mar 23 1996 14:09 | 23 |
| At the end of D shift, the Opt people put a warning on the system for
physical damage that occurred in the polish area. It appears that a
polypropylene flap (a flap that shields the MHTA rails from slurry)
was snapped off. It in turn wiped out four slurry drain tubes from
the top slurry drain ring. This, coupled with the three previously
broken tubes, left us with only two remaining tubes. We can not
afford to block off ~80% of the drain system. I was unable to find
the proper fittings in stock, therefore, I have "improvised" to the
best of my ability until we can procure the proper material. So, I
have inserted seven (7) 1/4" PVC nipples into the ring.
It's more than adequate, in fact, it may be an improvement in keeping
the polish area clean. The original tubes didn't extend low enough
into the bottom ring to properly contain the slurry waste. The newly
installed nipples DO extend into the ring, preventing the slurry to
be flung to the polish area walls. Oh boy, I can see my $500 cash
award for process improvement coming in the near future!
By the way, please do not use these PVC nipples as turning handles.
After all, the table does weight ~800 lbs and it IS only PVC fittings.
Doc
|
19.130 | quivver city | SUBPAC::LANDRY | | Wed Mar 27 1996 18:35 | 11 |
| Mid-afternoon today, the system majorly vibrated such that the EMO
was hit to prevent a crash. Upon inspection, the polish table was
found to lackenough slurry on it. There was some slurry in the middle
of the table, but not enough to make it out to the perimeter. We
can NOT confirm why this happened, but it's quite possible the quick
loss of slurry flow to the room happened within the time frame of
this occurance. After reboot, we manually flowed slurry to assure
a satisfactory amount was flowing, which there was. The system ran
like a champ the rest of the shift.
A Shift
|
19.131 | CMP.B2 shutdown for facilities tie-in completed | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Mar 28 1996 06:13 | 54 |
|
SHUTDOWN: 5:00 AM ON 3/28/96
STARTUP: 9:00 PM ON 2/28/96
CMP.B2 SHUTDOWN CHECKSHEET
--------------------------
Tech. Date
------- -------
1. Change RTC status to SCHEDULED DOWN. BJ 3/28/96
2. Remove all carriers. Any carrier that
has a wafer count higher than 500,
break the carrier down. If any carriers
have less than 500, place in DI water and
and mark a indicated. BJ 3/28/96
3. Remove all wafer receive cassetts. BJ 3/28/96
4. Remove and discard all polishing pads. BJ 3/28/96
5. Remove all PVA scrub brushes and submerge
in DI water. BJ 3/28/96
6. Perform a complete clean of each system
to remove dried slurry. Wipe down and dry
all wet surfaces. BJ 3/28/96
7. Remove and discard all slurry filters. BJ 3/28/96
8. Disconnect inlet and outlet lines on each
filter housing and remove housing. Connect
inlet and outlet lines together with a 3/8"
union. This will completely bypass the
filter housings. BJ 3/28/96
9. Close main incoming DI Water valve to each
system and open the DI Water return valve
(UDR). This will allow all incoming DI Water
to flow to UDR. BJ 3/28/96
10. Power down each system. BJ 3/28/96
11. Power down each Neslab TCU. BJ 3/28/96
|
19.132 | Receive elevator cross slotting | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Mar 29 1996 06:36 | 9 |
|
Put a warning on the tool. Receive elevator #2 is
still cross slotting. Product running on tool so I
couldn't get the tool. Ian from Speedfam is aware of
it.
BJ
|
19.133 | MOISTURE SENSOR ALARM | STRATA::WESTPHAL | | Sat Mar 30 1996 10:18 | 9 |
|
CMP.B2 - The tool was put up with an assist in reference to a moisture sensor
alarm. Found The drip pan under the unload station with water in it.
The leak was coming from where the fitting on the quick dump drain
screws into the main drain. This was caused by a tie wrap anchor
coming loose allowing the tubing to fall into the path of the
elevators' travel. The QD drain & cascade drain tubing were securely
tie wrapped out of the way and the pan and sensor were cleaned up.
|
19.134 | ELEVATOR 2 CROSS-SLOTTING PROBLEM | SUBPAC::CMP | | Sat Mar 30 1996 16:37 | 19 |
| TOOK A LOOK AT CROSS-SLOTTING PROBLEM ON ELEVATOR 2 AND FOUND 2
PROBLEMS.
1. C CLAMP THAT HOLDS THE FIBRE OPTIC CABLES IN PLACE, WHICH IS
LOCATED TO THE REAR OF THE FLOAT SWITCHES WAS CAUSING THE FLOAT TO
STICK DOWN. THIS MADE THE QUICK FILL SWITCH ON AND CHANGE THE WATER
LEVEL IN THE ELEVATOR, SO CAUSING CROSS SLOTTING.
I ROTATED THE CLAMP ON BOTH ELEVATORS 180 DEGREES SO THAT CABLES WERE
ROUTED FURTHUR AWAY FROM FLOAT.
2. THE CASSETTE PRESENT SWITCHES WERE LOOSE AND WERE ALLOWING THE
CASSETTES TO MOVE ON THE EVEVATOR DUE TO THEIR LOOSE FIT. THIS WAS
CAUSING MINOR ERRORS WITH SOFTWARE THINKING THE CASSETTE WASN'T PRESENT
BECAUSE THE SWITCHES WERE NOT ALWAYS BEING MADE.
READJUSTED THE SENSOR CLAMPS NEARER TO CASSETTE TO ENSURE POSITIVE
SWITCH ACTION AND MAKE CASSETTE A TIGHTER FIT ON BOTH ELEVATORS
|
19.135 | Robot to cassette #1 alignment | FABSIX::R_GEE | | Tue Apr 02 1996 07:48 | 8 |
| The robot missed a wafer from cassette #1. We found that
the wafers were being lifted too high when they exit the
cassette and dragging on the top of the slot. We lowered
the robot to cassette #1 slot #1 position slightly, all set.
B shift
|
19.136 | chiller remote completed | SUBPAC::CMP | | Tue Apr 02 1996 16:44 | 3 |
| Chiller now set and calibrated (24 deg C) for remote use.
Ian Lowbridge (Speedfam)
|
19.137 | Vacuum pump wiring | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Apr 05 1996 07:08 | 11 |
|
Re-routed vacuum solenoid wires out of the way
and neater looking.
Looked into the neutral wire on the vacuum capacitor
and found that the mount for the capacitor is once again
broke. The capacitor is up out of the way and secured.
The wire was ok. All of the wires have been secured as
well.
BJ
|
19.138 | Wafer crash, DI leak | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Apr 06 1996 06:07 | 19 |
|
Tool crashed a wafer from carrier #4 on the 2nd table. Little
pieces came out and trashed all five carriers which had to be
replaced. The hub to cup alignments were done as well as the
load positions. Everything looked ok. Everything was cleaned
up and new carriers and pads were installed. The tool ran fine
and passed MQC's.
The leak I discovered during the PM has been found and it
has been resealed. It was coming from the port between the two
drains. I siliconed both top and both. The tool will be DOWN
for the remainder of the shift and can be used at 10am after
Brad turns the DI feed to the scrub station brushes, to the
scrubber spray nozzles and re-installs the brushes which are
sitting it the elevator #1 tank.
D shift
|
19.139 | VACUUM PROBLEM? | LUDWIG::WESTPHAL | | Sat Apr 06 1996 20:38 | 14 |
|
CMP.B2 - Down for no wafers on carriers and DI leaking from the unload station.
Tool was down from this morning for DI leaking from the unload
station. The previous shift had siliconed the leak and asked for us
to verify the fix. Found it to still be leaking from the new silicone
weld. Stopped working on this issue because Production asked for the
tool. When the wafers were picked from the index station a "wafer
missing from carriers 1,2,3,4,5" error was recieved. Vacuum could be
felt on the wafer. Checked the vacuum and it was pulling 22 inches in
the pnuematics cabinet and 18 the multihead. The wafers seemed to be
loaded ok but the alignments were Checked anyway. The alignments
looked good but a small adjustment was made carriers #3 (15 out) and
#5 (20 out). Ran 5 more wafers and recieved the same alarm. Ran out
of time.
|
19.140 | Vacuum problem / DI water leak | FABSIX::R_GEE | | Sun Apr 07 1996 08:10 | 31 |
| "C" shift put the system down for a vacuum problem and a DI
water leak under the scrub station.
- We found the main vacuum gauge was only reading 22 lbs and
head #4 gauge was reading only 18 lbs. We cleaned the vacuum
pump diaphragm all the connectors and the in-line filter. The
main vacuum gauge now reads 30 lbs and head #4 25 lbs.
We began running wafers and found that carrier #2 misloaded
twice in a row. We checked the cup to spindle alignment
and the carrier to load cup alignment, both looked good.
Checked the load flipper to load cup alignment and found
that the wafers were not seating into the load cups far
enough. Adjusted the load flipper 75 steps back, all set.
- A hole in the steel main frame under the scrub station was
siliconed on "D" shift. We found that the leak just moved to
the next hole in the steel frame. The root cause for the leak
was a rip in the flex hose that houses the scrub station
power cable. The DI water was running down the inside of the
flex hose and down onto the main frame. We cut the power
cable, removed the scrub station and replaced the flex hose
then soldered and re-installed the assembly.
When we removed the top section of the scrub station we found
that all six stainless steel bolts had been stripped in.
One bolt broke off when we attempted to remove it. The others
needed to be replaced. The scrubber will work fine, but
a replacement part will need to be ordered.
B shift
|
19.141 | chiller fault | FABSIX::B_FINN | | Sun Apr 07 1996 20:17 | 7 |
|
Found B2 chiller with fault and display of 12 degrees, reset and
watched it recover to 24 degrees. aok rest of shift.
A shift
|
19.142 | SCRATCHES | LUDWIG::WESTPHAL | | Thu Apr 11 1996 20:35 | 5 |
| CMP.B2 - Down for scratches. Cleaned conditioners and process area but did not
pull pads as requested by process. Also drained and pulled the load
cups and wiped out the index table and unload tracks. The scrubber
brushes were checked and rinsed. Tool is presently getting MQCs
|
19.143 | Fatal Error | ASDG::POIRIER | | Sat Apr 13 1996 20:24 | 11 |
|
* Fatal Error:
The system experienced a fatal error at 5:00 p.m.
Gwen said that the screen went crazy. It looked like
lines of code zipping around the screen instead of
the menu screens. She EMO'd the tool and rebooted without
any further problems.
|
19.144 | Index table air cylinders | FABSIX::R_GEE | | Sun Apr 14 1996 07:50 | 7 |
| Errored for "All cups not down", found #3 & #5's air cylinders
sticking. Lubricated the air cylinder shafts and adjusted the air
pressure for cup #5, all set.
B shift
|
19.145 | secondary pad replaced / scrub station | FABSIX::R_GEE | | Mon Apr 15 1996 08:34 | 16 |
| * There was water leaking from the paddle wafer pedestal area.
Removed and inspected the wafer pedestals housing. Inspected,
cleaned and reseated the o-rings. Re-assembled the unit and
re-aligned the unload flipper fingers.
* Repaired the flex hose that houses the unload flipper cabling,
it was cracked.
* During the secondary polish an annoying squealing sound was
noticed. Before the scheduled MQC's we changed the secondary
pad.
B shift
|
19.146 | Loud noises, no slurry flow | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Apr 17 1996 05:57 | 13 |
|
During polishing, a loud noise was emitting from the
table due to poor slurry flow. The slurry lines on both
pump showed excessive wear. Both were replaced as they
were going to be done because of the PM due this week.
After everything was tested, air was coming through the
lines. It ended up being that the slurry pumps downstairs
shutdown due to a clogged filter. After the filters were
swapped over, the flow was back to normal and without any
air in the lines.
BJ
|
19.147 | Leak found, float switch not working | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Sat Apr 20 1996 01:03 | 14 |
|
During the weekly PM yesterday, DI water was found
once again in the front containment pan. A lot of it!
We eliminated the scrubber and DI was still coming down.
When then by-passed elevator #1 and drained all the water
out of it. The leak stopped.... We found the elevator #1
DI water fill float not working. We traced it to a broken
wire #13 at BB-10. After the wire was repaired, the switch
was once again working and as of right now, the is no leak.
D shift
a.k.a. Digital Water Rescue Team
|
19.148 | Pad change due to fab being dumped | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Sat Apr 20 1996 08:09 | 8 |
|
the fab dumped at 4:15 and B2 was still running
product. The tool was placed in wet idle upon return
but both pads were slightly dried up and needed to be
replaced.
D shift
|
19.149 | Unload flipper | FABSIX::R_GEE | | Sun Apr 21 1996 07:49 | 13 |
| The system errored for "wafer not detected in the unload flipper".
We found the wafer from unload cup #4 was only grasped by the
two rear unload fingers. We repositioned the wafer and pressed
start to continue. We checked all the alignments and didn't
find any problems, the system ran fine the rest of the night.
Five wafers got over polished due to the error. When the system
was restarted the five wafers on the carriers got polished a
second time. This wasn't noticed until after thickness measure.
B shift
|
19.150 | FLIPPER BOX CABLE | FABSIX::B_FINN | | Mon Apr 22 1996 19:43 | 26 |
|
System B2 unload station
System intermittently has strange errors like releasing the wafer at it's home
position on the way to the scrubber, and system lockup during unload.
Found outer corrugated cable on flipper box torn. Wafer drain is dumping water
on wires and into outer cable sheath. Wires are being stressed as
flipper gripper moves from cup to clean station.
Parts not in stock, FS is attempting to order "next day".
AC2535A1 TUBING, LIQ-T,1/",PLASTIC, DRAWING AA3037 ITEM 9
AC2534A1 FITTING, 90D EL, LIQ-T,1/2"TX Same drawing ITEM 10
This is not the same design flex tube as the scrubber uses. The end of
the tube is threaded and screws into the connector fitting. We had to
tape the sections together to removed the stress on the wires, as a
temporary water proof "band aid" until parts come,in hopefully tomorrow.
System has run the rest of the day fine.
A SHIFT/ RICH
|
19.151 | magnetic door closure vote | STRATA::RIDLON | Eliminate the obvious | Tue Apr 23 1996 02:46 | 6 |
| Installed magnetic door closures on the pneumatic cabinet. Let me know
what you guys think. I would like to install them on the cpu cabinet as
well but I'll wait for a majority vote.
B shift
|
19.152 | hung up | STRATA::RIDLON | Eliminate the obvious | Wed Apr 24 1996 07:42 | 9 |
| Tool was loading product and got hung up after loading three wafers
into the index table. No alarms, the screen just stayed at "waiting for
load flipper at home" and the index table did not move to the next load
cup. Rebooted and it worked fine the rest of the night. The secondary
pad is whistling when wafers are cleaning again. There are some bubbles
but we didn't replace it due to B1 being down most of the night
B shift
|
19.153 | Wafer breakage | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Thu Apr 25 1996 07:28 | 11 |
|
Early in the shift, wafers slid out of carrier #1 and
#4. The tool didn't crash however. Several hours later,
it did on the other hand. A product wafer slid out of #1
and took out #5 and #4 before the tool caught it. The entire
tool was cleaned up, pad replaced and all (5) carriers were
replaced due to the crash and high wafer count. The tool ran
flawless for the rest of the shift.
D shift
|
19.154 | Corrected wafer breakage passdown | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Thu Apr 25 1996 09:35 | 13 |
|
Early in the shift, wafers slid out of carrier #1 and
#4. The tool didn't crash however. Several hours later,
it did on the other hand. A product wafer slid out of #1
and took out #5 and #4 before the tool caught it. All of the
cup to head alignments were checked as well as the load
handling. Hi Tim! 8^) The entire tool was cleaned up, pad
replaced and all (5) carriers were replaced due to the crash
and high wafer count. The tool ran flawless for the rest
of the shift.
D shift
|
19.155 | FLY OUTS | LUDWIG::WESTPHAL | | Thu Apr 25 1996 21:15 | 12 |
|
CMP.B1 - Down for slipping wafers from the previous shift. Pad was changed and
the tool was cleaned up. The alignments were checked and the index
table and all carriers were adjusted. The tool slipped another wafer
at lift of but no wafers were broke. The reps. suspected that there
was a problem w/ the pads, so they tried breaking the pad in using a
modified conditioning recipe. Fifty wafers were run and It seemed to
have worked. While running MQCs 2 wafers came out of heads 1&5.
inverter errors were recieved and the screen locked up all at once.
The tool presently sets rebooted with all wafers off the system in wet
idle.
|
19.156 | mag door closures | STRATA::RIDLON | Eliminate the obvious | Mon Apr 29 1996 02:33 | 3 |
| Installed magnetic door closures on the cpu and di cabinets. I'll
have to scrounge up some more before doing b1.
|
19.157 | AB Boxes | FABSIX::R_GEE | | Sun May 05 1996 06:47 | 5 |
| Mounted the 2 AB boxes to the side of the computer
cabinet.
B shift
|
19.158 | FAILED MQC | FABSIX::B_FINN | | Sun May 05 1996 21:07 | 7 |
|
B2 failed MCQ's for particles, replaced polish and secondary
pads. System passed MCQ's aok.
A shift
|
19.159 | slurry & brushes | SUBPAC::LANDRY | | Tue May 07 1996 20:36 | 13 |
|
Early morning......
The system was EMO'd due to a lack of and/or no slurry flow.
This was evident as the system became load and vibratation was
observed. Rebooted the system. Ran a manual polish segment to
assure that slurry was flowing properly (it was). Also, changed
the scrubber brushes at this time. Returned the system to MFG
for MQC's. All is well.
Late afternoon, the room had a "power sag" which caused the room
lights to noticeably waver. The system was un-affected though.
Doc
|
19.160 | scratches | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat May 11 1996 06:18 | 9 |
|
Tool down for consumables when we came in. A set
of new carriers and scrub brushes were installed on the
tool. Tool passed MQC's and was turned back over to
production.
D shift
|
19.161 | VAC PROBLEMS/CRASH ON SECONDARY TABLE | ASDG::POIRIER | | Sat May 11 1996 20:47 | 30 |
|
Vacuum Errors:
-------------
The system experienced many vacuum errors during the shift.
We noticed that the vacuum seperator was 1/2 full of slurry.
The slurry was drained and the system ran for a while without
any errors. After a few hours we got an MAIN VAC TOO LOW
error. The vac gage at the pump was displaying 29 in/hg which
seemed unusual so we opened up the vac separator again and discovered
that the filter element was clogged with dried slurry. The element
was removed and the vac system has been performing normally ever since.
The manufacturing folks have been performing a double map at the
begining of each run to ensure that the tank is thoroughly drained and
that no slurry will be pulled into the pump. Please continue to do
this until a vacuum separator filter element can be located.
Also, the vac tank blow out cycle seems to last only a few seconds.
Shouldn't it be 15 seconds?
Wafer crash on the secondary table:
----------------------------------
The system broke a wafer on the secondary table at 7:00 p.m. The
contamination was pretty extensive so we are cleaning the entire
tool and changing the primary pad, secondary pad & scrubber brushes.
All cups have been removed from the index table and the tub thoroughly
wiped out and rinsed. PLEASE CLEAN THE CUPS CAREFULLY AS SOME OF THEM
HAVE SILICON PARTICLES DOWN INBETWEEN THE SLIDING PATS OF THE CUPS.
SOME OF THE CUPS MAY HAVE TO BE DISASSEMBLED.
|
19.162 | Crash on 2nd table (Cont.) | FABSIX::R_POIRIER | | Sun May 12 1996 07:50 | 14 |
|
Finished Load cup alignments started by "C" shift.
Load cup 4 needed slight alignment (256 counts). All
others ok. Took apart all Load/Unload cups to clean out
wafer particles. Load cup 3 had particles imbeded in the
middle of the cup. Replaced the cup with one from the
carrier rebuild room. Cleaned index table and scrubber,
replaced the scrubber brushes, both Polish pad and second
table pad. Cleaned conditioner.
Turned B2 back over to Operations for break-in and MQC's.
B2 Passed MQC's.
|
19.163 | switch boxes | STRATA::FINN | | Mon May 13 1996 21:09 | 10 |
|
Drilled hole in the side of the computer cabinet and ran cables for
remote monitor. Picked up 9 pin m/m 6 foot connector at radio shack and
completed wiring. Tested just now and screen lights up and displays
data, but touch screen is not working. started to troubleshoot but ran
out of time.
A shift / Tim
|
19.164 | passdown | STRATA::FINN | | Tue May 14 1996 21:33 | 18 |
|
Reps tried to upgrade CMP.b2 to 493 software. They found the keyboard
jack on cpu bad. No CPU's in stock. Kevin was trying to get one on
order. System running is fine but you cant access dos via keyboard.
Sent 30 pressure plates to speedfam today.
Continuation of yesterdays Problem (remote touch screen not working)
Found open in remote switch box for com1. Open is on pin seven of the
B side of the switch. Tried to clean but its still intermittently open
depending on how much pressure is placed on rotary switch. Pin seven
of B is the only pin that has this problem. John P. could you get a
return on this, not sure where you bought it?
A shift
|
19.165 | Monthly PM, crash and a the 24V supply | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu May 16 1996 08:22 | 13 |
|
The tool lost slurry slurry and is the suspect of a
crash on the 2nd table. A wafer looks like it cracked
on the polish and blew up on the 2nd table. Ashland
called the first time to say they were shutting down the
slurry but it went down unexpectedly a 2nd time. We
took the tool for the monthly and were making progress
until we lost all of the 24V in the tool. The tool is
dead in the water and we suspect it is the 24V power
supply. We are in the middle of a swap out now.
BJ
|
19.166 | Vacuum problems..... | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri May 17 1996 08:35 | 32 |
|
a long time ago, in a galaxy far far away....
opps... wrong story....
The tool was still down for a couple of PM issues.
The monthly PM was completed and carriers were put back
on the tool. While the pad was being conditioned, I had
the carriers rotating with DI rinse on. Only #1 was
spraying. After some troubleshootin', the 24V line that
feed the output OPTO board was accidentally placed in the
wrong hole from yesterdays 24V problem. Rinse #1 is not
in the same row as #2-#5. That was now back to normal.
OPS started breaking in the pad and we started get "no
wafers on carriers" error. The samething OPS were getting
just before the crash on Tuesday before the PM. So, we're
back to the original problem. After some more hard
troubleshootin', we noticed that there was vacuum when
multihead was at polish but not at load. Kinky line in
flex track. Yup... the new 1/2" clear vacuum line was
just a snudge too long and was kinking when the multihead
traveled. Took out the slack in the line and that was
all better. We then had #1, #2 and #5 not seeing vacuum
with a wafer in place and vacuum at 20". All of the vacuum
pressure switches we re-checked and slightly adjusted to
spec. The tool 40 wafers with I_O program #1 with no
problems. Tool is in verify but it needs a proper pad
break-in done.
BJ
|
19.167 | bad infranor idex table | FABSIX::B_FINN | | Mon May 20 1996 20:40 | 20 |
|
Index Table Time out error, Table would not move. Found Motor
#11's infranor failed for no power. Its breaker and infranor power
supply were fine. Changed with 14 amp infranor from F.S. office. That
Infranor had current error. Changed with infranor from carrier
rotation # 4. Index table was fine and problem followed to Carrier
#4. No other Infranor's in office or cabinet. Borrowed infranor from
B3 carrier #4 at F.S. request. System ran fine.
REp's have infranor coming in tomorrow.
During mqc's wafer cracked on table during segment 3. Changed
one carrier and pad.
Latter in the evening one wafer was left on pad at lift off.
Removed covers and checked all vacuum while running dummies, Pump was
28 and carrier with wafers was 24-25. Ran a couple of runs and gave
back with panel off so we can monitor. System is currently running.
A shift
|
19.168 | system pausing during polish | FABSIX::R_GEE | | Tue May 21 1996 08:09 | 9 |
| The system wouldn't load the next 5 wafers into the index
table until after it finished polishing the 5 on the carriers.
Rebooted the system and the problem went away.
B shift
|
19.169 | 493C upgrade completed | SUBPAC::CMP | | Tue May 21 1996 16:55 | 3 |
| installed 493C, set autoexec.bat for 493C boot.
tested 25 wafers and ran good with no hesitation.
Returned to production.
|
19.170 | Bootup problems / A/B box | FABSIX::R_GEE | | Sun May 26 1996 07:47 | 13 |
| Came in tonight and the touch screen had froze up,
tried to reboot the system and came up with 2 errors.
The first was a Dos error " DIVIDE OVERLOAD " the
second error was " ELODEV NOT INSTALLED ", it wouldn't
boot to the Initialization screen. By bypassing the
A/B box the problem went away. We couldn't troubleshoot
any further because ops. needed the tool. We believe
the A/B box needs to be replaced.
B shift & Dave B.
|
19.171 | carrier 2 no rotation | FABSIX::B_FINN | | Sun May 26 1996 20:22 | 11 |
|
System lost a wafer at lift off, manufacturing tech place it back
on carrier, but carrier 2 failed for rotation. System was powered down.
We brought system up. Check for breakers or infranor errors found
infranor for carrier 2 rotation had no power. Hit init and power was
applied to the infranor, STRANGE. tried to replicate but could not.
Ran test wafers aok 400lbs.
A shift
|
19.172 | Particles/Head #4 Flyout/Special Primary Pad On Tool | ASDG::POIRIER | | Wed May 29 1996 20:39 | 17 |
|
Particles:
----------
Both Primary & Secondary Pads were changed due to high particles.
Special Pad:
-----------
A Special Primary Pad was installed at 6:00 p.m. Do not remove this
Pad without calling Frank Krupa at home.
Wafer Flyout on Head #4:
-----------------------
A wafer fly out of Head #4 at lift-off. There was no breakage and we
recovered with no problems.
|
19.173 | INDEX TABLE | FABSIX::J_SWENSON | | Sat Jun 01 1996 20:57 | 12 |
| SYSTEM CONFUSED INDEXED WITH CUP CYLYNDER'S UP.
TABLE WAS JAMMED AGAINST CUP'S.
WENT INTO DCX1 INDEXED TABLE OFF OF CYLYNDERS.
CHECKED CYLYNDERS LOOKED O.K.
REMOVED CARRIERS THEY ARE SOAKING ON STAINLESS TABLE.
PAD IS IN ONLY WET PAD MODE.
CUP'S REINSTALLED
SYSTEM SITTING WITH TABLE AT HOME POS.
NEED TO VERIFY CUP TO CARRIER ALIGNMENT.
DOC AND HAWK
|
19.174 | Index table crash | FABSIX::R_GEE | | Sun Jun 02 1996 07:52 | 13 |
| The system went down for an index table to all cup cylinders
up crash. Checked the spindle to load cup alignment and the
carrier to load cup alignments, all looked good. Checked
the locate.cmp program and the size of speed493.exe in dos,
everything looked normal. Checked the error log and found
no errors logged for the crash. Replaced the pad due to high
wafer count, ran 25 wafers using no_touch. Gave the system back
to production, had no errors the rest of the night.
B shift
|
19.175 | crash......MQC wafers | SUBPAC::LANDRY | | Mon Jun 03 1996 20:34 | 7 |
| The system slipped an exercise wafer out of head four this afternoon.
It wiped out a few product wafers and a couple of carriers. The system
has been cleaned, new pad and some carriers installed. Pad
conditioned. But, currently waiting for MQC wafers to arrive to CMP.
Doc
|
19.176 | B2 the water fountain | SUBPAC::CMP | | Tue Jun 04 1996 17:52 | 2 |
| today the water line in the flex track ruptured and dumped water in
the tool. Parts are on order and will arrive in the A.M.
|
19.177 | new di supply line installed | SUBPAC::LANDRY | | Wed Jun 05 1996 20:36 | 6 |
| The new flex tubing for the MHTA DI water supply was installed today.
It was leak checked and proved water tight. A new pad was installed
and the original carriers re-installed. MQC's looked good.
Maintenance services provided by Hawk and Rich for this fix.
|
19.178 | Wafer breakage | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Jun 06 1996 07:48 | 22 |
|
The tool had just got done running a METAL-3 lot
without any problems. OPS started up a PMD touch-up
lot with (5) wafers. Carrier spit a wafer and the
wafer detect system didn't appear to pick the the
slide-out. The error that trip was MAIN VACUUM LOW.
Head #2 wasn't rotating either. Upon boot up, the
vacuum error was still present. We found breaker #13
tripped. Head was also not rotating. After a 2nd
power-up, everything was ok. The downforce on #2
was checked and was ok. The head to load cup alignment
and head to index table alignment were fine. Since
everything looked ok, we compared the voltage on
head #1 and #2 Rotation Infranors voltages. #1 was at
10V, #2 was at 9V. We tweaked #2 to 10V to see if we
saw any change. They ran break-ins without any problems.
OPS ran MQC's and passed particles and and uniformity but
the range is out. They are running more wafers now and
it looks like it will come in.
D shift
|
19.179 | rotation error on B2 | SUBPAC::CMP | | Thu Jun 06 1996 11:31 | 2 |
| Replaced infranor for head 2 rotation error. The infranor did not have
any LED's lit and as a result no output!!!.
|
19.180 | Waferfly | FABSIX::L_SHKOLNIKOV | | Fri Jun 07 1996 07:07 | 9 |
|
Tool was leaving wafers on the polish table after lift
off from carriers #1 and #5 (4)times tonight. It was causing
vacuum low alarms. This was an issue before, so I checked
Downforce, load cup to carrier alignment on the carr #1 and #5
and vacuum, looked good. I have changed the pad and carriers 1 and 5
and it seemed to have worked. The tool passed MQC's and is running
product without any problems.
|
19.181 | | ASDG::POIRIER | | Sat Jun 08 1996 09:41 | 2 |
|
|
19.182 | Unload Flipper Sensor Error / Water In The Box. | ASDG::POIRIER | | Sat Jun 08 1996 10:04 | 20 |
|
Unload Flipper Sensor Error:
---------------------------
The "Unload Flipper Wafer Present" Sensor was not enabled during a
product run at 4:00 p.m. on Friday. (DG-IN 7, PROX6)
We opened up the unload flipper box and found water inside. The
metal air cylinder bracket that sweeps over the 3 proximity
sensors was dragging water across all 3 sensors and confusing
the auto cycle program.
We dried up the water, cleaned the gaskets and applied a coating of
DOW Corning Electrical Compound to the mating surfaces and screw
holes. We avoided using Silicone Rubber due to the long curing time
and the difficulty it causes during disassembly.
The system ran for the rest of the shift with no unload flipper errors.
C Shift.
|
19.183 | system locked up | FABSIX::R_GEE | | Mon Jun 10 1996 08:05 | 9 |
| The system locked up in the middle of a 25 wafer run.
It completed polishing the wafers on the carriers then
froze. We had to reset the computer then remove the
wafers in manual.
B shift
|
19.184 | passdown | FABSIX::B_FINN | | Mon Jun 10 1996 20:38 | 19 |
|
System is in 24 hour required MQC. Today Multihead went out to
polish table and stopped. System alarmed for process door switch open.
Found rear right sliding door micro switch not made. Rich adjusted
micro switch. System failed for main vac operator powered system
down. At init failed for "No rotation head 3", manufacturing
tech hit any key to continue and software dumped him to DOS. I
reloaded Speed493 ran aok. Watched several wafer maps dump seems ok.
But when pressure equalizes at end of dump and isolation valve opens
pressure/vac drops close 15" (alarm point).
Tim would like sliding doors and stitches checked and repaired. (no done yet)
I spent rest day doing the same alignments to b "Pain in the
backside"! (The doors and rollers not Tim!)
System just passed mqc's
A shift
|
19.185 | rollers for doors | SUBPAC::CMP | | Tue Jun 11 1996 08:43 | 7 |
| Just a word on the rollers for the doors, we have an upgrade that will
take place when the parts are available, so at this time we will wait
for them to arrive instead of replacing with old style parts.
SF
|
19.186 | Remote Touchscreen Back In Service | ASDG::POIRIER | | Wed Jun 12 1996 20:22 | 7 |
|
* Rear Touchscreen Monitor:
------------------------
The defective A/B Box for the remote touchscreen monitor has been
replaced with a new one. The cables have been reconnected and the
monitor is fuctioning normally.
C Shift.
|
19.187 | Head #3 rotation, slurry chunks | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Jun 13 1996 00:24 | 14 |
|
OPS started running a lot after the tool was sitting in WET
IDLE for some time. While the MHTA was going to polish with (5)
send-aheads, a lot of slurry chunks were noticed on the 2nd table.
The tool was EMO'ed before polish. The tool sat off for 10 minutes
while the tables were sprayed off. At initialization, head #3 wouldn't
rotate. The rotation Infranor for head #3 cooked, no LED display with
a slight burnt odor. The Infranor was replaced at the volts set to 11.0.
The carriers were scrubbed clean an the pad rinsed off thoroughly.
At VERIFY, the tool mis-loaded a wafer on head #4. Started over and
it was ok. OPS ran particles wafers to verify the tool and the tool
passed and came back up.
D shift
|
19.188 | Main Vac Too Low Error / Leadscrew Covers Loose | ASDG::POIRIER | | Thu Jun 13 1996 21:07 | 20 |
|
* MAIN VACUUM TOO LOW ERROR:
-------------------------
The system errored for "Main Vac Too Low" this afternoon. A reboot
was necessary to recover from this condition. The system ran for the
rest of the shift with no problems.
* MULTIHEAD LEADSCREW COVERS:
--------------------------
A screw backed out of one of the covers yesterday causing the cover
to swing sideways. We replaced the screw only to discover that the
spacer was missing. It must have fallen into the tub under the table.
The spacer is flat on one side and spherical on the other. It looks
like a Speedfam part. We asked the reps to order replacement hardware
for the covers. Further inspection revealed additional loose covers.
C Shift.
|
19.189 | pad's changed, system cleaned | FABSIX::J_SWENSON | | Sat Jun 15 1996 20:21 | 7 |
| cleaned system.
changed primary and secondary pads.
primary pad count 1844
secondary pad count 1859
system passed mqc and is running production.
c-shift
|
19.190 | moisture sense module offf | STRATA::RIDLON | Eliminate the obvious | Mon Jun 17 1996 01:20 | 11 |
| Found the leak detector in the slurry pump box to be red. Checked the
detector in the fluids cabinet and it was also red (and in water).
Checked Hardware.cmp file and the moisture sensors were enabled. Checked
digin 27 and this was not highlighted. Checked voltage at the digin
opto-isolator and both sides had ~24vdc. Checked the leak sensor
module, LCK-1000, in the fluids cabinet and this was off. Turned it on
and the tool alarmed for moisture detected. Dried up water and placed
new breast feeding pads under the sensors.
B-shift
|
19.191 | cross slotting | STRATA::RIDLON | Eliminate the obvious | Mon Jun 17 1996 05:30 | 7 |
| Wafers cross slotting in elev 1. Ended up lowering the water level,
reducing the water track flow, and lowering elev 1 home by 150 counts.
Manually unloaded a cassette with no cross slotting. Also checked the
next cassette of production wafers and there was no cross slotting.
B-shift
|
19.192 | moisture sensor alarm | SUBPAC::LANDRY | | Mon Jun 17 1996 20:48 | 9 |
| The Moisture Sensor alarmed today..........
Traced problem to the Fluids cabinet. Found a leaking fitting coming
out of the exit cooling line valve within the cabinet. Retaped the
fitting and re-installed. Changed the oppropriate sensor paper.
Re-powered the moisture detection power unit. All set. Checked for
leaks several hours later, none to be found.
A
|
19.193 | | FABSIX::B_FINN | | Tue Jun 18 1996 20:05 | 19 |
|
System is in 24 hour required MQC. Today Multihead went out to
polish table and stopped. System alarmed for process door switch open.
Found rear right sliding door micro switch not made. Rich adjusted
micro switch. System failed for main vac operator powered system
down. At init failed for "No rotation head 3", manufacturing
tech hit any key to continue and software dumped him to DOS. I
reloaded Speed493 ran aok. Watched several wafer maps dump seems ok.
But when pressure equalizes at end of dump and isolation valve opens
pressure/vac drops close 15" (alarm point).
Tim would like sliding doors and stitches checked and repaired. (no done yet)
I spent rest day doing the same alignments to b "Pain in the
backside"! (The doors and rollers not Tim!)
System just passed mqc's
A shift
|
19.194 | slurry sensor added | FABSIX::B_FINN | | Tue Jun 18 1996 20:07 | 4 |
|
Installed the slurry detection sensors on b2. Calibrated and
tested. aok
A shift
|
19.195 | Wafer crash | FABSIX::L_SHKOLNIKOV | | Thu Jun 20 1996 07:48 | 8 |
| The system alarmed today for carrier #3 not rotating.
The system slipped an product wafer (on last sec. in segment 5)
out of head two this night. It took out #5 and #4 before the tool
caught it. The system has been cleaned, new pad and (3) new carriers
installed. Pad conditioned. Currently running MQC.
d-shift
|
19.196 | down force calibrated due to head 2 removal rates. | FABSIX::J_SWENSON | | Sat Jun 22 1996 20:11 | 0 |
19.197 | down force calibrated due to head 2 removal rates | FABSIX::J_SWENSON | | Sat Jun 22 1996 20:22 | 25 |
| Informed that head 2 had high removal rates.
Observed carrier offsets to be out of the +/- 30 range
carrier 1 2 3 4 5
+ 38 - 37 + 33 + 29 + 20
Performed Down Force Calibration for all five heads.
carrier 1 2 3 4 5 pre cal
365 lbs 404 lbs 377 lbs 368 lbs 390 lbs
carrier 1 2 3 4 5
377 +/- 1% --------------------------------> post
System in wet idle waiting for MQC'S
C shift
|
19.198 | Broken wafer | FABSIX::R_GEE | | Tue Jun 25 1996 09:29 | 16 |
| A monitor wafer slid out from under carrier #5 at the end of a
5 sec. polish (step #5). Due to the mess and high wafer count
it was determined that all 5 carriers should be replaced.
465 wafers on the pad
199 on head #5
We replaced the primary pad and all 5 carriers. The system has
been cleaned including the conditioner. We didn't change the
secondary pad or condition the primary pad.
The system is in wet mode and down_repair in workstream.
B shift
|
19.199 | passdown | FABSIX::B_FINN | | Tue Jun 25 1996 20:37 | 10 |
|
Installed 3/8" clamp for slurry sensor and calibrated, test
aok while room was shut down for slurry today. This amp is calibrated
to switch position "C".
A shift
|
19.200 | rotation, door switch and pad issues | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Jun 26 1996 07:22 | 17 |
|
The tool first errored for heads not rotating at segment #3
after touch down. While trying to recover, the tool said there
was a process door open but they were all closed. All doors were
checked and the culprit was the double doors in back. The left
microswitch wasn't making good contact, repaired same.
Later on, at touchdown, there was a loud bang that sounded
like the belt jumping and tool hung up for heads not rotating.
Leon went to hose off the pad when he noticed the top pad all
separated from the Suba pad. I don't know what was holding
down the pad but it wasn't by much. After the pad was replaced,
the tool ran fine for the rest of the morning.
D shift
|
19.201 | Particles | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Jun 29 1996 08:10 | 10 |
|
Tool never came up from ENG experiment. Tool
had massive particles. The pads were changed,
everything cleaned. If you use perfect particle
wafers on no touch and if the brushes are in
question they aren't going to take off something
that wasn't there. I swapped the top brushes
with the bottom brushes and particles passed...
wow..... BJ
|
19.202 | Wafer crash | LUDWIG::FINN | | Mon Jul 01 1996 20:44 | 8 |
|
System crashed while engineering was running 5 breakin wafers. Crash
occurred during segment 5. All five carriers and pad were replaced.
Wafer detect alarm did not function, it was found to be jumpered out.
Cup and carrier aliments were checked and found to be aok. Wafer detect
system is now operative.
A shift
|
19.203 | Start up B2 | LUDWIG::SWENSON | | Fri Jul 05 1996 20:29 | 12 |
|
powered system up no problems
installed 5 new carriers
performed conditioner arm cal, found to be approximately @72 lbs @ 67%
set to 76. 8 +/- 1 lb
cut primary pad.
retrieved new brushes from stockroom for scrubber.
C Shift
|
19.204 | Conditioner down force | FABSIX::R_GEE | | Tue Jul 09 1996 06:11 | 8 |
| Calibrated the conditioner down force using the procedure
in note 23.1. With the analog output at 4095 the force
was 54 lbs. we set it to 77 lbs.
B shift
|
19.205 | New Blocky Conditioner installed | STRATA::KRUPA | | Thu Jul 11 1996 18:15 | 2 |
| New Blocky conditioner installed on evening of 7/10/96.
|
19.206 | SCRUBBER TROUBLE | STRATA::WESTPHAL | | Sat Jul 20 1996 20:38 | 8 |
| CMP.B2 - Down for high particles. The scrub brushes were not rotating. The
scrub assembly was pulled and the gear on the motor shaft was found
to be stripped. Some bearings are also bad and most of the screws
were damaged. The upper and lower scrub assemblies were swapped with
CMP.B3's because it was down. Speedfam has been notified and the
parts will be ordered. The tool was cleaned up and returned to
production.
|
19.207 | scrubber brushes changed | SUBPAC::LANDRY | | Sun Jul 21 1996 19:32 | 6 |
|
The brushes on the scrubber were changed out today! What a
pleasure it was to work with the "user friendly" 3M adhesive backings.
A shift
|
19.208 | srubber not working | FABSIX::L_SHKOLNIKOV | | Wed Jul 24 1996 08:02 | 12 |
| CMP.B2 - Went down for high particles. The scrub brushes were rotating
but lower scrubber assy wasn't going up. The scrub assembly was
taken apart and teflon coupling that connect lower and upper motor
aasy was found to be stripped. Lower motor shaft sticking inside of
coupling and not going up. It needs to be replaced. I will speak with
reps in the morning. For now I temporarily swapped with B1 which
went down for high particles. The tool was cleaned up and returned
to production. Running MQC's now.
Leon
|
19.209 | particles passed | FABSIX::R_GEE | | Tue Jul 30 1996 07:51 | 10 |
| The system failed mqc's for high particles. We followed the
game plan issued by process engineering. Put a precut primary
pad from lot #6EN-JP4 on. Ran the regular breakin and mqc's
with steps #8 & #9's ID OD set at 12.7 The system passed mqc's
well within spec. released to production.
B shift
|
19.210 | b2 scrubber | SUBPAC::CMP | | Thu Aug 01 1996 15:46 | 2 |
| checked and adjusted scrubber rpm, (rpm were 3x more than normal)
|
19.211 | scrubber gremlins | LUDWIG::RIDLON | Eliminate the obvious | Sat Aug 03 1996 23:03 | 10 |
| Noticed that the scrubber wasn't operating properly. It was going in
fine but the motor wasn't turning on and the base wasn't going up.
All the proper sensors were being made at the proper times. Excerised
it using the Dig/out and Anl/out screens but it didn't start working
until the motor was physically turned by hand. After that it worked
fine no matter what screen was used. Told ops to keep an eye on it.
B-shift
|
19.212 | Scrubber belt snapped | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Mon Aug 05 1996 14:47 | 8 |
|
We noticed a product wafer wasn't rotating during
the scrub process. While running the scrubber manually,
the bottom brushes weren't spinning. Upon disassembly,
it was discovered that the drive belt had snapped. A new
belt was installed and test.
A shift
|
19.213 | guess. | STRATA::RIDLON | Eliminate the obvious | Tue Aug 06 1996 07:34 | 8 |
| Tool was down when we came in for particles. Both pads were pulled by
A-shift. The first set of mqc's failed. The carriers were scrubbed
and the polish area was given a once over again. The second set of
mqc's weren't measured as a wafer came off of carrier 1 on lift off.
A third set of mqc's will be attempted shortly.
B-shift
|
19.214 | MONTHLY WAVIED | STRATA::WESTPHAL | | Fri Aug 09 1996 20:22 | 4 |
|
CMP.B2 - The monthly has been wavied until next week.
|
19.215 | High particles | FABSIX::R_GEE | | Mon Aug 12 1996 08:25 | 12 |
| The system failed mqc's for high particles. Found the secondary
pad was fraying in the center and around the outside edge. The
pad was cut to fit, originaly ment for cmp.b3. We cleaned
the conditioner and index table, replaced the secondary pad
with one precut for cmp.b2 and particles passed. The particle
counts for this tool are running high, above 1000, with cmp.b1
& b3 running around 300.
B shift
|
19.216 | Broken wafer / vibration | FABSIX::R_GEE | | Tue Aug 13 1996 08:22 | 20 |
| The system broke 2 wafers on the polish table.
After a new pad was installed the system passed mqc's, we
began running product. When we started running the 1 second
touchup runs a loud vibration was noticed, usually coming off
and on in steps #6 thru #8. On this run the vibration occurred
in step #6, the vibration was so bad that one of the process
doors opened and the system stopped. We were watching the system
at the time and everything looked normal except for the loud
vibration. After closing the door we attempted to restart the run,
when the wafers went down onto the polish table #5 slid out and
broke on head #4.
We think the vibration is created by very little slurry on the table
while running only a 1 second touchup and the new water polish steps.
There was no vibration noticed during the initial long polish.
B shift
|
19.217 | Sorry A-shift. | STRATA::RIDLON | Eliminate the obvious | Wed Aug 14 1996 08:06 | 12 |
| Broke wafers on pad. Operator didn't know what segment but apparently
the wafer sense didn't work and the tool had to be emo'd. Two infranors
had to be replaced upon power up (motor 10 and 13). After tool was
cleaned up breakins were ran. It dropped a wafer on lift off in segment
9 out of carrier 2. Under carrier 2 was found a sliver of wafer which
may contributed to the problem. Carriers 3 and 4 were replaced after
the initial crash the others were deemed ok. Leaving in wet mode for
A-shift eval.
B-shift (it was a lousy night).
|
19.218 | MONTHLY PM PARTIALLY COMPLETED & DEFERRED TO WW8. | ASDG::POIRIER | | Sat Aug 17 1996 20:46 | 13 |
|
Monthly PM Deferred to WW8:
--------------------------
We have partially completed the WW7 Monthly P.M. Due equipment
availability, we could not complete the Monthly pm today.
The partially completed checksheet has been posted in the Incomplete
PM Checksheet note.
|
19.219 | vacuum pump relocated | FABSIX::R_GEE | | Tue Aug 20 1996 08:26 | 7 |
| Relocated vacuum pump to the black box below the floor.
Test fine.
B-shift
|
19.220 | Head #5 Fly -outs / Vacuum not present at touch down | ASDG::POIRIER | | Thu Aug 22 1996 20:45 | 11 |
|
Head #5 Dropping wafers / Vacuum trouble:
----------------------------------------
Just finished a test run. The multihead covers were removed on head #5.
The vacuum gage displayed 21 in hg as the poish cycle started. As soon
as the multihead arrived over the polish table, and the carriers
started to rotate, the vacuum dropped to zero. The wafer in carrer #5
flew-out at touch down. D shift, take a look at the rotory union on
head #5.
C Shift.
|
19.221 | | STRATA::WESTPHAL | | Thu Aug 22 1996 21:00 | 10 |
|
CMP.B2 - Down for losing wafers from head #5. The problem repeated itself
several times today. A different carrier was installed on #5 with the
same result. Removed the covers and read the gauge with vacuum drawn
on head #4 & #5. The vacuum was 26 at the carrier and 28 at the tank
for both. Ran out of time. Engineering requested that all the
fittings for head #5 be checked and the vacuum be rechecked at the
spindle #5 again with the carrier rotating.
|
19.222 | Carrier #5 Vacuum problems | FABSIX::O_WHITFORD | | Fri Aug 23 1996 07:02 | 35 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
-< CMP >-
================================================================================
Note 19.222 CMP.B2 EEG PASSDOWNS 222 of 222
FABSIX::O_WHITFORD 21 lines 23-AUG-1996 04:45
-< Carrier #5 Vacuum >-
--------------------------------------------------------------------------------
Took over from C shift.
Began by removing covers and panels for carrier 3 to make a
comparison to.
Loaded 5 wafers and noted carrier #5 @ 3"hg lower than #3. Ran up
carriers to approximate RPM during process with no change in vac which
pretty much eliminated the rotary union in the spindle.
Moved carriers to the "in" position and observed vac on carrier #5
drop to 0" then return to @ 22"hg. Observed the wiring bundle while the
arm was in motion looking for kink the vacuum line. When I touched
the bundle, reaching for the vacuum line, the solenoid chattered and vac
immediately dropped out.
Traced the wiring to the solenoid and found a broken wire in the 24V
line leading to the vac solenoid valve. The wire broke at the tiewrap
which anchored the bundle to the lower plate of the multihead. This was
the point where the wires made the greatest flex while the arm was
traversing from the "in" position to the range of oscillation. We cut into
the outer insulation and spliced the broken wires. Heat shrank the repaired
area to preclude moisture damage.
Exercised with no further problem. Turned over to ops for
verification. 25 wafers ran with no losses from carrier #5. MQC's,
uniformity, particles etc. came out well withing spec.
Leon S. Dave W.
|
19.223 | Wafer Misload on Head #3. | ASDG::POIRIER | | Sat Aug 24 1996 16:57 | 9 |
|
Wafer Misload:
-------------
The system failed to load a wafer on Head #3 this afternoon. We didn't
see exactly what happened. Its possible that Load Cup #3 may be
sticking or the lift cylinder might be degrading. We can't get the tool
at this time to check it out.
C Shift.
|
19.224 | PASSDOWN CMP B-SHIFT 8/25/96 | FABSIX::M_PROVENCHER | | Sun Aug 25 1996 07:56 | 108 |
|
CMP
___
monitor wafer usage
-------------------
Type of monitor Number used Comments
--------------- ----------- --------
MQC-12K MONS 16 MQC'S
BREAKIN WAFERS 42 misc
CMP.A1 (Strasbaugh 6DS-SP) - UP/IDLE
CMP.A1 24hr MQCS COMPLETED: @ 2:00 PM 8/24
CMP.A1 24hr MQCS NEXT DUE : @ 2:00 PM 8/25
CMP.A2 (Strasbaugh 6DS-SP) - UP/RUNNING XM0515 PMD
CMP.A2 24hr MQCS COMPLETED: @ 7:30 PM 8/24
CMP.A2 24hr MQCS NEXT DUE : @ 7:30 PM 8/25
CMP.B1 (Speedfam CMP-V) - UP/IDLE
CMP.B1 24hr MQCS COMPLETED: @ 3:00 AM 8/25
CMP.B1 24hr MQCS NEXT DUE : @ 3:00 AM 8/26
CMP.B1 12hr MQCS COMPLETED: @ 3:00 AM 8/25
CMP.B1 12hr MQCS NEXT DUE : @ 3:00 PM 8/25
CMP.B2 (Speedfam CMP-V) - UP/MQC
CMP.B2 24hr MQCS COMPLETED: IN PROCESS
CMP.B2 24hr MQCS NEXT DUE : " "
CMP.B2 12hr MQCS COMPLETED: @ 4:00 AM 8/25
CMP.B2 12hr MQCS NEXT DUE : @ 4:00 PM 8/25
CMP.B3 (Speedfam CMP-V) - : UP/IDLE
CMP.B3 24hr MQCS COMPLETED: @ 5:00 AM 8/25
CMP.B3 24hr MQCS NEXT DUE : @ 5:00 AM 8/26
CMP.B3 12hr MQCS COMPLETED: @ 5:00 AM 8/25
CMP.B3 12hr MQCS NEXT DUE : @ 5:00 PM 8/25
CMP-SCRUB.A1 (On-track scrubber) - UP/IDLE MQC'S PERFORMED
CMP-SCRUB.A1 12hr MQCS COMPLETED: @ 6:00 AM 8/25
CMP-SCRUB.A1 12hr MQCS NEXT DUE : @ 6:00 PM 8/25
CMP-SCRUB.A2 (On-track scrubber) - UP/IDLE MQC'S PERFORMED
CMP-SCRUB.A2 12hr MQCS COMPLETED: @ 6:00 AM 8/25
CMP-SCRUB.A2 12hr MQCS NEXT DUE : @ 6:00 PM 8/25
CMP-SCRUB.A3 (On-track scrubber) - UP/IDLE MQC'S PERFORMED
CMP-SCRUB.A3 12hr MQCS COMPLETED: @ 6:00 AM 8/25
CMP-SCRUB.A3 12hr MQCS NEXT DUE : @ 6:00 AM 8/25
Hot/B lots in area:
__________________
* NONE*
Hot lots coming to area:
-----------------------
* none *
Static Lots in area:
-------------------
* NONE*
Down Equipment:
--------------
* NONE*
Scrap lots:
----------
* NONE*
Hold Lots:
---------
* XD0519 M1 HOLD FOR F.KRUPA, SEE REMARKS
Monitor status:
--------------
5 BOXES of 12k mqc wafers.
Number of product lots in room:
------------------------------
1 LOT
New specs/revs posted:
---------------------
* SEE TDCO BY CMP.B1*
Lots/wafers processed:
---------------------
*XL0525 PMD 24 WAFERS (HOT)
JD0350 M3 20 WAFERS
JD0427 M3 24 WAFERS
# of carriers for strasbaugh/speedfam:
--------------------------------------
12 SPEEDFAM CARRIERS
4 STRASBAUGH CARRIERS
? STRASBAUGH PLATES
|
19.225 | broken wafer on unload flip | FABSIX::O_WHITFORD | | Thu Aug 29 1996 07:35 | 22 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
-< CMP >-
================================================================================
Note 19.222 CMP.B2 EEG PASSDOWNS 222 of 222
FABSIX::O_WHITFORD 21 lines 23-AUG-1996 04:45
--------------------------------------------------------------------------------
Called for broken wafer on unload station.
Unload flipper had tried to place a wafer onto the scrubber station before
the transfer robot had fully lowered out of the way.
Cleaned up wafer fragments & replaced all 6 scrubber brushes due to
visible wafer chunks in them from the crash.... vacuumed out load station,
replaced the secondary pad and removed and scrubbed all load and unload cups.
Re booted the computer and exercised the load/unload sequence with
no further problems.
Dave & Leon
|
19.226 | carrier replacement | FABSIX::O_WHITFORD | | Thu Aug 29 1996 07:38 | 1 |
| Replaced carrier #5 due to poor uniformity.
|
19.227 | Crash after crash | FABSIX::L_SHKOLNIKOV | | Fri Aug 30 1996 08:26 | 12 |
| The system broke the wafers couple times this night.
The first time a wafer slid out from carrier #5 taking #4.
I cleaned up the mess, replaced the pad and 2 carriers.
and checked the load cup position alignment -looks good.
We started MQC and after the first cycle the tool crash
again. The unload flipper dropped the wafer between flipper
box and paddle. Unload flipper alignment has been checked as
well. Claned up the mess again, tool passed MQC and released
to production.
Leon
|
19.228 | CB #6 Replaced/Index Table Infranor Replaced/ Elevator #1 Drain Tube Loose | ASDG::POIRIER | | Fri Aug 30 1996 20:29 | 45 |
|
Breaker #6 Replaced:
-------------------
The Speedfam Rep replaced the 15 amp breaker with a 20 amp breaker.
This was a scheduled upgrade from Speedfam. (Vacuum pump kept tripping
the 15 amp breaker at power up).
Index Table Infranor Replaced:
-----------------------------
The Infranor Motor Controller for the Index Table "blew" during power up.
The Controller was replaced with a new one.
Unload Flipper:
--------------
The Unload Flipper dropped a product wafer while in the home position.
The wafer fell on edge down inside the drop off area. The transfer paddle
moved back to its home position and smashed the the wafer.
The Flipper Arm screws were loose. The black plastic knobs on the
Flipper Arm screws were tightened a 1/4 turn with pliers. They never
seem to remain tight when tightened by hand.
Load Cups:
---------
The load cups were not rising smoothly during the load sequence. All
lift cylinder shafts were cleaned with alcohol and lubricated.
Unload Elevator #1:
------------------
The wafers were occasionally getting stuck half way into the cassette
during unloading. The Speedfam rep indicated that the water level might
be slightly low. The left drain tube was found loose. It looks like the
reducing bushing that the left drain tube connects to is loose. (Bottom
of tank) This loose (left) tube was lower than the right tube, which
was tight. This may explain the low water level in the tank. We didn't
have an opportunity to reset the tube today, maybe D shift can get to
it tonight.
C Shift.
|
19.229 | Slurry delivery system upgrade done | SUBPAC::LANDRY | | Mon Sep 09 1996 20:09 | 41 |
| Today, the system was available for many hours, so I performed the
slurry upgrade to the system. Quite the task. In summary, here's
whats up with this:
* The upgrade was to remove the "peace-pipe" slurry tube going to the
MHTA, change out the MHTA distribution spout, and install new clear/
flexible tubing to the MHTA.
* Upon checking out the schematic, there was a plethora of items to
change out. This included the feedthrough in the polish section
tub. Well, the present feedthrough was welded in-place with the
weld virtually impossible to get at. The print also showed a P-trap
to install in the tub section.
- The feedthrough was not changed out. We decided to use the
existing feedthrough. This necessitated that the original pipe
on the feedthrough had to be modified, big time. The pipe was
cut to a short section and the top was welded to form a seal.
Two holes were drilled and tapped on the pipe, where two fit-
tings were installed.
- The P-trap was not used in this senario, as B3 doesn't have this
feature either.
* The slurry distribution tube on the MHTA was then installed along
with several guides/brackets. Two runs of clear tubing were run
the feedthrough pipe fittings to this new distribution tube.
* Ran a manual polish cycle to test the flow. Seems to be working
just fine. No leaks to be found.
* Please note, the slurry system was only worked on from the polish
tub area and up. No plumbing was changed under the tub or to the
pumps/rinse.
As a side entry, let me say this. The poly welder we have is pretty
useless until we get some self-feeding tips. I borrowed the poly
welder I had purchased while working in the HPA group. Made life a
little easier.
Doc
|
19.230 | paddle recalibrateed | LUDWIG::RIDLON | Eliminate the obvious | Tue Sep 10 1996 06:57 | 8 |
| Noticed that the paddle to track 1 was hitting the left side of the
cut out as it placed a wafer in the track. Checked it at track 2
with the same results. The paddle at home location was also off to the
left although it wasn't hitting the wafer posts. Recalibrated the
paddle to track 1, track 2 and home. Tested fine.
B-shift.
|
19.231 | Broken wafer | FABSIX::R_GEE | | Tue Sep 10 1996 08:38 | 21 |
| Broke a wafer between the unload flipper and the unload paddle
at the beginning of our shift. The wafer came from unload cup #3.
Be believe the unload flipper failed to release the wafer into
the scrub station, when it went to it's home position the wafer
was dropped vertically between the unload flipper and the unload
paddle. When the unload flipper went to place the next wafer
into the scrub station it broke the dropped wafer (??).
Cleaned up the broken wafer. Checked the unload flipper positions,
fingers and lockdown screws, everything looked good. Checked the
unload cups and found them in the wrong index table feedthru's,
placed them in there designated feedthru holes. Checked the unload
air cylinder pedestal, it also looked o.k. Ran 150 wafer monitoring
all the unload with nothing out of the ordinary found.
B-shift
|
19.232 | Load flip initialization failure | FABSIX::L_SHKOLNIKOV | | Thu Sep 12 1996 08:15 | 19 |
| Load flipper failed to initialize.
Investigated problem and tried exercising the load flipper in manual
mode. Regardless of start postion or command, flipper always drove
toward index table and ignored all flags. Swapped cables from unload
flip to the load flip and had the same problem, thus eliminating the
controller as the suspect.
Removed load/unload flip assembly and drained copious amounts of water
from the internal workings. The load and unload home sense electrical
connectors are badly corroded and the rest of the wiring was damp.
Dried it out as best we could and are in the process of replacing the
corroded pins & sockets.
Asembly will need to be aligned and sensors adjusted when tyhe wiring
repair is completed. May also want to check/dry up any remaining water
in the load/unload box.
Dave & Leon
|
19.233 | Slurry Pressure Drop | STRATA::SWENSON | | Sat Sep 14 1996 17:28 | 3 |
| At approximately 1230 B2 slurry loss alarmed. Beeped ashland and return
caller said to contact onsite rep. no response. Slurry returned to
normal 10 minutes later. No product was in process at the time.
|
19.234 | dump valve upgrade | SUBPAC::CMP | | Thu Sep 19 1996 19:42 | 8 |
| installed the positive close feature on both dump valves, all tools now
have this upgrade.
Note: If bubbles are noticed it may be time to rebuild the valve, this
is due to wear on the O rings, but it should be a long time before that
happens.
Rich (Mr. Goodwrench)
|
19.235 | FLEX TRACK/ CABLES | STRATA::RIDLON | Eliminate the obvious | Sun Sep 22 1996 05:14 | 7 |
| CABLE AND FLEX TRACK: FLEX TRACK APPEARED TO HANG UP ON MULTI HEAD .
FLEX TRACK BROKE AND CABLE CONNECTIONS PULLED OUT OF BULKHEAD
CONNECTORS. REPAIRED FLEX TRACK AND TEMPORARY FIX ON CABLES . WILL
OR4DER NEW CABLES OR CONNECTION, WILL LOOK INTO FOR MONDAY
DELIVERY.RELACED PAD.
KEVIN:
|
19.236 | Pad profile | STRATA::KRUPA | | Wed Sep 25 1996 18:09 | 12 |
| Profiled last pad off tool. Pad had 1750 wafers changed on 9/17/96.
location thickness
-------- ---------
ID 61 & 62 mils
1/4 radius 62 & 63 mils
1/2 radius 64 & 64 mils
3/4 radius 65 & 66 mils
OD 66 & 67 mils
Pad is fairly uniform after 1750 wafers. No change in conditioning
made.
-Frank
|
19.237 | pad profile continued | STRATA::KRUPA | | Wed Sep 25 1996 18:13 | 1 |
| Note ID/OD for last entry was set at 9.3".
|
19.238 | HIGH PARTICLES | LUDWIG::WESTPHAL | | Thu Sep 26 1996 20:29 | 9 |
|
CMP.B2 - Down for high particles. Suspected that particles might be coming
from the 2nd table. Changed both pads. Cleaned the entire process
area. The Conditioner was cleaned. Index table drained and wiped
out. One of the drains were also clogged. The same was corrected.
Tool again failed MQCs for particles. Changed the carriers and wiped
out the unload track. Tool presently sets needing MQCs.
|
19.239 | High particles. | FABSIX::L_SHKOLNIKOV | | Fri Sep 27 1996 08:34 | 10 |
| The system failed mqc's for high particles. A new pad installed at
shift change and failed. The tool has been cleaned again and it is
failed. Last time two out of five carriers out of spec. with 1200
particles on them - spec. calls for 1000. Left the tool down for
process engineering decision. Last MQC was run about 7:00 am.
Looks like it' common slurry problem just move from A tools to
Speedfam.
D shift
|
19.240 | grounding | SUBPAC::CMP | | Fri Sep 27 1996 17:08 | 1 |
| The grounding of the mapper and the robot is now completed.
|
19.241 | Wafer break (lost slurry) | FABSIX::R_POIRIER | | Tue Oct 01 1996 06:10 | 31 |
|
SPEEDFAM WAFER CRASH FORM (to be filled out by MFG + EEG + PSS, together)
1) System _ CMP.B2 _
2) Recipe _ 1P_ILD1 _
3) Step # _ 4 _
4) Time into step _ 2 _
5) Type of wafers that broke _ dummy _
6) Carrier which wafer slid out of _ 4 _
7) Other carriers that were damaged _ 5 _
8) Alarms displayed by polisher _ Carrier 4 rotation error
_ Carrier 5 rotation error
9) Did any other polisher error for slurry or crash _yes_, sysytems_A1,A2,B3
10) Did wafer detect work _ yes _
11) # of wafers on pad _ 1160 _
12) # of wafers on carrier _ 277 _
13) Was there any vibration prior to the crash _ yes _
14) Was there plenty of slurry (for step 1-5) or water (step 6-9) present on
the table _ no _
15) Check step hight of the carrier from question #6 .0075 - .008
16) Please note any other unusual circumstances:
After the crash we noticed that there was some slurry on the table,
but not much and there wasn't any slurry in the line. Thats when B3 slipped
a wafer due to low slurry.
We lost slurry for about an hour.
17) Put this check sheet in CMP note #
|
19.242 | WAFER SLIDE OUT | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Oct 02 1996 07:43 | 11 |
|
The tool was running TUP200 on some product and at touchdown,
the wafer under #5 slid out and the wafer detect tripped.
Fortunately it was a 1 sec. touchdown and head #4 was just coming
up. The wafer did suffer some scratching. The carriers and pad
appeared to fine. Both carriers have 91 wafers on them as well
as a recent pad. The crash form at 34.last will give the details.
D shift
|
19.243 | DOWN FORCE/ SLURRY DILIVERY CHANGED | STRATA::WESTPHAL | | Fri Oct 04 1996 20:45 | 8 |
|
CMP.B2 - Scheduled down. Engineering requested the EP regulators be calibrated
and the DF done. The same was completed. The slurry delivery tubing
inside the process area was replaced. A few more items were taken
care of for the weekly PM. Tool sets needing the old carriers put
back on. The pad change and 2 items on the pm completed.
|
19.244 | COMPLETED PAD CHANGE & PM ITEMS | FABSIX::O_WHITFORD | | Sat Oct 05 1996 07:14 | 20 |
| PICKED UP FROM C SHIFT.
RAN MULTIHEAD INTO CARRIER #5, DIDN'T NOTICE THE ALIGNMENT PIN
STILL IN THE SPINDLE.... DOH!!!!!!
CHECKED ALIGNMENT FOR LOAD POSITIONS OF ALL 5 SPINDLES, ALL LOOKED
GOOD.
INSTALLED CARRIERS, REPLACED PAD, AND COMPLETED 2 REMAINING ITEMS
TO FINISH OFF THE PM.
RETURNED TOOL TO PRODUCTION. FAILED FOR UNIFORMITY 2X.
CHECKED DOWNFORCE CALIBRATION, WAS IN SPEC. 267~287 BETWEN ALL 5
CARRIERS. REPLACED ALL 5 CARRIERS AND REPLACED PAD AGAIN WITH ONE FROM
DIFFERENT LOT.
IN MQC AT 06:15
D SHIFT
|
19.245 | Vacuum error came and went ????? | FABSIX::J_SWENSON | | Sat Oct 05 1996 20:28 | 16 |
|
Carrier # 3 wold no come in below 10%.
Changed out carrier #3 and ran MQC's. Carrier #3 still out.
swithched carriers 2 and 3 to verify and MQC"S passed.
head 2 @ 10%
Tool was then put in wet idle for lunch. After lunch the tool
during the 3rd segment of ild3 came up with vacuum error.
(carrier 1-4 no wafer carrier5 o.k.)
Verified I/O's o.k. manually actuated carriers 1-5 vacuum
sequence. o.k. Ran carriers 1-5 to isolate. O.K. Checked main
vacuum guage and witnessed no pressure increase.
After verifying that every thing was o.k. the tool no longer
had an error. Ran ten verification wafers no touch process and
the Tool did not error. No root cause known, left down to pass
on to next shift, for more verification .
|
19.246 | vacuum fix | SUBPAC::CMP | | Sun Oct 06 1996 01:40 | 5 |
| Vacuum error on head 3, found the main vacuum a little low at 23,
checked at the head with loaded wafers and found in the 16 - 17
range. removed the filter and the main went back to 25 - 26 and
the head to 20 - 21. installed a new filter and plastic cup due
to damage of seal surface.
|
19.247 | more on vacuum errors | STRATA::RIDLON | Eliminate the oblivious | Sun Oct 06 1996 02:02 | 11 |
| Came in with b2 being down for vacuum errors. Ran no-touch and found
that head 3 would drop off first (digin 58). Replaced carrier 3 but had
the same results. At about this time Rich called in (C-shift must have
paged him) and asked what was up. With the quick carrier change not
fixing the problem coupled with A1 and A2 being down I asked him to
come in. Figures it was the stupid vacuum filter being clogged. Oh
well. Thanks to Rich for coming in. Maybe a vacuum filter change should
be added to the weekly pm???
B-shift
|
19.248 | Wafer detect tripped | FABSIX::R_GEE | | Sun Oct 06 1996 20:43 | 11 |
|
The system errored for a wafer detected on the table, the
carriers went up and left all 5 wafers on the table. Found
the slurry lines had caught the DE spray nozzle that sprays
the side of the polish table and pointed it upward. When the
side spray turned on it hit #5 wafer detect sensor and activated
it, re-adjusted the nozzle and slurry lines.
Doc & Bob
|
19.249 | Pad profile 1470 wafers | STRATA::KRUPA | | Fri Oct 11 1996 12:00 | 10 |
| Pad was profiles after 1470 wafers on CMP.B2. Pad_step was measured
to be 67 mils. Pad profile was:
location thickness
id 68
1/4 radius 70
1/2 radius 71
3/4 radius 71
od 72
Pad conditioner setting left at 9.3". No changes made.
-Frank
|
19.250 | bad board | SUBPAC::CMP | | Fri Oct 11 1996 19:52 | 5 |
| During a PM the tool lost the wafer detect board. ordered a new one
from stock and replaced.
Rich (Mr. Goodwrench)
|
19.251 | Removal rate | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Oct 16 1996 07:56 | 7 |
|
Tool removal rate was getting down around 900A, replaced
pad.
D shift
|
19.252 | particles | STRATA::RIDLON | Eliminate the oblivious | Tue Oct 22 1996 06:20 | 8 |
| After Ashlands work on the slurry system the tool failed particles X2.
Point of use filters on b3 and a1 failed (caused crashes). The mqc
recipe was modified (added 7 seconds to step 8 and 3 seconds in step 9)
but particles still failed. Running out of options and operators (they
have a meeting at 5) and mqc wafers.
B-shift
|
19.253 | NO DISPLAY | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Oct 31 1996 04:49 | 8 |
|
After the tool unloaded its last (5) wafers, the
screen went blank. I went to the monitor in back and the
display was blank as well. A softboot wouldn't bring anything
back as well. After a hard boot, everything was fine.
D shift
|
19.254 | PADS/CARRIERS CHANGED | LUDWIG::WESTPHAL | | Fri Nov 01 1996 19:37 | 4 |
| CMP.B3 - While down at external both pads and all carriers were changed
out due to high wafer counts on all of them. The tool was also cleaned
up and the conditioner cleaned.
|
19.255 | wafer crash. | FABSIX::L_SHKOLNIKOV | | Sat Nov 02 1996 07:44 | 8 |
| The tool crashed a monitor wafer tonight. The wafer slide out from
carr#5 and crush on#4 - 10 sec.in segment 3. It was a 3-d slide out
from this carrier. First two times it didn't crash. So we checked the
carrier#5 to load cup alignment and it was off. We moved this carrier
on 384 steps out. The tool runs fine after that and it running MQC's
now.
D shift
|
19.256 | doing fine | LUDWIG::RIDLON | Eliminate the oblivious | Sun Nov 03 1996 05:04 | 2 |
| Passed on from C-shift as having low removal rates. Ops did MQC's and
they passed.
|
19.257 | Wafer break (Double Load) | FABSIX::R_POIRIER | | Sun Nov 03 1996 07:13 | 15 |
|
B2 was ready to run some send aheads. It ran a set (5)
of dummy wafers first. When unloading the dummies,
carrier 5 did not unload it's wafer. Tool then loaded
the send aheads and tried to polish. AS soon as the
double loaded carrier hit the pad the send ahead came
out, taking out carrier 4 also. Replaced 4 and 5 and
manually loaded and unloaded wafers with new carrier.
No problem found. Also replaced the pad. Conditioned
new pad and returned tool to Operations.
B shift
|
19.258 | all carriers not down | FABSIX::R_GEE | | Sun Nov 03 1996 19:26 | 14 |
| The system errored for "all heads no arriving down". When
we got to the system carrier #1 was up and all the others
were rotating on the table. It was froze in segment #9 of
the TUP_B process, we had to stop the run to halt the polish.
Unloaded the wafers in manual, homed the multi-head and index
table.
Tried to run some dummies to duplicate the problem but we
came up with a "low main vacuum", found the vacuum was on to
all 5 heads. Checked the Wafer Tracking screen, it still
thought there were wafers on the carriers. Had to reset
the computer to recover, ran over 100 wafers after that
with no errors.
A shift
|
19.259 | I/O cables, filter elements | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Nov 08 1996 07:05 | 11 |
|
Whilest completing the monthly PM, we noticed
that the I/O cables going from the bulk-head to the
MHTA were getting worse. We told Kevin C. about it and
he said they were on order. These cables will not last
much longer. Also, could someone on days look into getting
more filter elements for inside of the vacuum pumps. All
(4) are flaking and falling apart.
D shift
|
19.260 | Wafer loss board out to lunch! | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Nov 09 1996 06:30 | 17 |
|
In the middle of the lot, the tool faulted for INVERTER ERROR.
After that, the tool refused to power back up. When it finally
did, it kept erroring for the INVERTER. We did a little tracing
and found K2 wasn't under voltage and there was no LED display on
the wafer loss breakout board. We jumped out the detection system
and the tool powered up and initialized fine. We checked stock on the
tube as well as manually and there was none. I paged Kevin and he
confirmed there was none stock since the last one was used last month
on B2 as well. He ordered (3) boards with at least one of them coming
in ASAP of the next flight out. He will be in here today to pick it
up and install it. The tool is up and running. We are just doing the
the old method of standing by the ye 'ol EMO. There is a WARNING on
the tool.
D shift
|
19.261 | Wafer Loss Detection Board Replacement | ASDG::POIRIER | | Sat Nov 09 1996 18:47 | 11 |
|
Wafer Loss Detection Board:
--------------------------
The Wafer Loss Detection Board failed last night on "D" shift. Kevin
from Speedfam called this afternoon, he's planning to pickup the
board at Logan around 6:00 p.m. He expects to make it here by
the 8:00 p.m. passdown meeting. He'll be replacing the the board
and reactivating the wafer detect circuit with "B" shift tonight.
C Shift.
|
19.262 | Wafer Loss Detection board | FABSIX::R_POIRIER | | Sun Nov 10 1996 20:57 | 5 |
| Kevin came in Sat. Nite 8:30 and replaced Wafer loss detection board.
Returned tool to OPs.
B Shift
|
19.263 | lock up | FABSIX::B_FINN | | Mon Nov 11 1996 19:07 | 9 |
|
B2 locked up between runs. Pad condition was over, but not on
polish table, table spinning, 5 polished wafer were in unload cups
and 5 new wafers in carriers, carriers were home. No visual alarms,
no alarms on screen just audable. Tried to exit but screen was locked.
AOk after reboot. Ran system to test aok
A shift
|
19.264 | Head #5 load position drifting | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Nov 15 1996 06:20 | 10 |
|
A product wafer slid out head #5 at touchdown. An
alignment was done of 35 and it was out 128 steps. This
is confirmed that it is drifting and will need TLC by SF
ASAP and E.E. will be aware in the A.M.
damn I love abbreviations late in the morning!
D shift
|
19.265 | head 5 alignment repair | SUBPAC::CMP | | Fri Nov 15 1996 12:50 | 5 |
| Today an infranor was replaced for the problem that head 5 had lost
it's alignment several times over the past several weeks.
We will monitor for the same problem and if it develops again
change the dcx board. The last course of action would
be to replace the motor.
|
19.266 | Head #5 Carrier to Cup Alignment Drift Problem | ASDG::POIRIER | | Fri Nov 15 1996 19:38 | 20 |
|
Head #5 Carrier to Cup Alignment Drift Problem:
----------------------------------------------
The Infranor motor controller for head #5 oscillation was replaced and
set to 13 volts today. The oscillation arm still shakes slightly as it
moves in and out. We cycled the arm and the alignment remained good.
We then replaced the DCX board for oscillation and tested the arm
for 20 min running the OSCGO program. Although the arm made a squealing
noise during the oscgo program, the alignment checked out AOK after the
test.
The system was put back UP per Tim Dyer's action plan. Maybe head #5
needs to be simulated under load when polishing before the alignment
drift occurs. Time will tell, keep a close watch on it. We will need to
address the squealing noise at a future time.
C Shift.
|
19.267 | Head #5 Oscillation Drift / Action Plan | ASDG::POIRIER | | Sat Nov 16 1996 20:00 | 47 |
|
Head #5 fly outs:
----------------
Wafers were repeatedly flying out of head #5 this morning. We checked
the carrier to cup alignment on head #5 and it was AOK. (Go figure)
We also ensured that the cup height was correct on cup #5.
We ran the OSCGO program to exercise the oscillation on arm #5. The
arm was chattering & squealing on every "IN" stroke. We called Tim D.
and Speedfam about this condition and they agreed that we should
replace the DCX control module (daughter card) for head #5 oscillation.
The board was replaced and the chattering/squealing stopped.
We then attempted to recheck the carrier to cup alignment when the
system failed to initialize for an "inverter communication " error.
The tool was also emiting a burning oder from the computer cabinet.
We found an ic on the wafer loss detection board scorched. We are
attempting to locate a replacement board. There are none in the fab 6
stock room but Tim D. & Rich Barrett are sure there are two in house.
We're handing off the search to "B" shift.
Note: THE WAFER LOSS DETECTION SYSTEM IS CURRENTLY JUMPED OUT!
BE SURE TO REACTIVATE THE SYSTEM WHEN THE NEW BOARD IS INSTALLED.
Please follow the latest action plad per Tim Dyer:
1) Locate one of the missing Wafer Loss Detection Boards and install
it. (If one cannot be located, call Tim Dyer at home.) Do not
run the tool without wafer loss detection.
2) Run 50 wafers for a verification run. If Arm #5 performs normally,
turn the tool over to Mfg for MQC's. If it misloads or throws a
wafer, keep the tool down and go to step 3.
3) Order the Oscillation motor & gear reducer from the Fab6 stock
room and replace them. Repeat verification run.
Speedfam Drawing # AA2265
Item #'s: 33 & 45.
33 = Motor, AC Servo AC2781A1
45 = Reducer, zero backlash AC2782A1
C Shift.
|
19.268 | gotta have connections. | FABSIX::M_RIDLON | Eliminate the oblivious | Sun Nov 17 1996 06:30 | 12 |
| Couldn't find any wafer loss boards in the stock room, sf's office, or
Annie Hargrave's office. We were able to troubleshoot the board from
last weeks failure and as luck would have it we had some replacement
ic's in house. Located a PACE station (for you non-navy pukes a PACE
station is a solder extraction station) and removed/replaced the faulty
ic's (U2 and U4). Powered up and initialized fine. Attempting to follow
through with the game plan now, however, from the looks and sound of
arm 5 we'll be skipping the 50 wafer test and going for the motor/gear
replacement.
B-shift
|
19.269 | arm 5 | STRATA::RIDLON | Eliminate the oblivious | Sun Nov 17 1996 07:28 | 9 |
| Took the covers off of arm 5 (brain fart) which turned out to be
fortunate because we found that a set screw for the gear on the
rotation motor had come loose and the gear had fallen down. The other
set screw was in a shaft key way which allowed the carrier to still
rotate. Anyway the oscillation is still shakey and sounds lousy. We'll
pass it on to A-shift as the time is getting late.
B-shift
|
19.270 | motor problem, head 5 | SUBPAC::LANDRY | | Sun Nov 17 1996 20:14 | 25 |
|
The system is currently down for head 5 motor problems. We have just
drawn a motor and zero reducer backlash from stock. There are in the
room. We had many problems with this system today while trouble
shooting. Head three bound up on us in the "home" position. We could
not drive it off by using DCX commands. We had to physically crank the
motor to get it to where it would move. After we moved it off, we tried
to home the heads. When it was time for 3 to home, it drove the motor the
opposite position. ??? So, we tried driving the head in DCX and it still
responded backwards. ??? We ended up changing the motor as the Infranor
responded as should. The new motor fixed this fault.
Head 5 oscillation was checked eventually. It seemed to have a big differ-
ence between running in manual in/out as apposed to running the OSCGO
program. OSCGO ran much smoother, with some shimmering. Manual in/out
was very rough.
While trying various things to isolate the problem, the Infranor for head 5
was lighting up on all three LED's, meaning overtemp. Changed the Infranor,
no change. Swapped the daughter board, no changed. Ohm'd out the cable, all
is fine there. There is 3.9 volts from motor 7 resolver shield to TC pins
29 & 30. As compaired to other TC's, which were at zero (o). Field service
recommends changing the motor out.
A shift
|
19.271 | jerks are not us | STRATA::RIDLON | Eliminate the oblivious | Mon Nov 18 1996 06:20 | 16 |
| Removed the motor and gear reducer for head 5. INstalled the new ones
and checked oscillation and the jerking is gone. Cup to carrier
alignment was right on as well with no adjustments needed. New pads and
carriers installed and broke in. No wafer fly outs from head 5 or
others. Everything was going fine until the 4th set of the first 25
wafers. Thats when the polish table decided to stop rotating. Rich
called in about then to check up on things and gave us some ideas to
check. One of them was to jump out the wafer detect board. EVen with
both green leds lit the relay for the 24 volt interlock circuit can
still not be made which will prevent the table rotation. Sure enough
thats what the problem was. We pulled the board out to trouble shoot
and it started working fine. We decided to try again to see if its just
a brain fart. More later.
B-shift
|
19.272 | table stall / due to inverter temp | STRATA::FINN | | Mon Nov 18 1996 19:19 | 14 |
|
Plagued by random inverter errors. Errors occurred both in automatic and
manual modes. Occurred with wafer detect board active and jumpered out.
Also happened with/out a load (carriers) on the table. Problem was traced
to two non operative muffin fans on the inverter. These were changed by F.S.
Ran 70 wafers without rotation errors or wafer flyouts. Gave to
manufacturing for MQC's. It passed both particles and uniformity. System
is up and idle.
A shift
Ran 70 wafers
|
19.273 | inverter fan fix | SUBPAC::CMP | | Tue Nov 19 1996 10:13 | 6 |
| Monday morning a problem with inverter developed. During a polish the
motor would stop and an error was displayed on the LCD screen.
The error could be cleared by a power off and the motor would run
for a short time before it happened again. We found that the cooling
fans had stopped working, so we replaced them and brought the tool
back up. The polish table is now working fine.
|
19.274 | unoformity problem | FABSIX::O_WHITFORD | Quticherwhinin' | Wed Nov 27 1996 07:33 | 3 |
| After PM, had uniformity problems with carrier #1 (449 wafers) Replaced
carrier #1 at ops request, return to them for verification.
|
19.275 | Index table vibration | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Nov 30 1996 06:18 | 10 |
|
There was a massive vibration coming from the
index table after the 1st wafer was loaded and the
table was rotating to LOAD #2. The rear process
doors were opened and causing the tool to start and
stop. Found the air line that feeds the door lock
pistons was pulled off on the left inside door. After
the line was re-attached, the tool ran smooth.
D shift
|
19.276 | strange event (CPU ?) | FABSIX::B_FINN | | Sun Dec 01 1996 19:30 | 28 |
|
System was passed down to us with a report of the following
problem "multi head would drive to polish table extend carrier arms
and ramp carriers, but would not descend to pad in segment 3. No
errors would occur until screen said timeout waiting for carriers NOT
on pad" sic. System had been power failed and recipe changed to
io_breakin by Roger and it still failed the same way.
We changed recipe to io_mqc and it ran fine, checked both
recipes for errors found none swapped back to io_breakin and system
ran fine. Ran 20 more wafers for 30 sec polish, aok. FM
Later in the day the operator removed system from wet mode to
check the pad count. The system did not decelerate at polish table, big
bang, EMO was hit before multi head reached carrier rebuild room. Doc rebooted
aok, and placed in wet mode, then took it out of wet mode to check pad
count and screen locked up, rebooted and upon reboot received error message
"elodev file not found" Check monitors in front and back(a/b) modes, failed
both ways. Checked connection at a/b box and switch positions. Checked for
monitor cable errors. Found neither "red fail" light nor "green pass"
light were illuminated on the CPU board. Green would be best but
either is expected.
Powered down reseated CPU. System ran fine. Ran test wafers aok.
A shift
|
19.277 | BROKEN BELT | FABSIX::B_WESTPHAL | | Wed Dec 11 1996 18:32 | 4 |
| CMP.B2 - Down for scrubber not operating properly. A complaint of the wafer
not spinning while the scrubber was opertaing. Found that the bottom
brushes were not working due to a broken belt. The same was replaced
and the scubber was tested. The tool was returned to production.
|
19.278 | Crash in Segment 10 (Lift-off) | FABSIX::R_POIRIER | | Tue Dec 24 1996 05:48 | 28 |
|
B2 crashed last nite. After loading first set of break-ins
Multi-head moved out to polish table. It alarmed for "No
Wafers on Carriers 1-4. Wafers were on carriers, so we hit
continue. Alarm went away and carriers started polishing.
In the 10th segment carrier 5 did not lift off the pad.
All other carriers lifted off. The table kept turning, and
there was still no alarm. At some point the wafer was crushed.
White ring on carrier 5 polished almost down to the film.
Alarm finally came for "Main Vacuum Low". No wafer detect
alarm ever came on.
Carrier 5 was cooked and carrier 4 was replaced because it
had chips all around white retaining ring.
Changed pad and carriers 4 and 5. Tested carrier 5 in Manual
and it works fine. Also checked Wafer detect and it works ok.
Put tool in verify and gave it back to Operations. They ran
boat of break-ins and no problems noted. Tool is ready for MQC's.
B Shift
|
19.279 | Misload on head #5. - Rebooted System | ASDG::POIRIER | | Thu Dec 26 1996 19:26 | 13 |
|
Misload on head #5:
------------------
The tool misloaded a wafer on head #5. The system was rebooted, and a
no touch recipe was run. Head #5 loaded properly. The carrier to cup
alignment checked out AOK. The index table rotation also checked AOK.
A wafer misloaded last night on D Shift as well. Rebooting the system
eliminated the problem both times. (Software?) Keep a close watch for
this reoccuring problem.
C Shift.
|
19.280 | unload flipper problem | SUBPAC::LANDRY | | Sun Dec 29 1996 19:46 | 19 |
| Had a problem with the uload flipper today. The problem started when
running the last run of a lot. The Multihead banged a few times at
the the polish end (after polishing). The system was softbooted to
home everything. We manually tried to unload 5 wafers. It did the
first wafer fine, but the flipper never returned to home after placing
the wafer in the scrub station. The flipper only returned about 75%
of the way back. Tried homing the flipper, and all worked, until we
tried to unload the remainder of the wafers. At this point, the problem
re-appeared. Talked with Dave Rocha and Kevin Coleman. They both saw
this problem late in the week. The found that we have to hardboot the
system, go directly into unload 5, and the system works fine. Tried
their advice, and it worked. No further unloading problems happened
the remainder of the shift.
In the middle of this activity, also tried working the fix through the
calibration screen. It exhibited the same problems with homing as pre-
booting did.
A shift
|
19.281 | lots o-stuff | FABSIX::M_RIDLON | Eliminate the oblivious | Mon Dec 30 1996 07:04 | 27 |
| Had another problem with the unload flipper this evening. While trying
to correct the multihead lost its postioning. AFter soft booting, bkin
wafers were ran but the first 5 crashed and took out all 5 heads.
Recovered and mqc's failed twice for carrier 5 uniformity. Since we've
been having some weird problems with this tool of late we decided to do
the weekly to see if we could find anything out of the ordinary. We
found DI/O cables 7 and 8 to be almost pulled out their connections on
the right side bulkhead fittings. We took apart the flex track and
repositioned all 8 cables. DI/O calbes 5 through 8 have places in the
outer jacket that have been ripped through the inner wire insulators
are fine. We ohmed out the 4 cables to make sure that no shorts or
openes exsisted. After reinstalling the cables and retierapping (these
cables were supposed to be changed due to the connectors being torn off
about 3 months ago) we tried to bring the tool up but it would not
reboot due to "elodev" not being installed. We called Dave Beecher for
some help (local DOS guru) and found that the cpu wasn't seeing the
touch screen controller. This was finally isolated to COM2 on the cpu
being faulty (I forgot to mention that we experienced a power glitch
while trying to boot up which probably took out COM2). We switched the
touch screen control cable to the COM1 port on the cpu and everything
booted up fine. NOTE: The original AUTOEXEC.BAT file was copied to
AUTOEXEC.DJB and AUTOEXEC.BAT was reconfigured to go through COM1.
I thinks that everything.....
B-shift w/ Dave Beecher
|
19.282 | wafer loss board and other stuff | FABSIX::M_RIDLON | Eliminate the oblivious | Tue Dec 31 1996 07:19 | 20 |
| System alarmed with no error message displayed on screen. Both Ops
and EE personnel observed carrier 5 stuck to the pad while the others
were up. The table was spinning at the desired speed but the carriers
were not rotating. EE's first thought for an explanation was that
the wafer loss detect board didn't operate properly. This was tested
using the same recipe and 4 dummies (carrier 5 was trashed). While the
carriers were on the pad a wafer was manually placed under the wafer
loss sensor. The wafer loss board led lit up for the sensor being
tripped but the wafer loss detect led didn't light up. The table and
carriers kept spinning and stayed on the table. With this in mind our
next thought was that maybe the wafer came out in step 3 while vacuum
was still on. The alarm could have been main vac low which may cause
the carriers to lift off the pad and possibly stop the slurry pumps.
Which would explain all observations seen.
The wafer loss board was replaced and tested good. No pad or carriers
were put on due to the upcoming shutdown.
B-shift
|
19.283 | flipper finger and CPU | SUBPAC::CMP | | Tue Dec 31 1996 14:37 | 4 |
| Replaced the unload flipper finger that was damaged last week.
Replaced the CPU and found a file that was damaged, repaired the file
and put back old CPU (no problem found) com port 2 works fine.
Reboot tool and test.
|
19.284 | Touch screen on the blitz again | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Jan 02 1997 08:30 | 8 |
|
The touch screen controller wouldn't install. Checked
a few things but nothing seemed wrong. Swapped over to
COM1 as the weekend guys did and it was now ok. Reps
notified.... The original AUTOEXEC.BAT is now backed up
and called AUTOEXEC.DJB
|
19.285 | vacuum problem | FABSIX::J_SWENSON | | Thu Jan 02 1997 17:53 | 29 |
| loading problem for carrier five
--------------------------------
- wafer not seating properly on carrier five
- checked go-no-go alignments. aok
- checked cup height alignment. aok
- found dried slurry in the mist eliminator for vacuum.
removed filtering unit and ran test wafers.
no error ocurred during the test.
- reinstalled filtering unit and started another test run.
tool failed during first five wafers.
- removed filtering unit ran a boat of wafers with no errors ocurring.
- returned tool to production .
1. The filter unit and guage need thier locations swapped.
The filter can be clogged and still read great vacuum.
C-Shift
|
19.286 | Pad change, removal rates going down again | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Jan 03 1997 04:56 | 6 |
|
Pad change at 180 wafers due to low removal rate.
D shift
|
19.287 | Wafer Breakage / Action Plan For Today | ASDG::POIRIER | | Fri Jan 03 1997 19:28 | 58 |
|
Wafer Breakage:
--------------
Eight product wafers were scrapped due to wafer breakage on the second
table. It appears that the wafer in head #5 slightly unseated from the
carrier during lift off on the primary table. (Not enough vac loss for
alarm) The wafer in head #5 broke apon contact with the secondary table.
A small piece of this wafer scratched the other 4 wafers. (The broken
wafer was cracked in half with a small piece broken off the edge)
The tool then unloaded 5 wafers and picked up another five. These five
were polished then scratched on the second table. This incident became
noticible when the unload flipper attempted to unload the broken wafer.
(Unload Flipper alarm)
Tim Dyer has published the following action plan for tonight:
-----------------------------------------------------------------------
CMP.B1,2,3 - OEM vacuum filters to be installed, and gauge put on BOTH sides
of element for DP readings. The filters from SR-94 have changed,
for reasons unknown and are clogging in 1-2 days.
CMP.B2 - 8 wafers from lot xd0590 broke/scratched on 2nd table. The ONLY
time we have experienced crashes on 2nd table is when an out of
pocket condition occurs during the lift off from the 1st table.
Test this theory by stopping the polisher immediately after lift
off from the primary table. Inspect all wafers for out of pocket
conditions. If they are in fact out, the vacuum must be improved or
the surface tension reduced.
DEC and SF reps will work to isolate a potential low vacuum problem.
The collapsing tubing in the multihead will be replaced, the vacuum
filter replaced and the pump rebuilt. This should improve the
vacuum at the carriers to ~25". If vacuum is in spec and wafers
continue to come out of pocket, change the last two step of the
recipes to 12.9 (from 12.7) and retest. If wafers remain in pocket,
check MQCs. If MQCs pass run marathon.
INCLUDING BREAKIN AND MQC WAFERS, CMP.B2 IS TO RUN A 100 WAFER MARATHON
BEFORE RELEASING TO MFG. THIS SYSTEM HAS BROKEN 12 PRODUCT WAFERS IN 4 DAYS
ANY QUESTIONS, GIVE ME A CALL.
------------------------------------------------------------------------------
Status at Shift Change:
-----------------------
1) Change OEM Vaccum Filters on All "B" tools. - Not Done (New Filters
in CMP room)
2) Stop tool after lift off from primary table and test for "wafer
out of pocket" condition (B2). - Ongoing
3) Replace thin walled tubing in multihead on B2. - Done
4) Replace Vacuum Pump on B2. - Ongoing (New pump on floor next to B2)
C Shift.
FYI: Richard Barrett from Speedfam will remain on site until 9 P.M. to
assist D Shift.
|
19.288 | test in progress (AOK) | SUBPAC::CMP | | Fri Jan 03 1997 20:59 | 10 |
| The test is running with good strong vacuum, on average 23 - 25 hg
during pickup and lift off. The wafers have a good fit to the carrier
when checked before touch down. There is however a loud noise at
lift off possibly due to very strong surface tension. The marathon
will continue with Bkin3 being used with bkin and MQC recipes to
finish off the run.
At this time the tubing on the multi head has been changed to the
heavier style which may account for the improved vacuum. The unload
flipper had to be adjusted today due to poor pickup.
|
19.289 | Vaccum issues, marathon ran ok. | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Jan 04 1997 06:55 | 20 |
|
Tool had already run 40 wafers ok on the tool to
verify the vacuum line replacement in the flex track.
There was great vacuum after lift-off and at load.
Two of the action items Tim wanted done weren't done
prior to starting the marathon so we stopped the
run and installed the vacuum filter and another vacuum
gauge. While installing the 2nd gauge, we found the
electric solenoid in-line with the liquid separator
had excessive slurry build-up in so we replaced that
as well. We didn't have a rebuild kit for the pump so
we replaced it with the pump that was next to the tool.
We continued on with the marathon run and it completed
fine without any issues. all loads and lift-offs held at
24.5 to 25 in/hg on all 20 runs. MQC's were run and passed
on particles and we are just waiting for the range and
deviation results to come in.
D shift
|
19.290 | vac check valve | LUDWIG::FINN | | Sun Jan 05 1997 19:30 | 11 |
|
Main vacuum error intermittently. Vacuum was good 28" and 26" on
either side of filter, but pumping speed was slow. Removed water from filter
glass trap. Dried filter. Did a manual flush of tank,(vacuum reservoir was
filled with water). Pulled check valve and cleaned. Found two pieces
of Teflon tape across the seat of the check valve. Cleaned rinsed and
reinstalled. Took pump apart and dried seals and diaphrams. Reassembled and
tested. AOK
A shift
|
19.291 | lock up | FABSIX::M_RIDLON | Eliminate the oblivious | Mon Jan 06 1997 06:36 | 17 |
| Operator found tool alarming whilst processing. Observations made were;
MHTA was over the index table, the pad conditioner was over (but not on)
the polish table, the table was rotating, the transfer paddle at track
1, no wafer was blocking the track sensors, the mhta had unloaded and
loaded fine, main vac was good, screen had locked up. All DIN's were
checked and nothing out of the ordinary was found. Those that should be
lit were lit and those that would indicate a problem were not lit.
EMO'd tool to recover since the screen locked up. While unloading the
wafers manually from the index table 1 broke. It broke real easily as
no excessive pressure was used and the fracture was straight across the
wafer starting from the notch. May have been a chip in the edge. PSS
was informed. Since the transfer paddle remained at track 1 while the
mhta unloaded and loaded, 5 wafers were ran with the "NOT_TOUCH" recipe
to insure proper operation of the unload module.
B-shift
|
19.292 | Particles.... | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Jan 08 1997 07:16 | 10 |
|
Tool failed MQC's for particles. We ended up
taking the tool for the Monthly PM and a thorough
cleaning. The PM was completed was turned over to
OPS. The particles passed but the range was slightly
out. OPS are currently trying to bring it in.
D shift
|
19.293 | flex track cable replacement | SUBPAC::CMP | | Thu Jan 09 1997 17:58 | 3 |
| Replaced the 2 cables from the bulkhead connector through the flex
track to the multi head. During initialization we lost the Infranor
for rotation 4, replaced and tested tool.
|
19.294 | cable replacement | SUBPAC::CMP | | Fri Jan 10 1997 15:58 | 2 |
| Installed cables 5 & 6 in the flex track to the multi head.
This completes the cable repair/replacement.
|
19.295 | check cassette | FABSIX::M_RIDLON | Eliminate the oblivious | Tue Jan 14 1997 07:42 | 7 |
| Operator found tool not running when it should have been. Isolated to a
broken cassette in elevator 1 (it wasn't tripping the sensor which
stopped the process once the wafers were done polishing). Replaced the
cassette and resumed.
B-shift
|
19.296 | chiller setpoint | LUDWIG::FINN | | Wed Jan 15 1997 19:50 | 10 |
|
Flow and temp alarms, for B2 Chiller. Found heater ON! And fan tripped.
Reset fan and turned heater off. Unit controlled down to 28'c but we found we
could not adjust pot for set point. Check set screws on cap aok removed and
tried to controlled by pot shaft still no control, Check wiring as best we
could, (unit is installed close to wall and cover does not open all the way.
Omhed out put it is adjusting. Need to find schematics. Unit is at proper
set point, but we can't adjust up or down.
|
19.297 | Chiller wouldn't operate. | FABSIX::L_SHKOLNIKOV | | Thu Jan 16 1997 07:50 | 7 |
| The tool went down for the chiller alarm. We could not reset the
chiller because on/off switch is totally destroyed and should be replaced
with a new one. Recirculation pressure gauge on chiller shows 0.
Temperature at the time being is 9'c - well below alarm setpoint and
can't be adjusted.
Leon.
|
19.298 | chiller fix | SUBPAC::CMP | | Thu Jan 16 1997 11:33 | 5 |
| Arrived and found the chiller idle. Put the chiller in manual mode and
disconnected from tool, adjusted the set point (chiller must be in
manual mode for this adjustment), powered the chiller off and
connected the cables, powered on the chiller, put in remote mode and
monitored. No problem was noticed and tool was brought back up.
|
19.299 | Chiller issue | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Jan 16 1997 17:24 | 15 |
|
The chiller wasn't working this morning. The low
alarm setting was at 8C and the SET POINT was set below
that at 3C. Thus... it will always alarm and not flow.
To change the SET POINT, in back of the chiller, set the
slide switch from REMOTE to MANUAL, set your SET POINT
at 24C and the slide the switch back to REMOTE. The REMOTE
is for the SPEEDFAM tie-in. The on/off/start switch is
totally junk. I've contacted Neslab and I getting a spare
parts listing sent to me so I can get some things for it.
BJ
|
19.300 | Conditioning Wheel Changed | ASDG::POIRIER | | Thu Jan 16 1997 18:55 | 9 |
|
Conditioner Wheel Replaced:
--------------------------
A new conditioning Wheel Was installed on the tool. The new diamond
wheel I.D.# is 218. The Total wafer count on the tool at the change
over was 60,477.
C Shift.
|
19.301 | Unload flipper issue | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Jan 31 1997 07:48 | 17 |
|
2nd ASSIST in 12 hours for two wafers in the scrubber.
Took tool down and watched 10 wafers unload. Noticed that
the unload flipper wasn't opening and closing smooth. The
cover to the unload flipper was removed and everything looked
ok. We gave the bimba a slight lubing and the fingers open
and close fine now. While running more VERIFY wafers, we
wanted to know why the beam sensor alarm didn't trip when
a wafer didn't break the beam. We pulled a wafer off of the
paddle before it go to wafer track #1 and the tool tried
sending nothing to elevator #1. It should of timed out and
alarmed it didn't. It just waited about 15 seconds and then
went for the next wafer. This will be brought to SF attention.
D shift
|
19.302 | flipper checks | SUBPAC::CMP | | Fri Jan 31 1997 10:02 | 3 |
| Checked operation of unload flipper, and found that it's OK. Checked
the operation of water track sensors and they checked out OK. Returning
tool to ops.
|
19.303 | UNLOAD flipper problems again | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Feb 01 1997 06:26 | 22 |
|
The unload flipper issue came right at us around 9pm.
From Dave (SF), if there handling issues again, go for
the bimba. After we replaced the bimba cylinder, we
were checking to make sure everything was ok when the
unload flipper started screwing up. When asked to HOME
coming from the UNLOAD cup, it would home fine. But when
the flipper was told to go HOME from the SCRUBBER, it would
find a HOME position about an 1" before where it should of
found it. The flipper box was removed and inspected. It was
discovered that home sensor was setting home on a set screw
on the cam. The home sensor amplifier was adjusted to not
pick-up the screw. The wafer present sensor for the
UNLOAD flipper wasn't always seeing a wafer. That sensor
was adjusted as well. The actually might have been part
of the problem all along. Everything was put back together
and verified. The tool loaded and unloaded wafers fine.
The tool came back up and has been running great since
without any problems.
D shift
|
19.304 | computer lock up | SUBPAC::LANDRY | | Sun Feb 02 1997 19:59 | 17 |
| The system's computer locked up today. The alarm sounded for an error in
the Unload area of the tool, but it was uncertain as to the exact termin-
ology, as the error vanished from view soon after the alarm. The only
indication to the Unload area being the problem was that a wafer was
found angled in a position such that the paddle wound have smashed it
should it have tried to transport it to the water tracks. Anyway, we
tried to resume running after repositioning the wafer onto the paddle,
but the computer had no response. The conditioning arm was still out
over the Polish Table, with the table spinning at conditioning speed.
We then softbooted, skipped initialization, exited to DOS and powered
down the tool. Waited, then brought the tool back up. Checked the
unloading sequence after the reboot. Found the unload grippers to be
below the wafer when trying to unload from the unload cup. Re-calibrated
the flipper at unload cup position. Tested several unload sequences
with positive results.
A shift
|
19.305 | unload cup pedestal crash | STRATA::FINN | | Sun Feb 02 1997 20:12 | 18 |
| While polishing system began loading the next series of 5 wafers.
It appears that a white Guide roller fell off the process door
(while load cup one was up) and rolled under the cup preventing it
from descending all the way to the base of the index table. When the
table indexed it crashed into the unload cup pedestal, bending it.
A second guide roller ( but no screws ) was found under the index
table. We backed the cup off in DCX mode. Now that the roller is
removed the pedestal clears the cup but is bent and does not make the
cup down sensor. Ran out of time, verbal Passdown given to B shift.
Left system powered down during Passdown.
A shift
|
19.306 | unload pedestal/doors modified | FABSIX::M_RIDLON | Eliminate the oblivious | Mon Feb 03 1997 00:09 | 11 |
| Found an unload cup pedestal in stock (3 remaining) and
installed/adjusted. We also took the liberty of modifying the process
doors a bit in the hopes that we can avoid this type of problem until
new doors can be installed. The modification consisted of removing the
two right doors drilling and tapping any existing holes that were
stripped and drilling and tapping two new holes on top and bottom
giving us 5 rollers instead of 3. The right front is a little tight
when moving but the right back moves well.
B-shift.
|
19.307 | Software drag | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed Feb 05 1997 05:56 | 9 |
|
When the OPS tech went to start-up the pad break-in
after the monthly was completed, the software had a brain
fart. The conditioner was conditioning when is wasn't
suppose to, the arms wouldn't move out ect. After a hard
boot, the tool ran smooth for the rest of the morning.
D shift
|
19.308 | Uniformity problems with head #2. | ASDG::POIRIER | | Wed Feb 05 1997 19:24 | 22 |
|
Head #2 Uniformity Out Of Spec:
------------------------------
The system experienced uniformity problems on head #2 today.
1st Attempt: DF offsets were reclaculated.
2nd Attempt: ID/OD offsets were adjusted.
3rd Attempt: Carrier #2 was replaced.
4th Attempt: Down force on all 5 heads were recalibrated.
5th Attempt: (Ongoing) Pad replaced, MQC's to commence.
The pad that was just removed from the tool was put on
during the Monthly P.M. last night. It pulled off with
very little force. The pad count is 105. We suspect the
pad may be the cause of head #2's uniformity problem.
Even though the unformity problem is isolated to head #2,
we're changing the pad after trying everything we can
think of related to head #2.
C Shift.
|
19.309 | head two removal high | FABSIX::J_SWENSON | | Fri Feb 07 1997 19:44 | 14 |
|
Head two only removal rate excessive
____________________________________
This only happens on product wafers.
Ops mqc wafers are right on the money, but when product is run the
removal is off on only head two.
All offsets are very close also.
will pass on to D-shift to check head 2 down force.
C-shift
|
19.310 | GOOD EXCUSE. | FABSIX::L_SHKOLNIKOV | | Sat Feb 08 1997 04:51 | 24 |
| COULDN'T FIX THE TOOL. FOUND NOTHING WRONG WITH IT.
ALL WAFERS WERE RUN FOR 120 SECONDS ON THE BELOW NOTED RECIPES
ILD1 RECIPE WAS THEN DEL AND ILD2 COPIED IN ITS PLACE. TWO LOTS
PROCESSED, ONE ILD1 AND 1 ILD2, BOTH LOTS PROCESSED WITH NO REMOVAL
RATE PROBLEMS FROM ANY ONE CARRIER. BOTH LOTS WERE DEVICE DC1043.
REMOVAL RATES AT DIFFERENT RECIPES
CARRIER MQC RECIPE ILD1 RECIPE ILD2 RECIPE
------- ---------- ----------- -----------
1 3396 2970 3149
2 3328 3018 3182
3 3288 3001 3170
4 3227 3126 3198
5 3314 3065 3200
AVRG 3309 3036 3179
RANGE 169 156 51
D - SHIFT
|
19.311 | unload flipper cal | SUBPAC::LANDRY | | Sun Feb 16 1997 18:09 | 9 |
| Had a couple of "unload flipper timeouts today". Observed it's
functionality and found that the flipper at unload cup height was
mis-adjusted. The flipper was driving into the cup. This caused the
periodic timeout of the flipper (unable to close grippers to make the
sensor). The pick up position was adjusted through the calibration
screen. Ran fine the remainder of the shift.
A Shift
|
19.312 | low slurry flow | FABSIX::R_GEE | | Mon Feb 17 1997 19:34 | 8 |
| Errored for low slurry flow. Found air in the slurry lines.
Called and paged Ashland chemical they said they did change filters
without notifying use, they were told that we must be notified.
Manually pumped slurry and restarted the tool.
A shift
|
19.313 | Vacuum and Infranor issues | FABSIX::B_JUBINVILLE | | Fri Feb 21 1997 06:32 | 18 |
|
Tool alarmed for LOW VACUUM and ARM #4 NOT AT HOME.
Found Infranor for #4 oscillation cooked, replaced same.
Proc. Eng. wanted head #4 and #2 swapped at next MQC
because there was a process concern with #2. After we
swapped them we tried starting tool back up but we kept
having vacuum problems. Thinking it was a head #4 problem,
we disabled head #4 and the vacuum problem went away. We
pulled the carrier off and blocked the hole on the spindle
and verified it was a carrier issue. When I picked up the
carrier (which was #2) to inspect it, (3) out of the (6)
M5? screws fell through the floor. :^) I checked the other
(3) and they were ready to come out as well. After we put
a new carrier on the tool, everything was fine. The new
carrier went on #2 and old #4 went being #4.
D shift
|
19.314 | CMP.B2 wafers missing error | FABSIX::J_SADIN | Freedom isn't free. | Tue Feb 25 1997 10:54 | 13 |
|
Machine alarmed for "wafers missing" at the end of 'b' shift (tool
was emo'd because conditioner would not stop conditioning pad). Found
wafers on only heads 2-5 (no wafer on head 1). Blew off remaining
wafers and checked alignments (all ok). Adjusted load flipper to
aligner position forward 20 steps. Ran 25 test wafers w/o incident.
Returned tool to production.
Also had to reboot chiller since pump had shut off after CMP.B2 was
emo'd.
a-shift
|
19.315 | slurry alarm | FABSIX::O_WHITFORD | Lead, follow or step aside. | Fri Feb 28 1997 06:10 | 13 |
| Had slurry flow alarm. Found a significant "bubble" in slurry #1 lines.
Exercised slurry system #1 and bubble purged itself.
Why B1 & B3 DIDN'T alarm? Beats me. Maybe B2 DID alarm as it was the
"idle" tool at the time. (With the others running, any bubbles would
be there and gone. THIS looked like it got trapped there after a pad
wetting cycle in wet mode....) Makes me wonder if Ashland did anything
to the slurry system or filters at or around 8:00 PM.....
Too came back on line o.k. after purging the lines. Also completed
weekly PM and replace weakened slurry pump hoses tonight during PM.
D shift
|
19.316 | Semi annual and index table | FABSIX::B_FINN | | Tue Mar 04 1997 19:21 | 13 |
|
System went down scheduled for semi annual pm and index
table replacement. All Index Table, carrier and flipper alignments
checked.
PM is well on its way to completion. Most of the lubrication on was
not done due the absence of a grease gun. Incomplete pm check sheet
was mailed out and entered into CMP note # 77.
A shift/ Dave Rocha
|
19.317 | flipper finger adj. | SUBPAC::CMP | | Tue Mar 11 1997 17:15 | 4 |
| wafer slipped out of unload flippers. Checked adjustment and found that
the fingers were gripping wafer and pushing slightly forward, made an
adjustment to the fingers, ran wafers,and found it to run better.
Return tool to operation.
|
19.318 | multi head motor key | STRATA::FINN | | Wed Mar 12 1997 19:41 | 28 |
|
Operator called us off cmp.a2 to look at cmp.b2. Kevin Jubinville
had just hit the EMO. We found the multi head over the polish table with the
conditioner arm crashed into the multi head. Water was on at the clean station
secondary table spraying up to toward the top of the system.
We moved the multi head to load manualy and turned the system back on. The pad
conditioner homed, but over shot and caught the lip of the condtioner bird
bath. Multi head was also free wheeling (not moving) and could be moved by
hand.
Checked infranor and breakers. aok. Found key had fallen out of the multihead
shaft/pully. Had to remove (drill out rivits) exhaust and back pannel to gain
acess to key/pully. Reinstalled key, tightend compression fixture. Tried to
power up to check multi head travel but system would not power up. Found one
phase missing at fuse at T1. Replaced fused. aok. Reinstalled all hardware.
Adjusted conditioner position sensors. Pulled carrier 3 and inspected aok.
Ran test wafers aok, returned to manufacturing. Great team effort.
A shift/Speedfam
|
19.319 | Unload flipper box problems | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Thu Mar 13 1997 06:00 | 20 |
|
The tool went down at the beginning of the shift
for wafers breaking in the scrub station. After the
mess was cleaned up, we ran wafers to verify the handling.
Everything ran smooth until the tool was unloading 21-25.
The unload flipped had the wafer but it said it missed it.
Acknowledging it would just keep the tool running and the
flipper would go to get another wafer with a wafer already
in the fingers. After visual inspection of the unload
flipper box guts, the arm that swings over the proximity
sensors which tells the tool the status of the fingers was
binding up and not always making the sensors. After everything
was moving smooth, the amplifiers for the CLOSE and WAFER
PRESENT position needed to be tweaked. (25) wafers were
run after that and it ran great. OPS ran full MQC's on
the tool and everything passed fine. The tool came back
up and was running product.
D shift
|
19.320 | Chiller controller board | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Sat Mar 15 1997 07:24 | 13 |
|
Found another Neslab controller in stock so
we swapped it out with the one on the chiller
due to the broken switch. However, new controllers
need calibrations which we didn't have documentation
on so therefore, the new controller was floating between
17C and 28C. We ended up swapping out the on/off switch
on the old one with the new which is really the only
thing wrong with it. The chiller is now all set and
stabilized at 24C.
D shift
|
19.321 | UNLOAD FLIPPER MISSED A WAFER | FABSIX::B_WESTPHAL | | Sat Mar 15 1997 18:21 | 5 |
| CMP.B2 - Down for the unload flipper missing a wafer. Rebooted the tool and
ran a boat of wafers on the no_touch recipe to observe the unload
alignments. The flipper unloaded fine. Returned the tool to
production.
|
19.322 | | FABSIX::M_RIDLON | Eliminate the oblivious | Mon Mar 17 1997 05:57 | 6 |
| Came onto shift with A-shift working on carrier to cup alignments after
a crash. We were able to get the alignments in returned the tool to
production. It passed mqc's and hasn't had a problem since.
A & B shifts
|
19.323 | new doors | SUBPAC::CMP | | Tue Mar 18 1997 14:55 | 4 |
| Today we did the new style door upgrade. They work a lot nicer than the
old style.
Just a reminder, make sure that the new doors are fully
positioned closed to insure the interlocks are working.
|
19.324 | EOL for pads and carriers | FABSIX::R_POIRIER | | Wed Mar 26 1997 06:25 | 10 |
|
B2 reached EOL on pads and carriers. Changed
both pads and the carriers and did weekly Pm.
B Shift
|
19.325 | Timeout for one or more carriers not down | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri Mar 28 1997 08:33 | 20 |
|
Tool was running product when the tool timed out
for "one or more heads not down." Same errors as B1????
It was in segment #3 before downforce was applied.
The carriers and pad kept rotating while the screen and
process clock were froze. The INPUTS for the carriers
were checked and they were all lit, indicating that they
were all down. The tool was EMO'ed. After recovery,
the heads were cycled up and down 15 times and all of the
switches were being made. OPS ran 15 wafers for VERIFY
and they ran fine. On a side note, carrier #1 was missing
(4) screws out of the retaining ring. They were missing
in every other slot so I am leaning towards the fact that
they were never installed. I checked around under both
tables and there were no screws to be found. Please be sure
that when you install new carriers that all of the hardware
is present and at proper torque.
BJ
|
19.326 | Load Flipper Timeout Error | ASDG::POIRIER | | Wed Apr 02 1997 20:32 | 10 |
|
Load Flipper Timeout:
--------------------
The system errored for "load flipper timeout" once this afternoon.
Pressing the error message on the touch screen resumed the auto cycle.
FYI: The identical error occured last Thursday on C Shift. I don't
remember documenting it at that time.
C Shift.
|
19.327 | Small D.I. Water leak under index table | ASDG::POIRIER | | Thu Apr 03 1997 20:25 | 8 |
|
Leak Detection Sensor:
---------------------
The system errored for leak detection tonight. The sensor under the
index table was activated. A small puddle of D.I. water was wiped up
around the sensor. We were unable to determine the source of the leak.
C Shift.
|
19.328 | Load Flipper Timeout Error | ASDG::POIRIER | | Fri Apr 11 1997 20:16 | 12 |
|
LOAD FLIPPER:
------------
(This information was left out of yesterdays passdown)
The system experienced a "load flipper timeout error".
The load flipper failed to lift a wafer out of the
pre aligner. The alarm was silenced using the touchscreen
and load flipper resumed normal operation.
FYI: This exact error occured last week on C Shift.
|
19.329 | Dropped wafer from carrier #1 | FABSIX::R_POIRIER | | Sun Apr 13 1997 08:49 | 28 |
|
B2 dropped a wafer out of carrier #1 during
segment 7. Wafer did not break and ended up
in lower tub area. Wafer detect went off and
"main vacuum low" alarm. Screen was frozen
so we had to re-boot. Unloaded all wafers
and removed carrier #1. Did Load cup alignment
check for Cup #1. Alignment was fine. Checked
Step height for the carrier it ranged from
.003 to .008. Also noticed white ring was
scratched in a couple of places. This could
have happened when it lost the wafer. We then
replaced the carrier and returned tool to ops.
When they tried to run break-in a new alarm
occured. Time-out waiting for carriers down.
Carriers had loaded wafers and went out over the
polish table, carriers and table started rotating
but the carriers wouldn't go down onto the table.
Couldn't clear problem, so we had to re-boot the
tool. Started new set of break-ins and tool ran
fine.
B Shift
|
19.330 | Some problems with B2 | FABSIX::R_POIRIER | | Mon Apr 14 1997 08:40 | 36 |
|
We had a few problems with B2 tonite.
During a run, the unload cup missed
a wafer. This only happened once.
Also during a later run Carrier #2
didn't pick up a wafer at load. Hand
loaded and continued run. When run
was complete, we checked load cup
alignments. All were ok. Then on the
next run multi-head loaded wafers,
went out to polish table, started to
spin, as did the table. But carriers
never came down. Kept getting time-out
errors fro carriers not down and tool
would not let us quit run. Had to do
a soft boot. Next, during the following
run tool froze up. Unload flipper got
lost and didn't know what to do. Tool
timed out waiting for unload flipper.
I hit start and tool finished it's run.
When operator was done, I exited out to
DOS and shut down the tool for five mins.
Powered it back up and tool ran fine rest
of the night.
Except for Carrier #1. The removal rate for
#1 is way off the other four (higher). They
ran MQC's and verified it was high. They adjusted
the off-set, re-ran MQC's. Still out. But just
barely out. Awaiting A shift to decide on either
replacing #1 or adjusting Off-set again.
B Shift
|
19.331 | This was from A shift... it was CMP.B2 that had the problem | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Wed Apr 16 1997 08:39 | 27 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
-< CMP >-
================================================================================
Note 17.549 CMP.B1 EE/PE PASSDOWNS 549 of 549
SUBPAC::LANDRY 20 lines 16-APR-1997 07:15
-< wafer broke, computer mishaps, fuse blown >-
--------------------------------------------------------------------------------
This is a late entry, should have been entered 12 hours ago...
The system broke a wafer IN THE POLIS AREA, but NOT on the table.
A wafer flew out of one of the heads, made it off the table and broke
in the tub rinse section. No data was captured, as the system was
EMO'd. We verified head alignments, which weren't really bad, but
we did bring the Index Table over a couple of hundred motor counts.
After starting up the system, the system was initialized. But things
weren't well. The Multihead got lost, and slammed against the stops
at the polish table end. EMO'd that problem. Rebooted. Then
experienced problems with initialization with the Inverter. That was
traced to a faulty door switch. We got by that, then the system just
powered inself off. Would not repower up. Found a 3 A fuse on T2 had
blown. The system was turned over at this time to the oncoming shift.
Carriers have been removed and placed in water. The pad is also keep
wet with a spray gun. Good luck with this baby.
A shift
|
19.332 | Fuses... interlocks and bears oh my!!! | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Wed Apr 16 1997 08:44 | 51 |
|
Oh what a night... last December back in '63 , what a
polisher, what a night...
Came in to clean up the tool after a most wonderful
crash from A shift. Apparently a product wafer was left
on the table after lift-off and slid off the table where
is miraculously fell just right... Right into the drain
overflow track of the table and one of the PVC stems blew
away the wafer... Ripley's - Believe it or not!!!
The tool wasn't able to come back up because the tool
lost its 24VAC. That problem was traced to a blown 3A
circuit limiting slow blow fuse on one of the legs of
primary 240/24 AC line. Wow... That fuse was replaced
and the tool started working again. Whilst running some
VERIFY wafers, the tool errored for Inverter Comm. error
and rotation errors. Powered down tool and found the OTHER
3A fuse on the OTHER leg of the primary circuit was tripped.
that fuse was replaced as well, however.. we installed the
same fuse except the tool fuse holder requires an indentation
at the end of the fuse to properly go into the fuse holder.
The one that we found didn't have the indentation so we swapped
the 30A fuse and incoming/outgoing wires with the 3A wires and
fuse to accommodate the fuses. Rich B. will be getting the
proper fuses in the a.m. as there are none in stock. After
the fuse was replaced, the 8-sec relay wasn't pulling in so we
looked at the process doors. This is the only thing that is
in the circuit along with the wafer detect board that could
cause the problem. The wafer detect board was easy... That
was working fine and wasn't causing the problem. We tried
the process doors and looking at the tool, the right front
door switch was dead. We verified this by jumping the switch.
The switch was replaced and the tool started running again.
The tool ran for 2 hours then went down again for the same
Inverter Comm. error. The 8-sec relay was working after I
powered the tool back up. We went through all of the doors
again and found that if you pull on the rear doors, the right
switch would break contact. The switch was set for better contact.
The tool started running again, until it went to unload the wafers
from the first run. The flipper had the wafer but the computer
didn't think so. The INPUT wasn't lit. Another softboot
brought the flipper back to life. We are now attempting another
VERIFY...
D shift
|
19.333 | flipper fix | SUBPAC::CMP | | Thu Apr 17 1997 14:11 | 6 |
| Late note.
Worked with an unload flipper problem for flipper not open. Found the
pivot point target was not being seen by the sensor. Install shim
on the pivot point to bring the target closer to the sensor. Tested
and returned to production.
|
19.334 | Carriers & pads replaced - Eng. rings on #3 and #4 | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri Apr 18 1997 00:30 | 11 |
|
Both pads were replaced due to primary step height,
wafer count on both at 1890.
All (5) carriers were replaced. The two special
carriers with prototype clamp rings are on heads
#3 and #4.
the delrin ring carrier is on #3 and the PPS is on #4.
BJ
|
19.335 | New PPS ring experiment a flop... | FABSIX::L_SHKOLNIKOV | | Fri Apr 18 1997 06:56 | 32 |
|
During the first run after pad change the tool was alarming for
wafer detect activated. All wafer detect sensors have been checked.
Found sensor #4 pointed on the very edge of the carrier, so had to
adjust this one and rebooted the tool. By the the way, carriers #3
and #4 are experimental and were placed on the tool per EE request.
Both pads were replaced due to pad life and OPS were using the
break-in recipes. When the OPS went to run the I/O-BRKIN recipe,
the tool errored for carrier #4 not rotating, 6 sec. into the seg.#5.
Checked wafer detect - looked fine. Powered tool down. Tried to power
it back up - got no power on the screen. Monitor turned on but
nothing on the screen. All voltages, infranors,interlocks and
whatever else have been checked everything fine except CPU board
which has two LED's "pass" and "fail" . Both of them were ON. After
the switching the CPU reset switch several times and reseating the
CPU board itself the screen started work again. Started running breakins
During the second run got 6 sec. into seg.#5 got the same error carr#4 not
rotating. So we rebooted the tool again and swapped carr#4 with carr#1.
The tool errored for carr#1 during the next run. The wafer exposure has
been verified and we got "+.035" instead of" -.075" and new experimental
guide ring was polished all way down and looked nice and shiny. Since
the tool had run ~6 wafers in that carrier alone, the original exposure
could of been up near +7. Since B1 is going down in the a.m., we
put a near carrier on #1 and took the PPS style ring off and got
them running again. The delrin ring is still on #3. The PPS style
carrier is in the chase next to A1. Please leave for Bruce to
inspect.
Is Hal-Bop gone yet??
D-shift
|
19.336 | PROBLEMS PASSING MQCS | FABSIX::B_WESTPHAL | | Fri Apr 18 1997 21:46 | 8 |
|
CMP.B2- Down for losing a wafer and failing MQCs. The tool lost a wafer out
head #3 which was an experimental carrier. Because of the problems
the previous shift had W/ the experimental carriers it was decided to
put a regular carrier on. The tool then failed for head to head rates
and a questionable removal rate so a new pad was put on. The tool then
failed for particles but operations is running the MQC again.
|
19.337 | Slurry delivery problem | FABSIX::O_WHITFORD | Lead, follow or step aside. | Sat Apr 19 1997 07:22 | 18 |
| Tool errored for carriers 1 & 2 not rotating.
Excessive vibration prior to error. Ops EMO'd tool and when
investigated further, no slurry was on the polish table and slurry line
leading to MHTA were full of water and NOT slurry.
Checked slurry flow and adjusted for 500 mL/min. It was initially found to
be @150 mL/min low.
While observing tool operaton, it was noted that the flex track on the
right side of the MHTA forward edge was broken in several places and
the end point was shattered. Replace several links and the end point.
Now runs straight and true. Cleaned a lot of debris abraided from cable
insulation. Looks much better.
Ops ran 10 wafer verify with no problems. Back to ops.
D SHIFT
|
19.338 | Just for a short time. | FABSIX::L_SHKOLNIKOV | | Sat Apr 19 1997 08:35 | 5 |
| Almost forgot. The end bracket which connect automation frame and flex
track on cmp.b2 and actually was causing the problem we took from CMP.B1
temporarily till new bracket will rushed in.
D -shift.
|
19.339 | Wafer Breakage - Removal Rate on Head #1 | ASDG::POIRIER | | Sat Apr 19 1997 20:41 | 22 |
|
Wafer Breakage:
--------------
The system broke 2 monitor wafers at 8:05 a.m. this morning.
The system was 6 seconds into segment #7 of the IO_TUP recipe
when the crash occured. The wafer slid out of carrier #4 and
hit the wafer in carrier #3. The pad and 2 carriers were replaced.
The tool ran for the rest of the shift with no further problems.
Removal Rate out of Range - Carrier #1:
--------------------------------------
The removal rate on carrier #1 was way out of range while the other 4
were close. The carrier was replaced and the tool passed MQC's.
C Shift.
|
19.340 | Crash on secondary table | FABSIX::R_POIRIER | | Sun Apr 20 1997 08:39 | 24 |
|
B2 dropped a wafer out of carrier #2 or 3 while
on the the secondary pad. Don't know what segment
tool was in as it was EMO'ed. Carriers thought
they all still had wafers and unloaded and loaded
new wafers. Tool errored when Unload fingers didn't
find a wafer in unload cup 2. Meanwhile multi-head
went out to polish table and started polishing.
When operator saw the wafer crash in the index table
area he EMO'ed the tool.
We cleaned up the index table area and replaced the
pad. The wafers on the carriers all came off and
were on the polish table. A product wafer was broke
when I tried to lift it off the table. All carriers
and the pad were replaced. Checked load cup alignment
for #2 and 3. Both ok. Cleaned up, re-booted and
returned tool back to Ops.
B shift
|
19.341 | Door Switch strike again! | FABSIX::T_LANDRY | | Tue Apr 29 1997 20:10 | 22 |
| Today, we had to replace an Infranor on carrier 2 rotation.
It all started with a weird senario. The tool errored for a MHTA
Timeout. The head didn't move to the polish section at all when
commanded. Pressing a button to continue, it jumped to the Index Table
movement, where we proceeded to also get a Timeout. Not seeing any
other alarms exhibited, we exited out to DOS and powered down to
reboot. Upon reboot, we had a serious problem. We couldn't get the
screen to appear. So, we decided to check the CPU breaker, which was
fine. Having seen this problem elsewhere before, we checked the door
interlocks. Sure enough, the back double doors switches were not being
made. The doors were closed, but while pushing on the bottom, the
latches kicked in. With this door open, it quickly explained the
timeouts on the MHTA and Index Table. Somehow, the door switches
joggled out of position. Anyway, we powered down for a while and
rebooted. Upon this reboot, we got a rotation error on carrier 2.
The Infranor was dead, no power what so ever. We found one in the room
labled as possibly good. It was installed and tested fine. All is
back to normal.
A shift
|
19.342 | Carrier 2 problem | FABSIX::R_POIRIER | | Tue May 06 1997 07:40 | 18 |
|
Tool alarmed for robot not picking up wafers.
Then alarmed for "Waiting for wafer to break beam"
Screen froze up. We had to re-boot tool. While
rebooting, Carrier #2 would not rotate. Found the
Infranor to be bad. Replaced Infranor and re-booted.
Carrier #2 would start to rotate but then just stop.
Tool would not alarm. Tried to rotate carrier using
Manual screen, no luck. Infranor is fine. Disabled
carrier #2 and returned tool to operations, due to
heavy work load. They are running MQC's using 4
carriers.
B Shift
|
19.343 | carrier two rotation (revisited) | FABSIX::J_SADIN | Freedom isn't free. | Tue May 06 1997 14:02 | 13 |
|
Carrier 2 rotation problem (continued from B-shift):
Swapped rotation infranors 1 and 2 and the non-rotation problem
followed the infranor. Replaced infranor and all is well. Pad and
carriers replaced due to drying out/high counts. Also found connector
to oscillation motor (on top of multihead) was loose. Reconnected and
tightened.
a-shift
|
19.344 | Carrier #1 rot infranor bad | FABSIX::A_QUILLEN | Nobody will get away with anything | Wed May 07 1997 08:27 | 5 |
| System went down for carrier #1 not rotating. The infranor was out and
remained so following re-boot and plug connections check. Replaced same
and system is back to production.
B shift
|
19.345 | Double loaded wafer, broke both! | STRATA::STANDING | | Sat May 24 1997 19:57 | 6 |
| Tool went down hard for breaking two wafers. As near as I can tell
carrier #1 loaded up two wafers and broke the first one over the index
table, then broke the second one over the polish table. Cleaned up the
mess and replaced both pads and all five carriers. Ran a boat of 25
using the notouch rcp and ran with out error. Returned tool to
production.
|
19.346 | Pad change | FABSIX::R_POIRIER | | Sun May 25 1997 06:10 | 7 |
| Tool failed MQC's for scratches. Replaced carriers,
and both pad. Also cleaned index table. Gave tool
back to OPs and it passed.
B shift
|
19.347 | overtravel | LUDWIG::FINN | | Mon Jun 02 1997 20:30 | 20 |
|
System shuttered when it failed to see either sensor for table at
polish and kept driving. Operator EMO the tool to make the room stop shaking.
Powered up and system ran fine in manual out and back. We noticed the bottom
scrub brushes were not spinning. We began to address this unrelated error and
disassembled the scrubber to check the belt. Turned out to be a bad belt.
Meanwhile back on the farm, we check out the multihead proximity sensors and
verified the operation of the inhibit switches which should protect system
from overtravel. They all work fine! Had numerous boot problems i.e. CRT
screen not coming on, elevator 1 not homing (not seeing home), and carrier 3
infranor showing no power or error light. Power supply to infranor was good.
We replaced this infranor and the problem of the dead infranor switched to the
pad conditioner. What the ....! System acted like it was randomly not seeing
i/o inputs. Reseated DCX Card and CPU Card. Booted aok and ran fine. Oh ya the
problem seen on front and remote screens cleared if we rebooted with the
switch boxes in the A positions.
A shift
|
19.348 | Moisture sensor alarm | FABSIX::R_POIRIER | | Wed Jun 04 1997 07:13 | 13 |
|
We found a large puddle in the bulkhead (right side)
that tripped the moisture sensor. Traced leak up to
connector for the hand held spray line. Tightened
the connector, cleaned up the water and returned the
tool to Operations.
B Shift
|
19.349 | Defective scrubber motor. | LUDWIG::STANDING | | Thu Jun 05 1997 18:47 | 4 |
| CMP.B2 Went down due to the scrubber brushes upper and lower set not
turning. Disassembled unit and found the gear on the motor shaft had
stripped out. Replaced defective motor, reassembled scrubber. Checked
out operation of motor OK. Returned tool to production.
|